A Complete Bibliography of The Journal of the American Statistical Association: 1950–1959

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

09 November 2023 Version 1.04

Title word cross-reference

0.05 [Dun57b]. 1/X [GS54a, GS54b]. $500 [Wei50]. α = 0.01 [Dun57b]. χ^{2} [Ber55b]. F [Dun57b]. G [JR55b]. H [Wal59a]. R_{1.23} [WF57, WF58]. t [Fed59, Wal50a, Wal51].

-Distribution [Fed59]. -Statistic [Wal51]. -Test [Wal50a].

1-6000 [Ano52k]. 100th [Sto58b]. 109th [Ano50l]. 110th [Ano51h]. 111th [Ano52f]. 112th [Ano53h, Alc53]. 113th [Ano54g, Alc54]. 114th [Alc55]. 115th [Alc56]. 116th [Alc57]. 117th [Alc58b]. 118th [Alc59]. 126-17 [Ano54k]. 1900- [Wil55a]. 1939-41 [Yer55, Yer56]. 1949 [Dub50]. 1950 [Mal50, MMN53]. 1951 [Ano52f]. 1952 [Alc53, Ano53h]. 1953 [Alc54, Ano54g]. 1954 [Alc55]. 1955 [Alc56]. 1956 [Alc57]. 1957 [Alc58b]. 1958 [Alc59].

376-20 [Ano54l].

6001-6100 [Ano53l]. 6501-6875 [Ano53m]. 6876-8125 [Ano53n].

876-15 [Ano54j]. 876-21 [Ano54m].

9001-13 [Ano53o].

Abbreviated [Ber57]. Ability [Tho52]. Absolute [HGCZ55]. Abstracts [Ano51a, Ano54n, Ano54o, Ano54p, Ano54q, Ano55d, Ano55l, Ano55m, Ano56l, Ano56m, Ano56n, Ano56o, Ano57m, Ano57n, Nic55, Woo51a, Woo51b, Woo51c, Woo52a, Woo52b, Woo52c]. Acceptance [Bre56, GF59, Tuk50a]. According [Ask57]. Accounting [Jon55b]. Accumulation [Fab52]. Accuracy [CR56, Fer53, Fer58a, GF59, HM58, Mye54a, Raj58, Whi54, Win59]. Activity [Sav54b]. Actuarial [Elv58b, PJ57a, PJ57b, Spo51]. Additional [Ano51a]. Additives [WK57]. Address [Hoa59]. Adjusted [Kar50]. Adjusting [EBS55]. Adjustment [Key51, SM58]. Adjustments [Eis56, SE57]. Advances [And53]. Advancing [Ask59]. Advertising [Gil50]. Aeronautics [Pre55]. Affecting [GWS50]. Affiliation [Lan59b]. Affiliations [Lan59a]. Africa [Sha52a]. after [Mye54b]. Against [Eps55, Stu54, Stu56, Noe56c]. Age [Coa55b, Coa55c, Man53, Mye54a]. Agricultural [Lee52]. Agriculture [Cro59]. Air [LB58]. Alfred [Dub50]. Allied [Gri59a]. Allocations [Eva51a]. Alternative [KM51, Wei58]. Alternatives [Eps55, Stu54]. American [Alc53, Alc54, Alc55, Alc56, Alc57, Alc58b, Alc59, Ano50l, Ano51h, Ano52f, Ano53h, Ano54g, Fit57, Fit58]. among [Edg56]. Amount [Ham59]. Analyses [Bur58]. Analysis [Ade58, Ade59, And54, And55a, BAR56, BGR56, BPR55, BS52, BK57, Bro52a, Coh55a, CG59, Dav52, Dun57b, Dyk58, Gil50, GG56a, Goo52, Goo54, Gre52b, GS59, Hou54b, Joh50b, JN51, Kin54, KW52, KW53, LB58, Lom54, Mei54, RGM58a, RGM58b, Sch52a, Sto55b, Tei55, Wag59b, Zel57, Zel58, GG56b, GG67, Bue56]. Analytic [Dem53b]. Analytical [Lah56]. Analyzing [GWS50]. Angeles [Wil55a]. Animal [Ree52]. Anniversary [Sto58b]. Annotation [Ano53d]. Annual [Alc53, Alc54, Alc55, Alc56, Alc57, Alc58b, Alc59, Ano50l, Ano51h, Ano52f, Ano53h, Ano54g]. Annum [Wei50]. Antisyphilitic [BDL^{+}54]. Application [Bab55, CM50a, CM50b, Eps55, GS59, JM57, Jon55a, MG55, Met50b, Spr50]. Applications [Coh55a, Gum54, Kim55a, Kim55b, Kru54, Min57a, Net52, Rob57b, Tat55, Tuk52]. Applied [Cro52, Eis56]. Apportioning [Wil54, Wil56b]. Appraisal [Abr58, Hef51, Mil53]. Appreciation [Mat51, Sto58b]. Approach [Foo58, Old54]. Approaches [Sim58]. Approximate [BH54]. Approximating [Jon53, Raf56b, Raf56a]. Approximations [Wal59a]. Aptitude [Vot52]. Arbitrary [Adl57]. Area [Bue56]. Areas [Coe55a, Coe55b, Vin53]. Arising [Lor59]. Ashby [Hou53].

Aspects [CM50a, CM50b, DN58, Mos56b, Smi54]. Assay [Ber55a]. Associate [BS52]. Associated [Kim53]. Association [Alc53, Alc54, Alc55, Alc56, Alc57, Alc58b, Alc59, Ano50l, Ano51h, Ano52f, Ano53h, Ano54g, Coc54b, Cut55, GK54, GK57, GK58, GK59, Kru58b, War56]. Asymptotic [Stu54, Wal56c]. Atlantic [Alc58b]. Attain [Dem53a]. Attempt [PS50]. Attitudes [Jar58, Mad51]. Attributes [HV55, RK52, RK53]. Auditing [Net52]. Authors [Ano58c, Ano58d, Ano58e, Ano58f, Ano59h, Ano59i, Ano59j, Ano59k]. Autocorrelation [And54, And55a]. Automatic [Fry56, Loc59]. Automobile [Raf51]. Autonomous [NY59a, NY59b]. Autoregressive [Wei55, Wei58]. Availability [BS50]. Available [Cla59, Lee52].

Back [Ano50a, Ano50b, Ano50c, Ano50d, Ano51b, Ano51c, Ano52a, Ano53a, Ano53b, Ano54a, Ano54b, Ano55a, Ano55b, Ano56a, Ano56b, Ano56c, Ano57a, Ano57b, Ano57c, Ano57d, Ano58a, Ano59a, Ano59b, Ano59c]. Balance [Dem53a, Fab54]. Balanced [BS52]. barX [Dun56a]. Based [Ask59, Ber53a, Cho59, Kar58b, Mit57a]. Basic [SM55]. Basis [BD59, Foo53]. be [Cro52, GG56a, GG56b, GG67, Jon59]. Bean [Foo53]. Bear [Tho52]. Bell [Bue56]. Berkeley [Met50a]. Bernoulli [GS54a, GS54b]. Best [And53, Ask57, GW56, SG57, Wal50a]. Beta [Cla53b, Wal59a]. Beta-Approximations [Wal59a]. Between [Bla54, Coc55, Cut55, Dem53b, Dem53a, Elk53, Fox51a, Gea53, Kit55, Kit56, OS59, dJ59]. Bias [BS50, Dem53a, Sob59, Ste51b]. Bibliography [Ano53c, Ano53d, Sav53, Sav58a]. Biltmore [Ano50l]. Binomial [Bue57, Coh59a, Mac59c, Mac59b, Raf56b, Raf56a, Rid55a, Rid55c, Som57]. Bio [Ber55a]. Bio-Assay [Ber55a]. Bioassay [Ber53a]. Birth [Sha50, Sto58b]. Biserial [Tat55]. Bivariate [Blu58b, Coh55b]. Blackjack [BCMM56, BCMM59]. Block [BS52, Zel57]. Blocks [KB53, Raf51]. Book [Ada54, Adl55, Alc58a, Ale57, Alt58, And50, And56, And55c, Ang57, Ans56, Ans59a, Ara52, Ark56, Arn52, Arr54a, Arr54b, Arr58a, Arr58b, Atk55, B.56, B.57a, B.57b, B.58, Bac53, Bac58, Bah57, Bai58, BO53, Bak56, Bal56, Ban58, Bar58, Bea57, Bea54, Bea50, Bel55, Ben57, Ben50a, Ben50b, Ben55, Ber56, Bey52, Bir55, Bir56, Bla51, Bla58c, Bla56, Ble58, Blu58a, Blu54, Bod57, Bod59, Boh55, Bor57, Bos58, Bou59a, Bou59b, Bow50a, Bow50b, Bow56, Box55, Bra56a, Bra51a, Bra51b, Bra54a, Bra56b, Bra58c, Bra59a, Bra59b, Bra55, Bre58, Bri50, Bri59, Bri56, Bro58a, Bro53a, Bro58b, Bro59, Bro53b, Bro56a, Bro56b, Bro52b, Bro52a, Bro57a, Bro58c, Bru55, Bry57, Buc54, Bud56, Bue56, Bur55, Bus56a, Bus56b, BC56, Cam57, Can50, Cha51, Cha52, Che55a]. Book [Che55b, Chi54, Chi59, Chi53, Chu57, Chu51, Cla50, Cla52, Cla53a, Cla57b, Coa55a, Coc52, Coc54a, Coc56, Coc57, Coh56, Coh57, Col57, Col56, Com59, Con52, Con56, Coo51b, Coo58, Cor50, Cor51, Cor55, Cor58, Cov51, Cox59, Cra50, Cra53a, Cra53c, Cro50, Cur54, Dal54, Dar51, Dar55, Dav59a, Dav59b, Day51, Day59, De 51, Dem55, Dem58, Den55, Der56, Des50, Dig57, Don56, Dor54, Dor50a, Dor56a, Dor56b, Dor56c, Dor55, Dow54, Dre58, Dud52, Dun54, Dun57a, Dun59a, Dur51, Dwa54, Dwy51, Dwy59, Eck58, Edw58, Eis53, Eis55, Eis59, Elv58a, Enk56, Eps54, Eva57, F.51a, F.51b, F.51c, F.51d, F.56a, F.56b, F.57a, F.57b, F.57c, F.58a, F.58b, F.58c, F.59, Fan58, Fel57, Fer55a, Fer58b, Fer50, Fey59, Fic55, Fin57, Fir51, Fis51, Fis54, Fol54, Foo59, Fox51b, Fox53]. Book [Fra51, Fra54, Fre50a, Fre51, Fre55, Fre59a, Fre50b, Fre59b, Fry51, Ful55, G.53, Gar52, Gar55, Gar56b, Ger53, Gil57a, Gil59, Gil57b, Gin52, Gir53, Glo56, Gog54, Gol58, Gol55a, Gol55b, Goo56, Goo53a, Goo58b, Gor56a, Gor56b, Gor58, Gou52, Gra59, Gra58, Gre52a, Gre53a, Gre52c, Gre56a, Gre57, Gre58a, Gre55, Gre58b, Gre54, Gre56b, Gre52d, Gri57a, Gri57b, Gur54, Gut56, Had55, Hae55a, Hae55b, Hae55c, Hae57a, Hae57b, Hae57c, Hag52, Hal50, Hal55a, Hal53, Hal55b, Ham54a, Ham52, Ham54b, Han56, Har51a, Har56a, Har55, Har59, Har52, Har51b, Har53, Har56c, Has56, Has51, Has53, Hau56, Hei56, Hen56, Hen57, Her57, Hic55, Hil55a, Him55, Hir55, Hoa50, Hoa57, Hod52, Hod55, Hoe55, Hol51, Hol58, Hoo57, Hor57, Hor55, Hos56]. Book [Hou53, Hou56, Hou54a, How57, How58, Hug52, Hul56, Hym57, Hym58, Ire52, Ire58, Isa57, Jac53, Jac56, Jaf52, Jaf56, Jah58, Jah54, Jam50, Jar56, Jeb51, Joh50a, Joh54, Joh56c, Joh56a, Joh56b, Jon51, Jon52, Jon56c, Jon56a, Jon57, Jon58a, Jos58, Kah56, Kai59a, Kal51, Kap59a, Kar56, Kar58a, Kat55, Kat56a, Kat56b, Kat56c, Kem54, Ken52, Ken56a, Ken56b, Ken58, Ken57a, Ken57b, Ker56, Kes56, Kim56, Kla58, Kle57, Kle52, Kni50, Kno54, Koc52, Koh56, Kon52, Kos53, Koz55, Kra59, Kru51, Kru58a, Kru59, Kul58, Kur51, Kuz50a, L.51a, L.51b, L.51c, L.54a, L.54b, L.55, Lac56, Lan56, Lau55a, Lau55b, LeN56, Lee51, Lee53, Lee56b, Lee56a, Lee59, Leh53, Ler57, Lev57, LF54, Lie53, Lie55, Lie58a, Lin54, Lin58, Lin59, Loc54, Loe51, Lof50, Lon56, Lor53]. Book [Lor52a, Lor52b, Low58, M.54, M.59, Mac52, Mac59a, Mac56, Mac55, Mad53, Mad55, Mad58a, Mad58b, Mad58c, Mai58, Mai56a, Mai56b, Mai54, Mal53, Man52, Man58, Mar58a, Mar57, Mar52, Mar53, Mar56, Mar58b, Mas59, Mat57, Mau58, May55, May58, May50, McA59, McC57, McD52, McG58, McI57, McI55, McP58, Mei58, Mer50, Mer51, Met50a, Mey59, Mic59a, Mil56b, Mil56c, Mil56d, Min57b, Mit51, Mit57b, Mod56, Mor54, Mor55, Mor51b, Mos51, Mos56a, Mur55a, Mut58, Nat57, Net58, New57, Nic51, Nis54, Noe56a, Noe57a, Noe59, Nor52a, Nor55b, Nor54a, Nor54b, Nut58, Nut59, O’R53, Old51a, Old55, Olk57, Orc55, Ost54, Ott57, Pab50, Pab55, Pad50, Pag57, Pal53, Par56, PW57, Par59, Per59, Pet55, Pet57a, Pet57b, Pfo55, Pot56, Pri53, Pri54b, Pri54a, Pyk57, R.51]. Book [R.53, Raf57, Rai55, Rap58, Ras56, Rau58, Rei52, Rei53, Rei58, Res58, Rev50, Ric56, Ric57, Ric59, Rio56, Rob55, Rob57a, Rob58, Rog56, Roo51, Roo52, Ros57, Ros59, Ros52, Rot55a, Rot55b, Rub58, Rud53, S.51, S.58, Sar56, Sav51a, Sav55, Sav56, Sav58b, Sav59b, Sav57c, Sax56, Sch55, Sch56a, Sch57a, Sch57b, Sch50a, Sch50b, Sch56b, Sco58, Sei56, Sew51, Sha57a, Sha52b, Sha57b, She58, Sho55, Shr54, Sie57, Sie59, Sil56, Sim51a, Sim53b, Sim53c, Sit53, Sle51, Sle57, Smi50b, Smi50c, Smi52, Smi55a, Smi55b, Smi58c, Smi58b, Smi58a, Sne51, Sne55, Sny56, Sol55a, Sol55b, Sol56a, Sol52b, Sol52a, Sol55c, Sol56b, Spe57, Spi51, Spi55, Spi57, Spr59a, Spr53, Spr59b, Ste57a, Ste52a, Ste51a, Ste52b, Ste53, Ste56, Ste50, Ste55a, Ste54, Ste58]. Book [SB52, Ste57b, Sti51, Sti52, Sto58a, Sto55c, SH51, Str53, Sui53, Sui58, Sup55, Sup58, Swa54, Sza54, Tae56, Tei56, Tei57, Tep50, Ter56a, Ter56b, Tho59, Tic58a, Tic58b, Tin58a, Tin58b, Tip52, Tip55, Tip59, Tol55, Tor52, Tru56, Tur57, Tyl54, Tyl58, Ull52, Ulm50, Val58, Van55, Vil52, Vil53, Vin51, W.51a, W.51b, W.51c, W.52a, W.52b, W.52c, W.52d, W.52e, W.53, W.54a, W.54b, W.54c, W.54d, W.54e, W.55a, W.55b, W.55c, W.55d, W.55e, W.56a, W.56b, W.56c, W.56d, W.56e, W.57a, W.57b, W.57c, W.58a, W.58b, W.58c, W.58d, W.58e, W.58f, W.58g, W.59, Wag59a, Wal56a, Wal57, Wal55, Wal56b, Wal58b, Wal54, Web59, Wei51, Wei59, Wei57, Wel54, Wel57, Wes58, Wes55, Wes51, Wil52, Wil56a, Wil55b, Wis58, Wit56, Wol51, Wol52, Wol55, Wol59, Woy50, Wri51, Wri52, Wri56]. Book [Wro56, Wyc58, Yer57, Yod52, Zei59, Zio55, Zuc57, dC56]. Books [Fel59]. Boston [Ano52f]. Both [Adl57, Mad59a, Mad59b]. Bound [Ber61, Wil59]. Bounded [Rei56]. Brain [Hou53]. Brains [Met50a]. Brazilian [Ste55b]. Bridge [WW53]. Briefer [Ano50e, R.50]. Britain [Mos55]. British [HR58]. Buffalo [Yer55, Yer56]. Building [Bla54, Wil55a]. Burdens [Tho52]. Business [Ano50l, Ano51h, Jon55a, LB58, Lom54]. Businessmen [Fer53].

C [Met50a]. Calculation [CM50a, CM50b, Kar50]. Calendar [Eis56]. Callbacks [PS50]. Canada [Mar55]. Cancer [BG52, Ber58, Cut55]. Cannot [GG56a, GG56b, GG67]. Car [dJ59]. Card [Gre53b, Ros53, Ros54]. Card-Duplication [Ros53, Ros54]. Cards [AC55, BAR56, BGR56]. Care [Dea52]. Cargo [Dee54]. Case [Adl57, Fis52, MMN53, War56]. Cases [Cro52]. Causal [Sim54]. Cause [Yer55, Yer56]. Cause-Specific [Yer55, Yer56]. Cells [Cra53b]. Censored [Coh55a, SG57]. Census [Coa55b, Coa55c, CG59, EBS55, Fol53, HHNS55, MM50, Mil53, Mye54a, Nor55a, Ste51b, VK52, Win59, Zar55, MMN53]. Cent [GS50]. Central [Sha52a]. Centroids [Ask59]. Century [Fit57, Fit58]. Certain [Cla59, Gre53b, Kai59b, Kat53, Man57]. Chairman [Tuk50a]. Changes [Ade55, Bur53, Key51, Mil51]. Characteristic [Dun57b, Gar56a, Kat53]. Characteristics [Wal52a]. Charlier [Ber57]. Chart [Wei54]. Charts [AL50, AL52, Dun56a, Dun57b]. Check [Kai59b]. Chemical [Dan53]. Chi [Ber55a, Coh55a, Ram58a, Ram59, Wil50]. Chi-Distributed [Coh55a]. Chi-Square [Ber55a, Ram58a, Ram59, Wil50]. Chicago [Alc53, Alc59, Ano51h, Ano53h]. Choice [Bro50, Wal50a, Wil50]. Chronic [Elv58b]. Circular [GGD53, Gum54, Mor59, Mor60]. City [Alc56, Alc58b]. Clair [Ber53b]. Class [Elk53]. Classes [BS52, Elk53, Sog54, Wil50]. Classification [BS52, Sim53a]. Classifications [GK54, GK57, GK58, GK59]. Classified [Coa55b, Coa55c]. closure [Tuk50a]. Cochran [CMT55]. Coefficient [Cow52, Gri58, Wei58, Wil59]. Coefficients [Cow58, Cow59, Dur54, Dur53, Lor55b]. Coffee [Ste55b]. Collection [CL54b]. Colonial [Jon55b]. Color [Coa55b, Coa55c]. Combination [Cra53b]. Combinations [Wel56]. Combining [Bir54a]. Comment [Tul59]. Comments [Wal59b, Whe50]. Commission [Ade55]. Common [Ban50]. Communities [Tho52]. Compact [Mac59c, Mac59b]. Comparing [Dun55, Hyr53a, Hyr53b]. Comparison [Bel54, Dun55, Eps55, Mor59, SAF54, Ste59a, Tho52, Mor60]. Comparisons [Dyk58, Rot52, Sch52a]. Competition [Mos52]. Complete [Bra58b, Bra58a]. Completeness [Sha50]. Complex [GF59]. Components [GW56, Kin54, Kit55, Kit56, Zel58]. Composed [Bak58]. Computation [BHP53a, BHP53b, Gri58, Jas56, WF57, WF58]. Computations [Kai59b]. Computer [Fry56, SE57]. Computers [Loc59]. Computing [Cow58, Cow59, Lor55b]. Comulative [Mac59c]. Concave [Hil54, Hil55b]. Concentration [Ade55]. Concept [BPR55]. Conceptions [Vin53]. Concepts [EBP55]. Condition [GT56]. Conditions [And53, Goo53b, Kuz50b]. Confession [Wat56]. Confidence [Bue57, Cro55, Cro56, Dun59b, Dur54, Jon58b, Kat53, Noe57b, Pro53, TK59, Wal50b, Woo52d]. Congress [Ano51h]. Consideration [War56]. Considerations [SH52]. Consored [Coh54]. Construction [Kap59b, Met50b, Rot58]. Consulting [Kim57]. Consumer [Gar56a]. Consumption [Fox51a]. Contigency [Bla58a]. Contingency [Bla58b, Cra53b]. Continuous [Ans58, Ans59b, Sav59a]. Contrasts [Hol57]. Contributions [Wal58a]. Control [AL50, AL52, Dun56a, Dun55, JM57, RGM58a, RGM58b, Ste59a, Wei53, Wei54, Wei56, Zar54]. Controlled [Gea53, GG56a, GK50, Ham50, Sch58, GG56b, GG67]. Coordinating [Ric54]. Copley [Ano52f]. Corporate [Cru56]. Corporation [RBJ51]. Correction [Adl51, Bea59, Ber61, CM50a, Cro52, Wal50a]. Corrections [Eis56]. Correlated [Sea59]. Correlation [Cow52, Foo53, Foo58, Lor55b, OS59, Sim54, Spu51, Tat55, Wei58, Wes52]. Corrigenda [Ade59, And55a, Ans59b, BAR56, BCMM59, Bir66, Bla58a, Bra58b, Che56, Coa55b, Coe55a, Cow59, Cro56, DH63, GG56b, GG67, GK57, GK58, GH63, Gra56, Hil55b, JR55a, Kim55a, KH63, Kit56, Lan59a, LR55a, Mac59c, Mad59a, Mic60, Moo56, Mor60, NY59a, Noe56b, PJ57a, Raf56a, Ram59, Rid55a, Rid56, Rid59, RGM58a, Sav58a, Sto55a, WF58, Wei56, WK56, Wil56b, Yer56, kem56]. Cost [RGM58a, RGM58b, dlGHN55]. Costs [Bla54, Wil55a]. Counterbalancing [Bra58b, Bra58a]. Counties [Bro55b, CS56]. Countries [Abr58, Osh57]. Counts [EBS55]. County [Jeb52, Mad50, Sch52b]. Course [Vot52]. Crime [Bea59]. Criterion [Ask57, Kin54, KW52, KW53]. Critical [JN51, Lee52, Mos53]. Cross [GK54, GK57, GK58, GK59, Hyr53a, Hyr53b, LB58]. Cross-Ranges [Hyr53a, Hyr53b]. Cross-Section [LB58]. Crude [Sto55b]. Cumulative [Mac59b, Man57]. Current [Dun56a, HHNS55]. Curtailed [Gar58]. Curve [BG52, Ber55a, CM50a, CM50b, Gar56a, Kar58c, Mil55, SM50, Sil57, Bro52a]. Curves [Dun57b]. Cut [Ask57, Sch52b, Spu51]. Cycles [Eis58, Fab54]. Cyclical [Hic58, Sav54b].

D [Bro52a]. D.C [Alc54, Ano54g]. Data [And55b, Bel54, Cla59, Dav52, Dea52, Dun56b, EBP55, Fis53, GS59, Joh50b, Kin53, KM51, Lom54, Lor55a, Man57, Mil53, SM58, Smi54, Tat55, VK52, Zar54]. December [Alc53, Alc54, Alc56, Alc59, Ano51h, Ano52f, Ano53h, Ano54g]. Decision [Sav51b, Sav54a]. Decisions [Spr50, Ste59b, Tul59]. Defective [Chu50, SM58]. Defining [Hol57]. Definitions [Wei58]. Deflation [Kap59b]. Degree [Fre50c]. Degrees [DR58]. Delimitation [Vin53]. Delivered [Alc53, Alc54, Alc55, Alc56, Alc57, Alc58b, Alc59]. Demand [Dal58, Gri59b, Hou54b]. Dependent [Dun59b, Moo57a]. Deposits [Gar53]. Derivation [Wei50]. Design [BD59, Bra58b, Bra58a, DH50, Dem56, Dun56a, Dun58, MMN53, NW59, WK57, Hou53]. Designing [Gol53, Moo51]. Designs [BS52, BF53, Jeb52, Mei54, Wal52b, Zel57]. Determination [Moo51]. Detroit [Alc57]. Development [Lie58b, Spo51, Yat51]. Developments [Mos55]. Deviate [HGCZ55]. Deviation [Eva51b, GS50, Noe55, Noe56b, Slo58]. Deviations [Coc55, Kar58c]. Diagrammatic [Li55]. Difference [Kit55, Kit56, Moo55, Moo56, Stu52]. Difference-Sign [Stu52]. Differences [KH59, KH63, Rao59]. Differential [Yer55, Yer56]. Differentials [Kis52, Mor51a]. Diffusion [Hic58]. Digits [Ano52k, Ano53l, Ano53m, Ano53n, Ano53o, Ano54j, Ano54k, Ano54l, Ano54m]. Disarray [Gri58]. Discrete [Ait55, Rid51b]. Discriminant [Joh50b]. Discrimination [Sto54, Sto55a]. Discussion [GK59]. Disease [Elv58b, MG55]. Dispersion [Coh55a]. Distinction [Dem53b]. Distributed [Coh55a, Dun59b]. Distribution [Ade58, Ade59, Ait55, And57, BH54, Bir52, Bur53, Fed59, GGD53, Gum54, Hal52, Mac59c, Mac59b, Mil51, Mos53, Mur55b, Pro53, Raf51, Rid51a, Rid51b, Rid53a, Rid53b, Rid55b, Rid56, Rid59, Sar55, SO59, Sto54, Sto55a, Stu54, TK59, Wag58, Wal56c]. Distribution-Free [Sto55a, Stu54, Sto54]. Distributions [Bab55, Bak58, Coh51, Coh55b, KR56, Rei56, Rid53c, Rid55a, Rid55c, SG57]. Divorce [Car53]. Do [BDL^{+}54]. Doolittle [Foo58]. Dosage [CM50a, CM50b]. Double [Dun58, HV55]. Double-Limit [Dun58]. Doubly [SG57]. Down [BHR53]. Drawn [Ste59b]. Due [BS50]. Dummy [Sui57]. Duplication [Ros53, Ros54]. Dwelling [Rei59]. Dwellings [KH58].

Each [Key57]. Earning [dJ59]. Econometric [Cro59]. Economic [Arr58c, Ber53b, BHP53a, BHP53b, Cla59, Dem53a, Dun56a, Fis52, Har56b, Hef51, Leb54, McL51, Rus57, Spr55, Vin53]. Economically [Bre56]. Economics [Cop55, Har56b]. Economy [BK54]. Edgeworth [Ber57]. Editing [Nor55a]. Edmund [Met50a]. Education [Wil51]. Educational [Zak56]. Effect [DR58, Fer56, Ros53, Ros54]. Effective [Wei55]. Effectiveness [AL50, AL52, Fis53, Noe57b]. Effects [Bra58b, Bra58a, Bro57b, Dun57b, Ste59b]. Efficiencies [Stu54, Stu56, Wal56c]. Efficiency [HV55, Ros53, Ros54, Rot52]. Elasticity [DR58]. Electronic [BHP53a, BHP53b, SE57]. Elementary [Fis52]. Elements [Mil55, Sim53a]. Empiric [Bak58]. Empirical [Dal58, Whi54].

Employed [EBP55]. Employing [Jeb52, Jon56b]. Employment [Bel50, JN51]. Engineering [Bur54, Tuk52]. Engineers [TW52]. England [Ern58]. Enumeration [MMN53]. Enumerative [Dem53b]. Epidemiology [MG55]. Equal [Dem56, Gar56a]. Equations [Bab55, HD53, Sui57]. Equipment [Goo53b]. Equivalence [Gay56]. Errata [Ano55c, Ano55d, BHP53b, DW53a, GS54a, Hyr53a, KW53, LLT54, Rid53b, RK53, Ros54, Sav54a]. Erratum [Ano50f, Ano50g, AL52, Moo51]. Error [Coh55a, Dem53a, KB53, Mad59a, Mad59b, Mor59, Nor55a, Pri55, Rei56, Mor60]. Errors [CL54b, DW53a, DW53b, HHMM51, Kai59b, Kim57, KL54, Lor59, MM50]. Eshref [Bue56]. Estimate [Ber55a]. Estimates [Abr58, And57, Ber55b, Bog50, Coh59a, Eva51b, Gay56, Hil54, Hil55b, JN51, Key57, Mad50, Mor59, Mor60, NY59a, NY59b, Rei56, SG57, Wei50, Zel58]. Estimating [Ber53a, Bro55b, Coh50, Coh51, Coh59b, CS56, Elk53, GS50, KL54, PJ57a, PJ57b, Sch52b, Sha50, She52, SH52, Sil57, Smi50a, Wei55]. Estimation [Bro55a, Coh54, Coh55a, DH50, Elv58b, Hal52, Jeb52, KM58, Lor55a, Nic57, Pri55, Qua58, Raj53, Rob57b, Sar55, Ste55b]. Estimator [Jas51b, Rid57]. Estimators [GH58, GW56, Mic59b, Mic60, Raj56, GH63]. Evaluation [BDL^{+}54, Dun56b, Lee52, Slo58]. Evidence [Cut55, Dem54]. Evolving [Dun52]. Examination [Pri55]. Example [Gil50, Lom54]. Exercises [Che56]. Expectations [Fer53, Fer58a, dJ59]. Expected [Coc55, GS54a, GS54b]. Expenditures [Coh59b, DR58]. Experience [BDL^{+}54, Kar58b]. Experiment [Hic58]. Experimental [And55b, BF53, Kem55, Old54, Tuk54, Wal52b, dlGHN55]. Experimentals [kem56]. Experimentation [Dan53, Dyk58, You51]. Experiments [BAR56, BGR56, Bur54, DH50, GG56a, GG56b, GG67, Moo51]. Exponential [Bir54b, Bir66, SG57, SM50]. Extended [Fed59, JR55a, JR55b]. Extensive [Bog50]. Extraneous [Dur53]. Extrapolation [DH50, Tuk52, dlG56]. Extreme [Kim55a, Kim55b]. Extremes [Hyr53a, Hyr53b].

F [Bro52a]. Factor [Gra55, Gra56, Gre52b]. Factorial [BAR56, BGR56, Dyk58]. Factors [Fab52, GWS50, Mye54b]. Failure [Dav52, Lom54]. Failures [Lom54]. Faith [Wat56]. Family [TL55]. Farm [Dal58, JN51]. Fatigue [FG54]. Federal [Ade55, Coh59b, Man53]. Females [Bra54b]. Ferrying [Dee54]. Fertility [Kis52, Lah56]. Fertilizer [Gri59b]. Figures [Coa55b, Coa55c]. Finding [And53, Hol57]. Finite [Bir52, Chu50, HT52, Mic59b, Mic60]. Finney [Bro52a]. Firms [Ade58, Ade59, Fer58a]. First [HR58, Vot52]. Fiscal [Cru56]. Fiscal-Year [Cru56]. Fisher [Hot51, Mat51]. Fisherian [You51]. Fit [AD54, Mas51, Wil50]. Fitting [Ask57, Kar58c, Mad59a, Mad59b, Man57, Sch58, SM50]. Fixed [Dun57b, Gol53, WK55, WK56]. Flow [Smi54]. Fluctuations [Sav54b]. Following [BG52]. Food [Bur58]. Force [Bel50, EBP55]. Forecast [Slo58].

Forecasting [HS52]. Forecasts [Dor50b, Gri51, Haj55]. Formation [Jas51a]. Formulas [Wil59]. Foundry [Sav54b]. Fractional [Moo57b]. Free [Sto55a, Stu54, Sto54]. Frequency [TL55]. Front [Ano50h, Ano50i, Ano50j, Ano50k, Ano51d, Ano51e, Ano51f, Ano51g, Ano52b, Ano52c, Ano52d, Ano52e, Ano53e, Ano53f, Ano53g, Ano54c, Ano54d, Ano54e, Ano54f, Ano55e, Ano55f, Ano55g, Ano55h, Ano56d, Ano56e, Ano56f, Ano56g, Ano57e, Ano57f, Ano57g, Ano57h, Ano58b, Ano59d, Ano59e, Ano59f, Ano59g]. Frontiers [Cox57b]. Function [Ber53a, Ber55b, Cla53b, Coc55, Cro55, Cro56]. Functional [Bur53, Gea53]. Functions [Hil54, Hil55b, Hoo58, Sim51b, TL55]. Further [GK59].

G [JR55a]. Games [Gil50, Gre53b]. Gamma [Wal56c]. General [Yer55, Yer56]. Generality [Cro55, Cro56]. Generalization [HT52]. Generalizations [God55]. Get [PS50]. Giant [Met50a]. Gompertz [SM50]. Goodness [AD54, Mas51, Wil50]. Goods [PJ57a, PJ57b]. Government [Coh59b]. Graduation [Lah56]. Gram [Ber57]. Graphic [Ask57, Ask59, Foo53, Gri58, Min57a, WF57, WF58]. Graphical [BPR55]. Great [Mos55]. Gross [Kai59b, Zel58]. Group [Bal51, Jon59, Kim53]. Grouping [Cox57a, Fis58, Kin54]. Growth [Ger57, NY59a, NY59b, Nut54]. Guide [Lan59a, Lan59b].

Harvest [Ste55b]. Having [Ait55]. Health [Dun52, NW59]. Height [Kar58b]. Hicks [Fis52]. Higher [Jas56]. Historical [Kru57]. History [MMN53]. Hollerith [HD53]. Homes [KL54, PS50]. Homogeneity [Bak58, Fis58]. Horatio [Mal50]. Hospital [Bel54]. Hotel [Ano50l, Ano51h, Ano52f, Ano54g]. House [Ano53h, Bla54, Wil54, Wil56b]. Household [NW59, PJ57a, PJ57b]. Households [Lie58b]. Housing [Wil55a]. Human [Mos55]. Hypotheses [Ban50, Sav57b, dlG56]. Hypothesis [Mar51, Wal52b].

Ignorance [Fer56]. II [Fit58, GK59, Kar58b]. III [Coh50, Raj53, Sea58]. Illinois [Alc59, Ano51h, Ano53h]. Illustrations [Rot52]. Immediate [Bra58b, Bra58a]. Immigration [Hut58]. Impact [Hot51]. Improving [Car53]. Inadmissible [Jon58b]. Incidence [GWS50]. Income [Bur58, Bur53, Cru56, DR58, Fis53, Ger57, Jas51b, Mil51, Mil53, Mil55, Osh57, RBJ51, Sel55, Wei50, Zel58]. Income-Food [Bur58]. Income-Maintenance [Fis53]. Incomplete [BS52, Cla53b, KM58, Lor55a, Nic57, Zel57]. Increase [Rei59]. Increases [Coe55a, Coe55b]. Increasing [Wei55]. Independence [Sav57b]. Independent [Bir54a, Edg56]. Index [Bla54, CR56, Ern58, Jas51a, Leo50, Rot58]. Indexes [Hic58, Mud55, Sim51b]. Individual [Fer58a]. Industrial [Sim53a, Sog54, Tou55]. Industry [Fer58a, Fry56]. Inequality [God55]. Inference [Kem55, kem56]. Inferences [Ste59b]. Influence [HM58, Yat51].

Information [Ano50q, Ano51q, Ano52l, Ano53p, Ano54s, Ano55o, Ano56q, Ano57o, Ano58k, Ano59p, Dur53]. Ingenieure [TW52]. Inhalation [Ham59]. Insensitive [Wal51]. Inspection [Ans58, Ans59b, Gol53, Lie52, LR55a, LR55b, Mos56b, Old51b]. Instead [Noe55, Noe56b]. Instruments [Smi50a]. Insurance [Man53]. Integrated [Ber55a]. Inter [Gri59b, Rot52]. Inter-Spatial [Rot52]. Inter-State [Gri59b]. Intercensal [CS56]. Intercorrelation [Wil59]. Interest [GS59, Sel55, Wic53]. Interference [Ham50]. Internal [Pri55]. International [Dun52, Jas50, Zak56]. Interpolation [Tuk52]. Interpretation [Sim54]. Interpretations [Bla58a, Bla58b]. Interpreting [Mye54b]. Intervals [Bue57, Cro55, Cro56, Dun59b, Kat53, Noe57b, Pro53, TK59, Wal50b, Woo52d]. Interview [NW59]. Interviewer [HM58]. Introduction [Sav53]. Inventory [Eis58, LL53, LLT54]. Inversion [Dur56, DW53a, DW53b, GS59]. Investigating [Jon56b]. Investigation [Bak58, Hoo58]. Investment [Abr58]. Iran [Win59]. Isopleth [SM55].

J [Bro52a, TL55]. J-Shaped [TL55]. James [Dub50]. Joint [Dur54].

Karl [Sto58b, Wal58a]. Karmel [Whe50]. Kendall [Adl57]. Kind [Kim57]. Kinsey [CMT55, Bra54b, CMT53]. Known [Wal50a]. Kolmogorov [Bir52, Mas51, Mil56a]. Kruskal [Wal59a].

Labor [Bel50, EBP55, Sie53, Zob50]. Labor-Force [Bel50]. Lady [Gri59a]. Large [Mar51, Wal50b, Wal51]. Large-Sample [Mar51]. Larger [Mar51]. Latent [Gre52b]. Latin [Bra58b, Bra58a, WK57]. Lattice [Mei54]. Leading [Fit57, Fit58]. Least [Ask57, Kar58c, Kim53]. Legal [Dem54]. Length [BS58, Wei55]. Less [Wei50]. Letter [Ano53d]. Level [Dal58]. Levels [GG56a, GG56b, GG67, Mac59c, Mac59b]. Life [Bar57, BS58, ES53, Eps55, FG54, Goo53b, PJ57a, PJ57b, Wal56c]. Life-Length [BS58]. Likelihood [And57, Ber55b, Coh51, Coh55a, CM50a, CM50b, KR56]. Limit [Dun58, Sea59]. Limits [Jon58b, Mit57a, TM59, Wei54]. Linder [TW52]. Line [Ask57, Man57]. Linear [Bab55, Coc55, DH50, Gea53, Kar58c, Moo57a, Qua58, Sar55, SG57, Wag58, Wag59b, Wel56]. Lines [Mad59a, Mad59b, Sch58]. Lists [DG59]. Literacy [Zar54]. Literary [Win59]. Literature [Lan59a, Lan59b]. Livestock [Nor52b]. Log [Mos53]. Log-Normal [Mos53]. Logarithmic [Coh51]. Logarithmic-Normal [Coh51]. Logistic [Ber53a, Ber55b, Sil57]. Long [BDL^{+}54, Ger57]. Loss [Dee54]. Lost [BDL^{+}54]. Lot [GT56, Mos56b]. Lotka [Dub50]. Lots [Chu50]. Low [BH54]. Lower [Ber61, Wil59]. Lung [Ber58, Cut55].

Machine [Jas56]. Machines [Met50a]. Magnitudes [Nor52b]. Mail [CL54b]. Mailed [EB56]. Main [Ano53c]. Maintain [Dun56a]. Maintenance [Fis53]. Maker [Hir56]. Making [Bog50, Hir56, Wei50]. Man [Ree52]. Manufacturer [Eis58]. Manufacturing [Ade55, Ern58]. Many [Jon59]. Mapping [SM55]. Maps [Cho59]. Mark [AC55]. Mark-Sense [AC55]. Market [War56]. Marketing [BF53]. Markov [MG55]. Marriage [Car53]. Mass [Ait55]. Massachusetts [Ano52f]. Matching [DG59]. Materials [BS58]. Mathematical [Coo51a, Foo53]. Matrix [Dur56, DW53a, DW53b, GS59]. Matter [Ano50a, Ano50b, Ano50c, Ano50d, Ano50h, Ano50i, Ano50j, Ano50k, Ano51b, Ano51c, Ano51d, Ano51e, Ano51f, Ano51g, Ano52a, Ano52b, Ano52c, Ano52d, Ano52e, Ano53a, Ano53b, Ano53e, Ano53f, Ano53g, Ano54a, Ano54b, Ano54c, Ano54d, Ano54e, Ano54f, Ano55a, Ano55b, Ano55e, Ano55f, Ano55g, Ano55h, Ano56a, Ano56b, Ano56c, Ano56d, Ano56e, Ano56f, Ano56g, Ano57a, Ano57b, Ano57c, Ano57d, Ano57e, Ano57f, Ano57g, Ano57h, Ano58a, Ano58b, Ano59a, Ano59b, Ano59c, Ano59d, Ano59e, Ano59f, Ano59g]. Maximum [And57, Ber55b, Coh51, Coh55a, CM50a, CM50b, Fis58, HGCZ55, Mur55b, Rid55b, Rid56, Rid59]. Mean [Bla58a, Bla58b, Jon56b, Mit57a, Moo55, Moo56, Rao59, Wei53, Wei56, Wil59]. Means [Cla57a, Dun59b, JR55a, JR55b, Jon56b, OS59, Wei54]. Measure [Spu51]. Measurement [Jar58, Leb54, Lor59, Mad51, McL51, Min57a, Mud55, Sim58, Zob50]. Measurements [Dal58, EBP55]. Measures [GK54, GK57, GK58, GK59, Hef51, Kru58b, Noe57b, Tho52, Woo52d]. Measuring [Fer53, Fis53, Ger57, Goo53b, Moo58, Nor55a, Nor52b, Smi50a, TJ57, War56]. Meat [Nor52b]. Mechanism [Dem53a]. Mechanisms [Dun52]. Medians [Woo52d]. Medical [Dea52]. Mediziner [TW52]. Meet [Cla57a]. Meeting [Alc53, Alc54, Alc55, Alc56, Alc57, Alc58b, Alc59, Ano50l, Ano51h, Ano52f, Ano53h, Ano54g, Ano50l, Ano51h]. Memory [Gra55, Gra56]. Mental [Lor59, MG55]. Merger [Nut54]. Method [Ask57, Ber53a, BHR53, CS56, Dur56, Elv58b, Foo53, Hol57, Min57a, SH52, SM58, Slo58, Spr50]. Methoden [TW52]. Methodological [Kap59b]. Methodology [Woo51a, Woo51b, Woo51c, Woo52a, Woo52b, Woo52c]. Methods [Ano50g, Ask59, Bir54b, Bir66, Goo53b, Jar58, Kru54, Mat51, PJ57a, PJ57b, Sch52b, SE57, TW52, Whi54, Wil54, Wil56b, Yat51, You51, Zar55]. Metropolitan [Coe55a, Coe55b, Dun56b]. Midrange [Rid57]. Migration [Pri55, SH52]. Milo [Mal50]. Minerals [Leo50]. Minimum [Ber55a, Ber55b, DH59, FG54, Spr50, TM59, dlGHN55, DH63]. Minutes [Ano50l, Ano51h]. Missing [And57, Edg56]. Mitchell [Ber53b]. Mixed [WK55, WK56]. Mobilization [McL51]. Mode [Jon53]. Model [BK54, BS58, Cro59, Fis52, Sav59a]. Models [Coo51a, Leb54, Tat55, WK55, WK56]. Modification [Adl57]. Modified [Foo58, SM50]. Moments [Jas56]. Monetary [Eis58]. Moneyflows [Men55]. Monotone [KR56]. Montreal [Alc55]. Morbidity [Bel54, Dea52]. Mortality [Mor51a, Mye54b, Sob59, Wal50b, Yer55, Yer56]. Most [DH50]. Movements [Mud55]. Mr. [Whe50]. Multi [KH59, KH63]. Multi-Stage [KH59, KH63]. Multiple [Cow52, Dun55, Dur54, Edg56, Foo53, Foo58, Kim53, Lor55b, RK52, RK53, Ste59a]. Multiple-Partial [Cow52]. Multistation [Lie52]. Multivariate [And57, CF58, JM57, Nic57, Rob57b].

N [Tuk52]. National [Car53, Jas51b, NW59, Osh57, Sim58, Zel58]. Natural [TW52]. Naturwissenschafter [TW52]. Negative [Rid55a, Rid55c]. Neighbor [Mar55]. Neighborhood [BH54]. Neighboring [Cra53b]. Net [Pri55, SH52]. Nineteenth [Fit57, Fit58]. Nomograph [Lor55b]. Non [BS50, Eps55, Gea53, LB58, Wal51, WK57]. Non-Additives [WK57]. Non-Availability [BS50]. Non-Business [LB58]. Non-Linear [Gea53]. Non-Parametric [Eps55]. Non-Randomness [Wal51]. Noncoverage [KH58]. Nonparametric [Ano53c, Ano53d, KM58, Sav53, Sav57a, Sav58a, Wal52b, Wal56c]. Nonresponse [Bro57b, Dem53a]. Nonwhite [Coe55a, Coe55b]. Normal [And57, Bak58, Ber55a, CL54a, Coh51, Coh55b, Dav56, Eps55, GGD53, Gum54, Hal52, HGCZ55, HD53, Mit57a, Moo57a, Mos53, Pro53, Rei56, Stu54, Stu56, TK59, Wal52a]. Normality [Moo57b]. Normally [Dun59b]. Normit [Ber55a]. Northern [Mar55]. Not-at-Homes [PS50]. Note [Ber61, Bro57b, Cox57a, Dur56, Dur53, GW56, Kar50, PS50, Rao59, Wei50, Whe50, dJ59]. Notes [Ano58c, Ano58d, Ano58e, Ano58f, Ano59h, Ano59i, Ano59j, Ano59k, Hut58, Kap59b, Kru57]. Notices [Ano50e, R.50]. Number [Bro55a, Goo52, Goo54, Jas51a, Kat53, Wil50]. Numbers [Coc55, Jon59]. Numerical [Bir52, Tei55]. Nutshell [Slo57].

Obeying [Qua58]. Objective [Cop55]. Observation [BDL^{+}54]. Observations [And57, Ber58, Coh59a, Edg56, KM58]. Observed [BDL^{+}54, Coc55]. Old [Man53]. Old-Age [Man53]. One [Dur53, Elk53, Gea53, Kin54, KW52, KW53, Mar51, Sch58, Ste51b]. One-Criterion [Kin54, KW52, KW53]. Operating [And53, Dun57b, Gar56a, Wal52a]. Operation [Dun58]. Operational [Goo53b]. Operations [Dee54, Gil50, Jon55a]. Opportunities [Bel50]. Optimal [TK59]. Optimum [BCMM56, BCMM59, Bre56, Bro55a, Eva51a, Kin54, Som57]. Order [SG57]. Ordinal [Kru58b]. Ordinates [Hil54, Hil55b]. Organizations [Jas50]. Orientation [War56]. Origin [Ait55, TJ57]. Orthogonal [Kim53]. Other [Ber61, Elk53, Mar51, Sav57b, Woo52d]. Outgoing [SO59]. Outlook [Leo58]. Output [Ans58, Ans59b, Sim58].

P. [Kar50]. Paasche [Jas51a]. Paired [Dyk58, Sch52a]. Palmer [Ano53h]. Panel [Sob59]. Paper [CL54a]. Papers [Alc53, Alc54, Alc55, Alc56, Alc57, Alc58b, Alc59, Ano53d]. Parallel [Goo58a]. Parameter [Coh54, Coh55a, NY59a, NY59b]. Parameters [Bue57, Coh50, Coh51, Coh59a, Lor55a, Nic57, Qua58, Raj53, Sar55, SG57, Sea58, Wei55]. Parametric [Eps55]. Parity [Kar50]. Partial [Cow52, Foo58]. Partially [BS52]. Participation [Bel50]. Particular [Ber55a, Jon59]. Partition [KB53]. Partnership [Hir56]. Patients [BG52]. Payments [Fab54]. Pearson [Coh50, Sto58b, Wal58a]. Percentage [BH54, Cla53b, Fed59, HGCZ55, Mil56a, SO59]. Percentiles [Wal56c]. Performance [And55b]. Personal [Sel55, Zel58]. Persons [BDL^{+}54, Wei50]. Philippine [Lie58b]. Phillips [Jas51a]. Photographic [JM57]. Physical [Leo50, Tou55]. Physicians [TW52]. Pitfalls [Dor50b]. Plan [Dun58, Mos56b, Sav59a, Sea59]. Planning [McL51, Ree52]. Plans [Gol53, LR55a, LR55b]. Plaza [Ano52f]. Plot [Mos56b]. Point [Hil54, Hil55b, Raf56b, Raf56a]. Points [Ask57, BH54, Cla53b, Dun57b, Fed59, HGCZ55, Mil56a, SO59]. Poisson [Bir66, Bir54b, Coh54, GS54a, GS54b, Rid53c]. Policy [Arr58c, Eis58, Hir56, Sim53a]. Polio [Bro55c]. Pollock [Mal50]. Polykays [Rob57b]. Polynomials [Kim53]. Population [Bog50, Bro55b, Coa55b, Coa55c, Coe55a, Coe55b, CS56, Dor50b, Dun56b, HS52, Haj55, HHNS55, Kat53, Mic59b, Mic60, Mit57a, NCD50, Not50, Rid51a, Rid51b, Rid53a, Rid53b, Rid57, Sch52b, Shr51, Som57, Sto54, Sto55a, Wal52a, Zar55]. Populations [Bak58, Bir54b, Coh50, Dav56, Moo55, Moo56, Mos55, Raj53, Sea58, Whi54, Bir66]. Position [Woo52d]. Positive [Ait55, GS54a, GS54b]. Positivity [Ber57]. Possessing [KR56]. Possible [Ste59b]. Post [MMN53]. Post-Enumeration [MMN53]. Power [Stu52]. Powers [Moo57b]. pp [CMT55]. Practical [Goo54, Kim55a, Kim55b, Zak56]. Precise [Ber53a, DH50]. Precision [GH58, GH63, Smi50a]. Predecessor [Jas51a]. Predicting [And55b, Vot52]. Prediction [Mad51]. Prepaid [Dea52]. Present [Coc54b]. Presentation [Dem54]. President [Spr55, Har56b, Rus57]. Presidential [Hoa59]. Price [Hef51, Mud55, Rot58]. Prices [Bla54, Fox51a, Wil55a]. Primary [Jeb52, Mad50]. Principles [CMT54, VK52]. Probabilistic [Bla58a, Bla58b]. Probabilities [Cho59, Dem56, Gre53b, Key51, Noe57b, Raj56, WW53]. Probability [Ait55, Ban50, BF53, CL54a, DN58, Dem53a, GK50, Mac59c, Mac59b, Tou55]. Probable [Mor59, Mor60]. Probit [Bro52a]. Problem [And54, And55a, Bar57, CF58, DG59, Kim53, LL53, LLT54]. Problems [And55b, Bal51, Coh59b, Dal58, Ger57, Lor59, NW59, Pre55, Ric54, Rot52, SM55, Tuk54, Wag58, Wes52, CMT53]. Procedure [Cow58, Cow59, Dun55, EB56]. Procedures [Jon55a, Kin53, Pra59, SM50, dlGHN55]. Proceedings [Ano50l, Ano51h, Ano52f, Ano53h, Ano54g]. Process [And55b, Dun56a, Ham50]. Processes [Bir54b, Bir66, MG55]. Processing [JM57, VK52]. Producer [Gar56a]. Product [Bue57, Gay56, Rid53a, Rid53b, Rid55b, Rid56, Rid59, She52, Zel58]. Production [And55b, CR56, Dun52, Ern58, Fox51a, GT56, Leo50, Nor52b, Old51b, Sav59a, Sim51b, War56]. Productivity [Sie53, Zob50]. Products [Dal58, Li55]. Program [Lie58b]. Programming [Bab55, Loc59, Wag58, Wag59b]. Programs [Fis53]. Projecting [Whi54].

Projections [Dor50b, HS52]. Properties [Ask59, Jon56b, Moo55, Moo56]. Proportion [Chu50]. Proportional [Key51]. Proportions [Elk53]. Prospects [Haj55, Jas51b]. Provisional [Zel58]. Psychological [Ano50f, Coo51a]. Public [Fis53, Wic53]. Publication [Ste59b, Tul59]. Publications [Ano50m, Ano50n, Ano50o, Ano50p, Ano51i, Ano51j, Ano51k, Ano51l, Ano52g, Ano52h, Ano52i, Ano52j, Ano53i, Ano53j, Ano53k, Ano54h, Ano54i, Ano55i, Ano55j, Ano55k, Ano56h, Ano56i, Ano56j, Ano56k, Ano57i, Ano57j, Ano57k, Ano57l, Ano58g, Ano58h, Ano58i, Ano58j, Ano59l, Ano59m, Ano59n, Ano59o]. Puerto [Rot52]. Punch [BAR56, BGR56]. Purchases [dJ59]. Purpose [Wei55].

Quadratic [dlGHN55, dlG56]. Qualitative [Joh50b]. Quality [AL50, AL52, JM57, SO59, Wei53, Wei56]. Quantal [Ber53a]. Quantification [Joh50b]. Quantity [Mud55]. Quarterly [BK54]. Questionnaire [CL54b, Met50b]. Questionnaires [EB56, Nor55a]. Quotient [Mur55b, Rid51a].

R. [Hot51, Mat51]. Radial [Coh55a]. Random [Ait55, Ano52k, Ano53l, Ano53m, Ano53n, Ano53o, Ano54j, Ano54k, Ano54l, Ano54m, Jon56b, Jon59, Mar51, RGM58a, RGM58b, She52, WK55, WK56]. Randomization [CF58, Hoo58, Kem55, kem56]. Randomized [KB53]. Randomness [Sav57b, Stu54, Stu56, Wal51]. Range [BH54, Mit57a, Noe55, Noe56b, Rid51b]. Ranges [Hyr53a, Hyr53b, OS59, Rid51a, Rid53a, Rid53b, Wei54]. Rank [Pra59, Wal59b]. Ranking [Dav56, Jar58, Sea58]. Rankings [Adl57]. Ranks [KW52, KW53]. Rapid [BK57]. Rates [Kar50, Kit55, Kit56, Lah56, Sto55b, Wal50b]. Ratio [DN58, Elk53, GH58, GH63, KR56, Mic59b, Mic60, Noe57b, Rei56, Rob57b, Wal50a]. Ratio-Type [GH58, GH63, Rob57b]. Rational [TJ57]. Ratios [KH59, KH63]. Received [Ano50m, Ano50n, Ano50o, Ano50p, Ano51i, Ano51j, Ano51k, Ano51l, Ano52g, Ano52h, Ano52i, Ano52j, Ano53i, Ano53j, Ano53k, Ano54h, Ano54i, Ano55i, Ano55j, Ano55k, Ano56h, Ano56i, Ano56j, Ano56k, Ano57i, Ano57j, Ano57k, Ano57l, Ano58g, Ano58h, Ano58i, Ano58j, Ano59l, Ano59m, Ano59n, Ano59o]. Recessions [Moo58]. Records [Bel54, Man53]. Recreation [Cla59]. Rectangular [Mur55b, Rid51a, Rid51b, Rid53a, Rid53b, Rid55b, Rid56, Rid59]. Rectifying [Ans58, Ans59b]. Redesign [HHNS55]. Redistribution [Shr51]. Reference [Ber55a]. References [GK59, Lan59a, Lan59b]. Refinement [AC55]. Regimes [Qua58]. Regions [Ber57, Dur54]. Registration [Sha50]. Regression [And54, And55a, Cow58, Cow59, Cro55, Cro56, Dur54, Dur53, Edg56, Eis56, Foo58, Mic59b, Mic60, Qua58, Stu56, Sui57, Wag59b, dlGHN55]. Regressions [Ber50]. Rejection [Kin53]. Related [Ano53c, Ano53d, Dal58, Lan59a, Lan59b, Sav53, Sav58a, dlG56]. Relating [And55b]. Relation [Fer58a, Gre52b, Ham59]. Relations [Fox51a]. Relationship [Bla54, Gea53, Wil55a, dJ59]. Relationships [Bur58]. Relative [Raj58, Rei56, Stu54]. Relatively [Ber53a]. Reliability [Fer55b]. Religious [Lan59a, Lan59b]. Remarks [Pra59]. Rent [Rei59]. Replacement [HT52, Kat53, Raj56]. Replacing [Wes52]. Replication [Dem56]. Replications [Mei54]. Report [Ade55, Bra54b, KHS^{+}55, Leo58, Rus57, Spr55, CMT53, CMT55]. Reporting [Cru56, Mye54a]. Reports [Ber58, Har56b]. Representation [BPR55, Li55]. Representatives [Wil54, Wil56b]. Reprints [Woo51a, Woo51b, Woo51c, Woo52a, Woo52b, Woo52c]. Reproduction [Kar50]. Reproductivity [Sto55b]. Required [GS50]. Requirements [Cla57a, GF59]. Research [Ano50g, AC55, Bal51, BF53, Cla59, Dun56b, Fol53, Kuz50b, MM50, Mat51, Tei55, Yat51]. Residual [SH52]. Resources [NCD50]. Respondent [Fer56]. Response [Ber53a, BD59, Bro52a, CL54b, CM50a, CM50b, Dem53a, HHMM51, KL54, MM50]. Responses [Fer55b]. Restriction [Coh55b]. Resultant [Dem53a]. Results [Dem54, Fer56, HM58, KM51, Ros53, Ros54]. Retail [Dal58]. Retirement [Mye54b]. Returns [RBJ51]. Revenue [Sel55]. Review [Ada54, Adl55, Alc58a, Ale57, Alt58, And50, And56, And55c, Ang57, Ans56, Ans59a, Ara52, Ark56, Arn52, Arr54a, Arr54b, Arr58a, Arr58b, Atk55, B.56, B.57a, B.57b, B.58, Bac53, Bac58, Bah57, Bai58, BO53, Bak56, Bal56, Ban58, Bar58, Bea57, Bea54, Bea50, Bel55, Ben57, Ben50a, Ben50b, Ben55, Ber56, Bey52, Bir55, Bir56, Bla51, Bla58c, Bla56, Ble58, Blu58a, Blu54, Bod57, Bod59, Boh55, Bor57, Bos58, Bou59a, Bou59b, Bow50a, Bow50b, Bow56, Box55, Bra56a, Bra51a, Bra51b, Bra54a, Bra56b, Bra58c, Bra59a, Bra59b, Bra55, Bre58, Bri50, Bri59, Bri56, Bro58a, Bro53a, Bro58b, Bro59, Bro53b, Bro56a, Bro56b, Bro52b, Bro52a, Bro57a, Bro58c, Bru55, Bry57, Buc54, Bud56, Bue56, Bur55, Bus56a, Bus56b, BC56, Cam57, Can50, Cha51, Cha52, Che55a]. Review [Che55b, Chi54, Chi59, Chi53, Chu57, Chu51, Cla50, Cla52, Cla53a, Cla57b, Coa55a, Coc52, Coc54a, Coc56, Coc57, Coh56, Coh57, Col57, Col56, Com59, Con52, Con56, Coo51b, Coo58, Cor50, Cor51, Cor55, Cor58, Cov51, Cox59, Cra50, Cra53a, Cra53c, Cro50, Cur54, Cut55, Dal54, Dar51, Dar55, Dav59a, Dav59b, Day51, Day59, De 51, Dem55, Dem58, Den55, Der56, Des50, Dig57, Don56, Dor54, Dor50a, Dor56a, Dor56b, Dor56c, Dor55, Dow54, Dre58, Dud52, Dun54, Dun57a, Dun59a, Dur51, Dwa54, Dwy51, Dwy59, Eck58, Edw58, Eis53, Eis55, Eis59, Elv58a, Enk56, Eps54, Eva57, F.51a, F.51b, F.51c, F.51d, F.56a, F.56b, F.57a, F.57b, F.57c, F.58a, F.58b, F.58c, F.59, Fan58, Fel57, Fer55a, Fer58b, Fer50, Fey59, Fic55, Fin57, Fir51, Fis51, Fis54, Fol54, Foo59, Fox51b]. Review [Fox53, Fra51, Fra54, Fre50a, Fre51, Fre55, Fre59a, Fre50b, Fre59b, Fry51, Ful55, G.53, Gar52, Gar55, Gar56b, Ger53, Gil57a, Gil59, Gil57b, Gin52, Gir53, Glo56, Gog54, Gol58, Gol55a, Gol55b, Goo56, Goo53a, Goo58b, Gor56a, Gor56b, Gor58, Gou52, Gra59, Gra58, Gre52a, Gre53a, Gre52c, Gre56a, Gre57, Gre58a, Gre55, Gre58b, Gre54, Gre56b, Gre52d, Gri57a, Gri57b, Gur54, Gut56, Had55, Hae55a, Hae55b, Hae55c, Hae57a, Hae57b, Hae57c, Hag52, Hal50, Hal55a, Hal53, Hal55b, Ham54a, Ham52, Ham54b, Han56, Har51a, Har56a, Har55, Har59, Har52, Har51b, Har53, Har56c, Has56, Has51, Has53, Hau56, Hei56, Hen56, Hen57, Her57, Hic55, Hil55a, Him55, Hir55, Hoa50, Hoa57, Hod52, Hod55, Hoe55, Hol51, Hol58, Hoo57, Hor57, Hor55]. Review [Hos56, Hou53, Hou56, Hou54a, How57, How58, Hug52, Hul56, Hym57, Hym58, Ire52, Ire58, Isa57, Jac53, Jac56, Jaf52, Jaf56, Jah58, Jah54, Jam50, Jar56, Jeb51, Joh50a, Joh54, Joh56c, Joh56a, Joh56b, Jon51, Jon52, Jon56c, Jon56a, Jon57, Jon58a, Jos58, Kah56, Kai59a, Kal51, Kap59a, Kar56, Kar58a, Kat55, Kat56a, Kat56b, Kat56c, Kem54, Ken52, Ken56a, Ken56b, Ken58, Ken57a, Ken57b, Ker56, Kes56, Kim56, Kla58, Kle57, Kle52, Kni50, Kno54, Koc52, Koh56, Kon52, Kos53, Koz55, Kra59, Kru51, Kru58a, Kru59, Kul58, Kur51, Kuz50a, L.51a, L.51b, L.51c, L.54a, L.54b, L.55, Lac56, Lan56, Lau55a, Lau55b, LeN56, Lee51, Lee53, Lee56b, Lee56a, Lee59, Leh53, Ler57, Lev57, LF54, Lie53, Lie55, Lie58a, Lin54, Lin58, Lin59, Loc54, Loe51, Lof50]. Review [Lon56, Lor53, Lor52a, Lor52b, Low58, M.54, M.59, Mac52, Mac59a, Mac56, Mac55, Mad53, Mad55, Mad58a, Mad58b, Mad58c, Mai58, Mai56a, Mai56b, Mai54, Mal53, Man52, Man58, Mar58a, Mar57, Mar52, Mar53, Mar56, Mar58b, Mas59, Mat57, Mau58, May55, May58, May50, McA59, McC57, McD52, McG58, McI57, McI55, McP58, Mei58, Mer50, Mer51, Met50a, Mey59, Mic59a, Mil56b, Mil56c, Mil56d, Min57b, Mit51, Mit57b, Mod56, Mor54, Mor55, Mor51b, Mos51, Mos56a, Mur55a, Mut58, Nat57, Net58, New57, Nic51, Nis54, Noe56a, Noe57a, Noe59, Nor52a, Nor55b, Nor54a, Nor54b, Nut58, Nut59, O’R53, Old51a, Old55, Olk57, Orc55, Ost54, Ott57, Pab50, Pab55, Pad50, Pag57, Pal53, Par56, PW57, Par59, Per59, Pet55, Pet57a, Pet57b, Pfo55, Pot56, Pri53, Pri54b, Pri54a]. Review [Pyk57, R.51, R.53, Raf57, Rai55, Rap58, Ras56, Rau58, Rei52, Rei53, Rei58, Res58, Rev50, Ric56, Ric57, Ric59, Rio56, Rob55, Rob57a, Rob58, Rog56, Roo51, Roo52, Ros57, Ros59, Ros52, Rot55a, Rot55b, Rub58, Rud53, S.51, S.58, Sar56, Sav51a, Sav55, Sav56, Sav58b, Sav59b, Sav57c, Sax56, Sch55, Sch56a, Sch57a, Sch57b, Sch50a, Sch50b, Sch56b, Sco58, Sei56, Sew51, Sha57a, Sha52b, Sha57b, She58, Sho55, Shr54, Sie57, Sie59, Sil56, Sim51a, Sim53b, Sim53c, Sit53, Sle51, Sle57, Smi50b, Smi50c, Smi52, Smi55a, Smi55b, Smi58c, Smi58b, Smi58a, Sne51, Sne55, Sny56, Sol55a, Sol55b, Sol56a, Sol52b, Sol52a, Sol55c, Sol56b, Spe57, Spi51, Spi55, Spi57, Spr59a, Spr53, Spr59b, Ste57a, Ste52a, Ste51a, Ste52b, Ste53, Ste56, Ste50, Ste55a, Ste54]. Review [Ste58, SB52, Ste57b, Sti51, Sti52, Sto58a, Sto55c, SH51, Str53, Sui53, Sui58, Sup55, Sup58, Swa54, Sza54, Tae56, Tei56, Tei57, Tep50, Ter56a, Ter56b, Tho59, Tic58a, Tic58b, Tin58a, Tin58b, Tip52, Tip55, Tip59, Tol55, Tor52, Tru56, Tuk52, TW52, Tur57, Tyl54, Tyl58, Ull52, Ulm50, Val58, Van55, Vil52, Vil53, Vin51, W.51a, W.51b, W.51c, W.52a, W.52b, W.52c, W.52d, W.52e, W.53, W.54a, W.54b, W.54c, W.54d, W.54e, W.55a, W.55b, W.55c, W.55d, W.55e, W.56a, W.56b, W.56c, W.56d, W.56e, W.57a, W.57b, W.57c, W.58a, W.58b, W.58c, W.58d, W.58e, W.58f, W.58g, W.59, Wag59a, Wal56a, Wal57, Wal55, Wal56b, Wal58b, Wal54, Web59, Wei51, Wei59, Wei57, Wel54, Wel57, Wes58, Wes55, Wes51, Wil52, Wil56a, Wil55b, Wis58, Wit56, Wol51, Wol52, Wol55, Wol59, Woy50]. Review [Wri51, Wri52, Wri56, Wro56, Wyc58, Yer57, Yod52, Zei59, Zio55, Zuc57, dC56]. Revision [Coa55b, Coa55c]. Revolution [You51]. Rico [Rot52]. Risk [Spr50, TM59]. Risks [Gar56a]. Rocks [Kru54]. Root [Dur56]. Ross [Hou53]. Runs [Wei53, Wei54, Wei56].

Sales [HS52]. Same [BDL^{+}54]. Sample [Bir52, CF58, Dem54, EBS55, Fer55b, Gol53, GS50, Jon53, Jon56b, Jon59, KH58, KM51, Kru57, Mac59c, Mac59b, Mar51, Mit57a, Moo51, OS59, PS50, Rid57, Wal50a, Wal50b, Wal51, Wel58]. Samples [BHR53, Coh50, Coh54, Coh55a, Coh55b, DG59, HGCZ55, Hyr53a, Hyr53b, Jon58b, KH59, Moo55, Moo56, Mur55b, Nic57, Raj53, Rid51a, Rid51b, Rid53a, Rid53b, Rid55b, Rid56, Rid59, SG57, Ste51b, KH63]. Sampling [BS50, BF53, Chu50, CMT54, Dem56, Dun58, EB56, Gar56a, Gar58, Gol53, GK50, HV55, HT52, Jeb52, Jon55a, Kat53, Key51, Key57, LR55a, LR55b, Mad50, Man53, Mos55, Mos56b, Raj56, Raj58, RK52, RK53, RBJ51, RGM58a, RGM58b, Sav59a, Sea59, SAF54, Sha52a, Slo57, Som57, Ste55b, Tuk50a, Tuk50b, Zar54, Zar55]. Saving [Zel58]. Scale [Sea58]. Scaling [Coo51a, Met50b]. Scheffe [Dyk58]. Schemes [Lie52]. Science [Pit57, Spo51, Yat51]. Sciences [Tou55]. Scientist [Ber53b]. Scientists [TW52]. Scoring [Gri51]. Seasonal [Eis56, Mud55, Rot58, SE57]. Seats [Wil54, Wil56b]. Section [LB58]. Secured [Fer55b]. Sedimentary [Kru54]. Seen [Jas51b]. Selected [Key57, Tho52, Whi54, Zel58]. Selecting [Jon59]. Selection [BD59, Bro50, Coh55b, GK50]. Sense [AC55]. Sensitivity [GG56a, GG56b, GG67]. Separate [Qua58]. September [Alc55, Alc57, Alc58b]. Sequences [Wei58]. Sequential [Bra58b, Bra58a, Chu50, DN58, Gar56a, Noe56c, Wei54]. Serial [Goo52, Goo54, Wei58]. Series [Ask57, Ber57, Goo58a, Mos52, Pre55, Tuk52, Wei55]. Service [PJ57a, PJ57b]. Set [Bab55, Moo57a]. Sets [Mei54]. Several [Dun55, She52, Wel56]. Sex [Coa55b, Coa55c]. Shaped [TL55]. Shevky [Bue56]. Shifts [Eis56]. Shoreham [Ano54g]. Short [Ask57, Sch52b, Spu51, Wei55, Wei58]. Short-Cut [Ask57, Sch52b, Spu51]. Should [Cro52]. Showing [Kat53]. Side [Ber61]. Sided [Sea59]. Sigmoid [Bro52a]. Sign [Blu58b, Coh59a, Ste59a, Stu52]. Signed [Pra59, Wal59b]. Signed-Rank [Wal59b]. Significance [Bir54a, Goo58a, Kim53, Ste59b, Tul59]. Simple [Ber53a, Mei54]. Simplest [Wal59b]. Simplifications [Dem56]. Simplified [SM50, Tuk50b, Wal59a]. Single [Gol53, SG57, Sea59]. Single-Sampling [Gol53]. Single-Sided [Sea59]. Singly [SG57]. Size [Ade58, Ade59, Bir52, Gol53, GT56, GS50, Key51]. Sizes [Mac59c, Mac59b, Moo51, Wal50a]. Skewed [Mil55, Sar55]. Slope [DH50]. Small [Bal51, BHR53, Chu50, Hyr53a, Hyr53b, Wel58]. Smaller [Wal56c]. Smallest [Ram58a, Ram59]. Smirnov [Mas51]. Smoking [Ber58, Cut55, Ham59]. Smoothing [Tuk52]. Social [Fab52, Gra55, Gra56, Jon55b, Lan59a, Lan59b, Bue56]. Solitaire [Gre53b]. Solution [HD53]. Solutions [Bab55, Wag58]. Some [And55b, Bal51, Ber58, Bur58, CM50a, CM50b, Cro52, Dal58, Dav52, DN58, Eps55, Fel59, Goo54, GWS50, Kap59b, Kin53, Mic59b, Mic60, Mos56b, Net52, NW59, Noe57b, Raj56, Raj58, SH52, Tuk50b, VK52, Wal52b, dlGHN55, And57]. Source [Dea52, Sel55, Ste51b, Wei50]. Sources [Bea59, Pri55]. Soviet [Fel59, Jas50, Sie53]. Spatial [Rot52, Vin53, War56]. Special [War56]. Specific [Yer55, Yer56]. Specification [Sea59, TM59]. Spurious [Sim54]. Square [Ber55a, Bla58a, Bla58b, Bra58b, Bra58a, Dur56, Kim53, Moo55, Moo56, Ram58a, Ram59, Rao59, WK57, Wil50]. Squares [Ask57, Li55]. Stability [Wal52a]. Stage [KH59, KH63, SAF54]. Stages [Dem56]. Standard [Eva51b, GS50, Noe55, Noe56b]. Standards [Kar58b]. State [Gri59b, Mad50, Whi54]. States [BK54, Coa55b, Coa55c, Cro59, Fis52, Hut58, Kis52, Mil51, Mye54a, Ric54, Sog54]. Stationary [Tuk52]. Statistic [Bir52, Wal51]. Statistical [Alc53, Alc54, Alc55, Alc56, Alc57, Alc58b, Alc59, And55b, Ano50g, Ano50l, Ano51h, Ano52f, Ano53h, Ano54g, Ano54n, Ano54o, Ano54p, Ano54q, Ano55d, Ano55l, Ano55m, Ano56l, Ano56m, Ano56n, Ano56o, Ano57m, Ano57n, Bal51, BPR55, Bir54b, BS58, Bro52a, Cla59, CMT53, Cox57b, Cut55, Fel59, Kim57, KM51, Kru54, Kuz50b, Lie58b, Mat51, McL51, Min57a, Nic55, Ric54, Sav51b, Sav54a, Spr50, Tei55, TW52, Vin53, Wil51, Woo51a, Woo51b, Woo51c, Woo52a, Woo52b, Woo52c, Yat51, Zel58, Bir66]. Statistically [Ber53a]. Statistician [Hir56]. Statisticians [Adl51, Fit57, Fit58, Hoa59]. Statistics [Ano50f, Ano53c, Ano53d, Bea59, BHP53a, BHP53b, Bro55c, Bur54, Car53, Che56, CL54b, Cop55, CG59, Dan53, Dun52, Fab52, Hot51, Hut58, Jas50, Lan59a, Lan59b, Lee52, Mil56a, Net52, Old51b, Old54, Osh57, Pit57, Sar55, SG57, Sav53, Sav57a, Sav58a, Tou55, Tuk54, Vot52, Wic53, Win59, Yat51, Zak56]. Statistische [TW52]. Status [Jas51b]. Steamships [HR58]. Stereotypy [Fre50c]. Stimulas [GG56b]. Stimulus [GG56a, GG67]. Stochastic [Ade58, Ade59]. Stochastically [Mar51, Moo57a]. Straight [Ask57, Mad59a, Mad59b, Man57, Sch58]. Strategy [BCMM56, BCMM59]. Stratification [DH59, Eva51a, DH63]. Stratified [SAF54]. Stratum [Key57]. Stream [Raf51]. Structure [Coc54b, Fer53, Fer58a, Gre52b, Men55, Vin53]. Student [Wel58, Fed59, Wal52b]. Studentized [HGCZ55, Ram58a, Ram59]. Studies [Lan59a, Lan59b]. Study [CMT55, Gri59b, KHS^{+}55, MG55, Tou55, Vin53, Whi54]. Sub [Elk53, Jeb52]. Sub-Class [Elk53]. Sub-Sampling [Jeb52]. Subject [Mad59a, Mad59b]. Subjective [TJ57]. Subsample [Jon56b]. Subsampling [Bro55a]. Success [Vot52]. Successive [Moo55, Moo56, Rao59]. Sum [Li55]. Summaries [Alc53, Alc54, Alc55, Alc56, Alc57, Alc58b, Alc59]. Surface [BD59]. Survey [Bel54, Fer56, HHNS55, HM58, KM51, Lie58b, MMN53, NW59, Ros53, Ros54, Spo51, Ste55b, VK52]. Surveys [BS50, Bro57b, Dem53b, Dem54, Fer55b, Gra55, Gra56, HHMM51, Sha52a]. Survival [BG52]. Survivors [Man53]. Survivorship [Elv58b]. Suspected [Kin53]. Symmetric [Mac59c, Mac59b]. Symmetry [Mil55]. symposium [Tuk50a, KHS^{+}55, CMT55]. Syphilis [GWS50]. System [Qua58, Ric54, Vin53]. Systematic [Hol57, Sar55]. Systems [GF59, SAF54].

Table [Mac59c, Mac59b, Mil56a]. Tables [BK57, Coh59a, Cra53b, Fed59, GS54a, GS54b, GGD53, HGCZ55, JR55a, JR55b, Mit57a, SG57, Tei55]. Tabular [BAR56, BGR56]. Tabulation [Bir52]. Tabulations [EBS55]. Target [Coh55a]. Tasting [Gri59a]. Tattersfield [Bro52a]. Tau [Adl57, Gri58]. Tax [Cru56, RBJ51, Sel55, Tho52]. Tchebychef [God55]. Tea [Gri59a]. Teaching [Bur54, Old54]. Technical [Smi54]. Technique [Bog50, Bro55b, GK50, GWS50, HD53]. Techniques [Eis56, Goo54, Raj58, SM55, Wag59b]. Term [BDL^{+}54, Ger57]. Test [AD54, AC55, Bak58, Blu58b, CR56, Coc55, Coh59a, DN58, JR55a, JR55b, Kru57, Lor59, Mar51, Mas51, Ram58b, Ste59a, Vot52, Wal59a, Wal50a, Wal56c, Wil50, dlG56]. Testing [Bar57, ES53, Eps55, GF59]. Tests [Ban50, Bir54a, Bre56, CF58, Dun57b, Eps55, Goo58a, Noe56c, Sav57b, Ste59b, Stu52, Stu54, Stu56, Tul59, Wal50b, Wal52a, Wal52b, Wal59b]. Their [KH59, KH63, Ste59b]. Theoretical [Che56, Mos56b, Spo51]. Theory [Gil50, GGD53, Kem55, Kim55a, Kim55b, Lor59, Rob57b, Sav51b, Sav54a, SM55, Sim51b, Wel58, kem56]. Therapy [BDL^{+}54]. There [Ber50, Dur53]. Think [Met50a]. Third [Kim57]. Those [Fer58a]. Ties [Adl57, Pra59]. Time [Gar53, Pre55, Tuk52, Wei55]. Time-Series [Wei55]. times [BK57]. Tinbergen [Arr58c]. Today [Hoa59]. Tolerance [Mit57a, Pro53]. Tomorrow [Hoa59]. Topics [Ano53c, Ano53d, Gri59a, Sav53, Sav58a]. Tract [CG59]. Tracts [Fol53]. Trade [Ade55]. Traffic [Raf51]. Transformation [Moo57a, Sim51b]. Transformations [Moo57b]. Transitory [DR58]. Transportation [Smi54]. Travel [LB58]. Treatment [BG52, Bro52a]. Treatments [Dun55, KM51, Ste59a]. Trend [Noe56c]. Trends [Kis52, Mor51a, Nor52b]. Trentile [Slo58]. Trials [Bro55c]. True [GS50]. Truncated [Coh50, Coh54, Coh55a, Hal52, Raj53, Rid53c, Rid55a, Rid55c]. Truncation [Cla57a]. Twelve [Mac59c, Mac59b]. Two [Ber50, Ber58, BS52, Bro50, Bue57, CF58, Elk53, Gay56, Gea53, Key57, Kit55, Kit56, Kru57, Mar51, Noe56c, Noe57b, Pri55, Qua58, SAF54, Sto54, Sto55a, Stu52]. Two-Choice [Bro50]. Two-Population [Sto54, Sto55a]. Two-Sample [CF58, Kru57]. Two-Stage [SAF54]. Type [Coh50, GH58, GH63, Ham59, Rob57b, Wei54, Raj53, Sea58]. Types [Cla59, Man57].

Unbiased [GH58, GH63, GW56, Mic59b, Mic60, Rob57b]. Uncongested [Raf51]. Underdeveloped [Abr58, Osh57]. Undergraduate [Wil51]. Unemployed [HP50]. Unequal [Mei54]. Uniformly [GW56]. Unimodality [Ber57]. Union [Sie53]. Unit [Jeb52, Mad50]. United [BK54, Coa55b, Coa55c, Cro59, Fis52, Hut58, Kis52, Mil51, Mye54a, Ric54, Sog54]. Units [Key57, Rei59]. Univariate [Sto54, Sto55a]. Universe [HT52]. Unknown [Hoo58]. Unpaired [Kru57]. Unpublished [Tei55]. Unsolved [Tuk54]. untitled [Ada54, Adl55, Alc58a, Ale57, Alt58, And50, And56, And55c, Ang57, Ano51m, Ano51n, Ano51o, Ano51p, Ano54r, Ano55n, Ano56p, Ans56, Ans59a, Ara52, Ark56, Arn52, Arr54a, Arr54b, Arr58a, Arr58b, Atk55, B.56, B.57a, B.57b, B.58, Bac53, Bac58, Bah57, Bai58, BO53, Bak56, Bal56, Ban58, Bar58, Bea57, Bea54, Bea50, Bel55, Ben57, Ben50a, Ben50b, Ben55, Ber56, Bey52, Bir55, Bir56, Bla51, Bla58c, Bla56, Ble58, Blu58a, Blu54, Bod57, Bod59, Boh55, Bor57, Bos58, Bou59a, Bou59b, Bow50a, Bow50b, Bow56, Box55, Bra56a, Bra51a, Bra51b, Bra54a, Bra56b, Bra58c, Bra59a, Bra59b, Bra55, Bre58, Bri50, Bri59, Bri56, Bro58a, Bro53a, Bro58b, Bro59, Bro53b, Bro56a, Bro56b, Bro52b, Bro57a, Bro58c, Bru55, Bry57, Buc54, Bud56, Bur55, Bus56a, Bus56b, BC56]. untitled [Cam57, Can50, Cha51, Cha52, Che55a, Che55b, Chi54, Chi59, Chi53, Chu57, Chu51, Cla50, Cla52, Cla53a, Cla57b, Coa55a, Coc52, Coc54a, Coc56, Coc57, Coh56, Coh57, Col57, Col56, Com59, Con52, Con56, Coo51b, Coo58, Cor50, Cor51, Cor55, Cor58, Cov51, Cox59, Cra50, Cra53a, Cra53c, Cro50, Cur54, Dal54, Dar51, Dar55, Dav59a, Dav59b, Day51, Day59, De 51, Dem55, Dem58, Den55, Der56, Des50, Dig57, Don56, Dor54, Dor50a, Dor56a, Dor56b, Dor56c, Dor55, Dow54, Dre58, Dud52, Dun54, Dun57a, Dun59a, Dur51, Dwa54, Dwy51, Dwy59, Eck58, Edw58, Eis53, Eis55, Eis59, Elv58a, Enk56, Eps54, Eva57, F.51a, F.51b, F.51c, F.51d, F.56a, F.56b, F.57a, F.57b, F.57c, F.58a, F.58b, F.58c, F.59, Fan58, Fel57, Fer55a, Fer58b, Fer50, Fey59, Fic55, Fin57, Fir51, Fis51]. untitled [Fis54, Fol54, Foo59, Fox51b, Fox53, Fra51, Fra54, Fre50a, Fre51, Fre55, Fre59a, Fre50b, Fre59b, Fry51, Ful55, G.53, Gar52, Gar55, Gar56b, Ger53, Gil57a, Gil59, Gil57b, Gin52, Gir53, Glo56, Gog54, Gol58, Gol55a, Gol55b, Goo56, Goo53a, Goo58b, Gor56a, Gor56b, Gor58, Gou52, Gra59, Gra58, Gre52a, Gre53a, Gre52c, Gre56a, Gre57, Gre58a, Gre55, Gre58b, Gre54, Gre56b, Gre52d, Gri57a, Gri57b, Gur54, Gut56, Had55, Hae55a, Hae55b, Hae55c, Hae57a, Hae57b, Hae57c, Hag52, Hal50, Hal55a, Hal53, Hal55b, Ham54a, Ham52, Ham54b, Han56, Har51a, Har56a, Har55, Har59, Har52, Har51b, Har53, Har56c, Has56, Has51, Has53, Hau56, Hei56, Hen56, Hen57, Her57, Hic55, Hil55a, Him55, Hir55, Hoa50, Hoa57, Hod52, Hod55, Hoe55, Hol51]. untitled [Hol58, Hoo57, Hor57, Hor55, Hos56, Hou56, Hou54a, How57, How58, Hug52, Hul56, Hym57, Hym58, Ire52, Ire58, Isa57, Jac53, Jac56, Jaf52, Jaf56, Jah58, Jah54, Jam50, Jar56, Jeb51, Joh50a, Joh54, Joh56c, Joh56a, Joh56b, Jon51, Jon52, Jon56c, Jon56a, Jon57, Jon58a, Jos58, Kah56, Kai59a, Kal51, Kap59a, Kar56, Kar58a, Kat55, Kat56a, Kat56b, Kat56c, Kem54, Ken52, Ken56a, Ken56b, Ken58, Ken57a, Ken57b, Ker56, Kes56, Kim56, Kla58, Kle57, Kle52, Kni50, Kno54, Koc52, Koh56, Kon52, Kos53, Koz55, Kra59, Kru51, Kru58a, Kru59, Kul58, Kur51, Kuz50a, L.51a, L.51b, L.51c, L.54a, L.54b, L.55, Lac56, Lan56, Lau55a, Lau55b, LeN56, Lee51, Lee53, Lee56b, Lee56a, Lee59, Leh53, Ler57, Lev57, LF54, Lie53, Lie55, Lie58a, Lin54, Lin58, Lin59, Loc54]. untitled [Loe51, Lof50, Lon56, Lor53, Lor52a, Lor52b, Low58, M.54, M.59, Mac52, Mac59a, Mac56, Mac55, Mad53, Mad55, Mad58a, Mad58b, Mad58c, Mai58, Mai56a, Mai56b, Mai54, Mal53, Man52, Man58, Mar58a, Mar57, Mar52, Mar53, Mar56, Mar58b, Mas59, Mat57, Mau58, May55, May58, May50, McA59, McC57, McD52, McG58, McI57, McI55, McP58, Mei58, Mer50, Mer51, Mey59, Mic59a, Mil56b, Mil56c, Mil56d, Min57b, Mit51, Mit57b, Mod56, Mor54, Mor55, Mor51b, Mos51, Mos56a, Mur55a, Mut58, Nat57, Net58, New57, Nic51, Nis54, Noe56a, Noe57a, Noe59, Nor52a, Nor55b, Nor54a, Nor54b, Nut58, Nut59, O’R53, Old51a, Old55, Olk57, Orc55, Ost54, Ott57, Pab50, Pab55, Pad50, Pag57, Pal53, Par56, PW57, Par59, Per59, Pet55, Pet57a, Pet57b, Pfo55, Pot56, Pri53, Pri54b]. untitled [Pri54a, Pyk57, R.51, R.53, Raf57, Rai55, Rap58, Ras56, Rau58, Rei52, Rei53, Rei58, Res58, Rev50, Ric56, Ric57, Ric59, Rio56, Rob55, Rob57a, Rob58, Rog56, Roo51, Roo52, Ros57, Ros59, Ros52, Rot55a, Rot55b, Rub58, Rud53, S.51, S.58, Sar56, Sav51a, Sav55, Sav56, Sav58b, Sav59b, Sav57c, Sax56, Sch55, Sch56a, Sch57a, Sch57b, Sch50a, Sch50b, Sch56b, Sco58, Sei56, Sew51, Sha57a, Sha52b, Sha57b, She58, Sho55, Shr54, Sie57, Sie59, Sil56, Sim51a, Sim53b, Sim53c, Sit53, Sle51, Sle57, Smi50b, Smi50c, Smi52, Smi55a, Smi55b, Smi58c, Smi58b, Smi58a, Sne51, Sne55, Sny56, Sol55a, Sol55b, Sol56a, Sol52b, Sol52a, Sol55c, Sol56b, Spe57, Spi51, Spi55, Spi57, Spr59a, Spr53, Spr59b, Ste57a, Ste52a, Ste51a, Ste52b, Ste53, Ste56, Ste50, Ste55a]. untitled [Ste54, Ste58, SB52, Ste57b, Sti51, Sti52, Sto58a, Sto55c, SH51, Str53, Sui53, Sui58, Sup55, Sup58, Swa54, Sza54, Tae56, Tei56, Tei57, Tep50, Ter56a, Ter56b, Tho59, Tic58a, Tic58b, Tin58a, Tin58b, Tip52, Tip55, Tip59, Tol55, Tor52, Tru56, Tur57, Tyl54, Tyl58, Ull52, Ulm50, Val58, Van55, Vil52, Vil53, Vin51, W.51a, W.51b, W.51c, W.52a, W.52b, W.52c, W.52d, W.52e, W.53, W.54a, W.54b, W.54c, W.54d, W.54e, W.55a, W.55b, W.55c, W.55d, W.55e, W.56a, W.56b, W.56c, W.56d, W.56e, W.57a, W.57b, W.57c, W.58a, W.58b, W.58c, W.58d, W.58e, W.58f, W.58g, W.59, Wag59a, Wal56a, Wal57, Wal55, Wal56b, Wal58b, Wal54, Web59, Wei51, Wei59, Wei57, Wel54, Wel57, Wes58, Wes55, Wes51, Wil52, Wil56a, Wil55b, Wis58, Wit56, Wol51, Wol52, Wol55, Wol59, Woy50, Wri51]. untitled [Wri52, Wri56, Wro56, Wyc58, Yer57, Yod52, Zei59, Zio55, Zuc57, dC56]. Up-and-Down [BHR53]. Upon [Ask59]. Urban [Fol53]. Usable [Jon59]. Use [AC55, BGR56, Bur54, CL54a, DN58, Hoo58, Hyr53a, Hyr53b, JR55a, JR55b, KW52, KW53, Mad50, Noe55, Noe56b, RGM58b, Rot58, SH52, Sui57, Tou55, Wei53, Wei56, BAR56, RGM58a]. used [Dun56a]. Useful [Goo53b].

Uses [EBP55, Sto55b]. Using [Kar58c, Moo57b].

Vaccine [Bro55c]. Validation [Bel54]. Value [GS54a, GS54b, GS50, KL54, Zak56]. Values [Ban50, Jon53, Kim55a, Kim55b, Mos53, Mur55b, Rid55b, Rid56, Rid59, TJ57]. Variable [Ait55, Gea53, Moo57b, Sch58]. Variables [Dun58, Dun59b, Edg56, Gar56a, Gar58, Gea53, GS54a, GS54b, LR55a, LR55b, Mad59a, Mad59b, Mar51, Moo57a, Sea59, She52, Sui57, Wes52]. Variance [BPR55, DH59, DH63, Dun57b, Dyk58, Eva51b, Gay56, GW56, Kai59b, Key57, Kin54, KW52, KW53, Sch52a, TK59]. Variances [Dav56, KH59, KH63, Ram58b, Rei56, Wal50a, Wel56]. Various [Moo55, Moo56]. Varying [DR58, Raj56, Rot58]. Velocity [Gar53]. Verification [Gri51]. Versa [Ste59b]. Versus [Ste59a]. Vice [Ste59b]. Vital [CG59, Sto55b]. Volume [Ano50q, Ano51q, Ano52l, Ano53p, Ano54s, Ano55o, Ano56q, Ano57o, Ano58k, Ano59p, Leo50].

W [Hou53]. Wage [CL54b]. Wallis [Wal59a]. War [Kar58b]. Washington [Alc54, Ano54g]. Wealth [Ger57]. Weather [Gri51, Slo58]. Weight [Kar58b]. Weight-Height [Kar58b]. Weighted [Hic58, Jon53]. Weighting [Ros53, Ros54]. Weights [Rot58]. Wendell [Bue56]. Wesley [Ber53b]. Whelpton [Kar50]. Where [Key57]. Which [Cro52, Wal51]. Who [HP50]. Width [Wil50]. Wiener [Tuk52]. Wilcoxon [Kru57, Pra59]. Will [Jon59]. Willard [Jas51a]. Without [HT52, Kat53, PS50, Dem56, Raj56]. Women [Bel50]. Work [RGM58a, RGM58b]. Worker [Rot52]. Workers [Mat51, Yat51]. World [Kar58b, Mos52, Not50].

Yates [Adl51, Cro52]. Year [Cru56, Not50]. York [Alc56, Ano50l]. Yugoslav [Zar55].

Zeros [Pra59].

References

Abraham:1958:IEU

[Abr58] William I. Abraham. Investment estimates of underdeveloped countries: An appraisal. Journal of the American Statistical Association, 53(283):669–679, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

          org/stable/2282045.

Appel:1955:RUM

[AC55] Valentine Appel and George Cooper. A refinement in the use of mark-sense cards for test research. Journal of the American Statistical Association, 50(270):557–560, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280976.

Anderson:1954:TGF

[AD54] T. W. Anderson and D. A. Darling. A test of goodness of fit. Journal of the American Statistical Association, 49(268):765–769, December 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281537.

Adams:1954:BRB

[Ada54] Leonard P. Adams. Book review: [untitled]. Journal of the American Statistical Association, 49(267):675–677, September 1954. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281151.

Adelman:1955:FTC

[Ade55] M. A. Adelman. Federal Trade Commission report on changes in concentration in manufacturing. Journal of the American Statistical Association, 50(271):660–664, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281157.

Adelman:1958:SAS

[Ade58] Irma G. Adelman. A stochastic analysis of the size distribution of firms. Journal of the American Statistical Association, 53(284): 893–904, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281959.

Adelman:1959:CSA

[Ade59] Irma G. Adelman. Corrigenda: a stochastic analysis of the size distribution of firms. Journal of the American Statistical Association, 54(288):810, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282506.

Adler:1951:YCS

[Adl51] Franz Adler. Yates’ correction and the statisticians. Journal of the American Statistical Association, 46(256):490–501, December 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280398.

Adler:1955:BRB

[Adl55] John H. Adler. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1391–1394, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281262.

Adler:1957:MKT

[Adl57] Leta McKinney Adler. A modification of Kendall’s tau for the case of arbitrary ties in both rankings. Journal of the American Statistical Association, 52(277):33–35, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281397.

Aitchison:1955:DPR

[Ait55] John Aitchison. On the distribution of a positive random variable having a discrete probability mass at the origin. Journal of the American Statistical Association, 50(271):901–908, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2281175.

Aroian:1950:EQC

[AL50] Leo A. Aroian and Howard Levene. The effectiveness of quality control charts. Journal of the American Statistical Association, 45 (252):520–529, December 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280720.

Aroian:1952:EEQ

[AL52] Leo A. Aroian and Howard Levene. Erratum: The effectiveness of quality control charts. Journal of the American Statistical Association, 47(260):685, December 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280785.

Alchian:1953:SPD

[Alc53] Armen A. Alchian. Summaries of papers delivered at the 112th Annual Meeting of the American Statistical Association in Chicago, December 27 to 30, 1952. Journal of the American Statistical Association, 48(263):615–644, September 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281013.

Alchian:1954:SPD

[Alc54] Armen A. Alchian. Summaries of papers delivered at the 113th Annual Meeting of the American Statistical Association in Washington, D.C., December 27 to 30, 1953. Journal of the American Statistical Association, 49(266):323–370, June 1954. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2280938.

Alchian:1955:SPD

[Alc55] Armen A. Alchian. Summaries of papers delivered at the 114th Annual Meeting of the American Statistical Association in Montreal, September 10–13, 1954. Journal of the American Statistical Association, 50(270):561–599, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280977.

Alchian:1956:SPD

[Alc56] Armen A. Alchian. Summaries of papers delivered at the 115th Annual Meeting of the American Statistical Association in New York City, December 27–29, 1955. Journal of the American Statistical Association, 51(275):501–527, September 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281439.

Alchian:1957:SPD

[Alc57] Armen A. Alchian. Summaries of papers delivered at the 116th Annual Meeting of the American Statistical Association in Detroit, September 7–10, 1956. Journal of the American Statistical Association, 52(279):361–383, September 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280907.

Alchian:1958:BRB

[Alc58a] Armen A. Alchian. Book review: [untitled]. Journal of the American Statistical Association, 53(284):1047–1048, December 1958. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281986.

Alchian:1958:SPD

[Alc58b] Armen A. Alchian. Summaries of papers delivered at the 117th Annual Meeting of the American Statistical Association in Atlantic City, September 10–13, 1957. Journal of the American Statistical Association, 53(282):565–594, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281877.

Alchian:1959:SPD

[Alc59] Armen A. Alchian. Summaries of papers delivered at the 118th Annual Meeting of the American Statistical Association, Chicago Illinois, December 27–30, 1958. Journal of the American Statistical Association, 54(286):484–509, June 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281787.

Alexander:1957:BRB

[Ale57] Kenneth O. Alexander. Book review: [untitled]. Journal of the American Statistical Association, 52(280):596, December 1957. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281725.

Alt:1958:BRB

[Alt58] Franz L. Alt. Book review: [untitled]. Journal of the American Statistical Association, 53(282):624–625, June 1958. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281900.

Anderson:1950:BRB

[And50] Paul H. Anderson. Book review: [untitled]. Journal of the American Statistical Association, 45(250):308, June 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280698.

Anderson:1953:RAF

[And53] R. L. Anderson. Recent advances in finding best operating conditions. Journal of the American Statistical Association, 48(264): 789–798, December 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281072.

Anderson:1954:PAR

[And54] R. L. Anderson. The problem of autocorrelation in regression analysis. Journal of the American Statistical Association, 49(265):113– 129, March 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281039.

Anderson:1955:CPA

[And55a] R. L. Anderson. Corrigenda: The problem of autocorrelation in regression analysis. Journal of the American Statistical Association, 50(272):1331, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281218.

Anderson:1955:SSP

[And55b] T. W. Anderson. Some statistical problems in relating experimental data to predicting performance of a production process. Journal of the American Statistical Association, 50(269):163–177, March 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281103.

Andrews:1955:BRB

[And55c] Fred C. Andrews. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1337–1338, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281230.

Anderson:1956:BRB

[And56] R. L. Anderson. Book review: [untitled]. Journal of the American Statistical Association, 51(276):667–669, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281504.

Anderson:1957:MLE

[And57] T. W. Anderson. Maximum likelihood estimates for a multivariate normal distribution when some observations are missing. Journal of the American Statistical Association, 52(278):200–203, June 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280845.

Angell:1957:BRB

[Ang57] James W. Angell. Book review: [untitled]. Journal of the American Statistical Association, 52(280):599–602, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281729.

Anonymous:1950:BMa

[Ano50a] Anonymous. Back matter. Journal of the American Statistical Association, 45(249):??, March 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280445.

Anonymous:1950:BMb

[Ano50b] Anonymous. Back matter. Journal of the American Statistical Association, 45(250):??, June 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280701.

Anonymous:1950:BMc

[Ano50c] Anonymous. Back matter. Journal of the American Statistical Association, 45(251):478, September 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280314.

Anonymous:1950:BMd

[Ano50d] Anonymous. Back matter. Journal of the American Statistical Association, 45(252):??, December 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280738.

Anonymous:1950:BN

[Ano50e] Anonymous. Briefer notices. Journal of the American Statistical Association, 45(251):473–475, September 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280312.

Anonymous:1950:EPS

[Ano50f] Anonymous. Erratum: Psychological statistics. Journal of the American Statistical Association, 45(249):148, March 1950. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280443.

Anonymous:1950:ESM

[Ano50g] Anonymous. Erratum: Statistical methods in research. Journal of the American Statistical Association, 45(250):308, June 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280699.

Anonymous:1950:FMa

[Ano50h] Anonymous. Front matter. Journal of the American Statistical Association, 45(249):??, March 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280423.

Anonymous:1950:FMb

[Ano50i] Anonymous. Front matter. Journal of the American Statistical Association, 45(250):??, June 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280674.

Anonymous:1950:FMc

[Ano50j] Anonymous. Front matter. Journal of the American Statistical Association, 45(251):??, September 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280289.

Anonymous:1950:FMd

[Ano50k] Anonymous. Front matter. Journal of the American Statistical Association, 45(252):??, December 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280717.

Anonymous:1950:PAS

[Ano50l] Anonymous. Proceedings: American Statistical Association 109th Annual Meeting: Hotel Biltmore, New York, New York: Minutes of the Annual Business Meeting. Journal of the American Statistical Association, 45(250):270–286, June 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280686.

Anonymous:1950:PRa

[Ano50m] Anonymous. Publications received. Journal of the American Statistical Association, 45(249):148, March 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280444.

Anonymous:1950:PRb

[Ano50n] Anonymous. Publications received. Journal of the American Statistical Association, 45(250):309–310, June 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280700.

Anonymous:1950:PRc

[Ano50o] Anonymous. Publications received. Journal of the American Statistical Association, 45(251):476–477, September 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280313.

Anonymous:1950:PRd

[Ano50p] Anonymous. Publications received. Journal of the American Statistical Association, 45(252):586, December 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280737.

Anonymous:1950:VI

[Ano50q] Anonymous. Volume information. Journal of the American Statistical Association, 45(252):587–594, December 1950. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2280716.

Anonymous:1951:AA

[Ano51a] Anonymous. Additional abstracts. Journal of the American Statistical Association, 46(256):517, December 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280403.

Anonymous:1951:BMa

[Ano51b] Anonymous. Back matter. Journal of the American Statistical Association, 46(255):416, September 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280539.

Anonymous:1951:BMb

[Ano51c] Anonymous. Back matter. Journal of the American Statistical Association, 46(256):540, December 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280422.

Anonymous:1951:FMa

[Ano51d] Anonymous. Front matter. Journal of the American Statistical Association, 46(253):??, March 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280088.

Anonymous:1951:FMb

[Ano51e] Anonymous. Front matter. Journal of the American Statistical Association, 46(254):??, June 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279917.

Anonymous:1951:FMc

[Ano51f] Anonymous. Front matter. Journal of the American Statistical Association, 46(255):??, September 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280506.

Anonymous:1951:FMd

[Ano51g] Anonymous. Front matter. Journal of the American Statistical Association, 46(256):??, December 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280392.

Anonymous:1951:PAS

[Ano51h] Anonymous. Proceedings: American Statistical Association 110th Annual Meeting: Congress Hotel, Chicago, Illinois, December 27, 1950: Minutes of the Annual Business Meeting. Journal of the American Statistical Association, 46(254):242–250, June 1951. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279925.

Anonymous:1951:PRa

[Ano51i] Anonymous. Publications received. Journal of the American Statistical Association, 46(253):145–146, March 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280117.

Anonymous:1951:PRb

[Ano51j] Anonymous. Publications received. Journal of the American Statistical Association, 46(254):277–278, June 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279942.

Anonymous:1951:PRc

[Ano51k] Anonymous. Publications received. Journal of the American Statistical Association, 46(255):414–415, September 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280538.

Anonymous:1951:PRd

[Ano51l] Anonymous. Publications received. Journal of the American Statistical Association, 46(256):538, December 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280421.

Anonymous:1951:Ua

[Ano51m] Anonymous. [untitled]. Journal of the American Statistical Association, 46(253):129–130, March 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280104.

Anonymous:1951:Ub

[Ano51n] Anonymous. [untitled]. Journal of the American Statistical Association, 46(253):130, March 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280105.

Anonymous:1951:Uc

[Ano51o] Anonymous. [untitled]. Journal of the American Statistical Association, 46(253):140, March 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280113.

Anonymous:1951:Ud

[Ano51p] Anonymous. [untitled]. Journal of the American Statistical Association, 46(253):140, March 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280114.

Anonymous:1951:VI

[Ano51q] Anonymous. Volume information. Journal of the American Statistical Association, 46(256):539–550, December 1951. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2280391.

Anonymous:1952:BM

[Ano52a] Anonymous. Back matter. Journal of the American Statistical Association, 47(258):354, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280776.

Anonymous:1952:FMa

[Ano52b] Anonymous. Front matter. Journal of the American Statistical Association, 47(257):??, March 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279973.

Anonymous:1952:FMb

[Ano52c] Anonymous. Front matter. Journal of the American Statistical Association, 47(258):??, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280739.

Anonymous:1952:FMc

[Ano52d] Anonymous. Front matter. Journal of the American Statistical Association, 47(259):??, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281308.

Anonymous:1952:FMd

[Ano52e] Anonymous. Front matter. Journal of the American Statistical Association, 47(260):??, December 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280778.

Anonymous:1952:PAS

[Ano52f] Anonymous. Proceedings: American Statistical Association 111th Annual Meeting: Hotel Copley–Plaza, Boston, Massachusetts, December 27, 1951. Journal of the American Statistical Association, 47(258):305–312, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280754.

Anonymous:1952:PRa

[Ano52g] Anonymous. Publications received. Journal of the American Statistical Association, 47(257):111–112, March 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279995.

Anonymous:1952:PRb

[Ano52h] Anonymous. Publications received. Journal of the American Statistical Association, 47(258):352–353, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280775.

Anonymous:1952:PRc

[Ano52i] Anonymous. Publications received. Journal of the American Statistical Association, 47(259):581–582, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281340.

Anonymous:1952:PRd

[Ano52j] Anonymous. Publications received. Journal of the American Statistical Association, 47(260):709, December 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280797.

Anonymous:1952:RD

[Ano52k] Anonymous. Random digits (1-6000). Journal of the American Statistical Association, 47(260):710–714, December 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280798.

Anonymous:1952:VI

[Ano52l] Anonymous. Volume information. Journal of the American Statistical Association, 47(260):715–728, December 1952. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2280777.

Anonymous:1953:BMa

[Ano53a] Anonymous. Back matter. Journal of the American Statistical Association, 48(261):168, March 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280899.

Anonymous:1953:BMb

[Ano53b] Anonymous. Back matter. Journal of the American Statistical Association, 48(262):384, June 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281307.

Anonymous:1953:BNSa

[Ano53c] Anonymous. Bibliography of nonparametric statistics and related topics: Main bibliography. Journal of the American Statistical Association, 48(264):850–898, December 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281079.

Anonymous:1953:BNSb

[Ano53d] Anonymous. Bibliography of nonparametric statistics and related topics: Papers by annotation letter. Journal of the American Statistical Association, 48(264):898–906, December 1953. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281080.

Anonymous:1953:FMa

[Ano53e] Anonymous. Front matter. Journal of the American Statistical Association, 48(261):??, March 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280874.

Anonymous:1953:FMb

[Ano53f] Anonymous. Front matter. Journal of the American Statistical Association, 48(262):??, June 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281280.

Anonymous:1953:FMc

[Ano53g] Anonymous. Front matter. Journal of the American Statistical Association, 48(264):i–vi, December 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281065.

Anonymous:1953:PAS

[Ano53h] Anonymous. Proceedings: American Statistical Association 112th Annual Meeting, Palmer House, Chicago, Illinois, December 28, 1952. Journal of the American Statistical Association, 48(262):350– 357, June 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281292.

Anonymous:1953:PRa

[Ano53i] Anonymous. Publications received. Journal of the American Statistical Association, 48(261):166–167, March 1953. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280897.

Anonymous:1953:PRb

[Ano53j] Anonymous. Publications received. Journal of the American Statistical Association, 48(262):382–383, June 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281305.

Anonymous:1953:PRc

[Ano53k] Anonymous. Publications received. Journal of the American Statistical Association, 48(264):929–930, December 1953. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281093.

Anonymous:1953:RDa

[Ano53l] Anonymous. Random digits (6001-6100). Journal of the American Statistical Association, 48(261):167, March 1953. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2280898.

Anonymous:1953:RDb

[Ano53m] Anonymous. Random digits (6501-6875). Journal of the American Statistical Association, 48(262):383, June 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281306.

Anonymous:1953:RDc

[Ano53n] Anonymous. Random digits (6876-8125). Journal of the American Statistical Association, 48(263):672, September 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281029.

Anonymous:1953:RDd

[Ano53o] Anonymous. Random digits (9001-13,750). Journal of the American Statistical Association, 48(264):931–934, December 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281094.

Anonymous:1953:VI

[Ano53p] Anonymous. Volume information. Journal of the American Statistical Association, 48(264):vii–940, December 1953. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281064.

Anonymous:1954:BMa

[Ano54a] Anonymous. Back matter. Journal of the American Statistical Association, 49(265):208, March 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281063.

Anonymous:1954:BMb

[Ano54b] Anonymous. Back matter. Journal of the American Statistical Association, 49(266):412, June 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280959.

Anonymous:1954:FMa

[Ano54c] Anonymous. Front matter. Journal of the American Statistical Association, 49(265):??, March 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281030.

Anonymous:1954:FMb

[Ano54d] Anonymous. Front matter. Journal of the American Statistical Association, 49(266):??, June 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280929.

Anonymous:1954:FMc

[Ano54e] Anonymous. Front matter. Journal of the American Statistical Association, 49(267):??, September 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281120.

Anonymous:1954:FMd

[Ano54f] Anonymous. Front matter. Journal of the American Statistical Association, 49(268):i–vi, December 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281532.

Anonymous:1954:PAS

[Ano54g] Anonymous. Proceedings: American Statistical Association 113th Annual Meeting: Shoreham Hotel, Washington, D.C. December 28, 1953. Journal of the American Statistical Association, 49(266):315– 322, June 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280937.

Anonymous:1954:PRa

[Ano54h] Anonymous. Publications received. Journal of the American Statistical Association, 49(265):204–205, March 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281061.

Anonymous:1954:PRb

[Ano54i] Anonymous. Publications received. Journal of the American Statistical Association, 49(267):679–681, September 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281153.

Anonymous:1954:RDa

[Ano54j] Anonymous. Random digits (12,876-15,125). Journal of the American Statistical Association, 49(265):206–207, March 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281062.

Anonymous:1954:RDb

[Ano54k] Anonymous. Random digits (15,126-17,375). Journal of the American Statistical Association, 49(266):410–411, June 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280958.

Anonymous:1954:RDc

[Ano54l] Anonymous. Random digits (17,376-20,875). Journal of the American Statistical Association, 49(267):682–684, September 1954. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281154.

Anonymous:1954:RDd

[Ano54m] Anonymous. Random digits (20,876-21,875). Journal of the American Statistical Association, 49(268):928, December 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281560.

Anonymous:1954:SAa

[Ano54n] Anonymous. Statistical abstracts. Journal of the American Statistical Association, 49(265):178–183, March 1954. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281044.

Anonymous:1954:SAb

[Ano54o] Anonymous. Statistical abstracts. Journal of the American Statistical Association, 49(266):371–377, June 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280939.

Anonymous:1954:SAc

[Ano54p] Anonymous. Statistical abstracts. Journal of the American Statistical Association, 49(267):640–646, September 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281135.

Anonymous:1954:SAd

[Ano54q] Anonymous. Statistical abstracts. Journal of the American Statistical Association, 49(268):907–915, December 1954. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281553.

Anonymous:1954:U

[Ano54r] Anonymous. [untitled]. Journal of the American Statistical Association, 49(267):670, September 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281147.

Anonymous:1954:VI

[Ano54s] Anonymous. Volume information. Journal of the American Statistical Association, 49(268):929–934, December 1954. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281531.

Anonymous:1955:BMa

[Ano55a] Anonymous. Back matter. Journal of the American Statistical Association, 50(270):631–638, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280996.

Anonymous:1955:BMb

[Ano55b] Anonymous. Back matter. Journal of the American Statistical Association, 50(271):1004, September 1955. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281197.

Anonymous:1955:E

[Ano55c] Anonymous. Errata. Journal of the American Statistical Association, 50(272):1437, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281279.

Anonymous:1955:ESA

[Ano55d] Anonymous. Errata: Statistical abstracts. Journal of the American Statistical Association, 50(272):1331, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281221.

Anonymous:1955:FMa

[Ano55e] Anonymous. Front matter. Journal of the American Statistical Association, 50(269):??, March 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281095.

Anonymous:1955:FMb

[Ano55f] Anonymous. Front matter. Journal of the American Statistical Association, 50(270):??, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280960.

Anonymous:1955:FMc

[Ano55g] Anonymous. Front matter. Journal of the American Statistical Association, 50(271):??, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281155.

Anonymous:1955:FMd

[Ano55h] Anonymous. Front matter. Journal of the American Statistical Association, 50(272):i–vi, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281199.

Anonymous:1955:PRa

[Ano55i] Anonymous. Publications received. Journal of the American Statistical Association, 50(269):265–266, March 1955. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281119.

Anonymous:1955:PRb

[Ano55j] Anonymous. Publications received. Journal of the American Statistical Association, 50(270):630, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280995.

Anonymous:1955:PRc

[Ano55k] Anonymous. Publications received. Journal of the American Statistical Association, 50(272):1423–1425, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281278.

Anonymous:1955:SAa

[Ano55l] Anonymous. Statistical abstracts. Journal of the American Statistical Association, 50(269):249–256, March 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281110.

Anonymous:1955:SAb

[Ano55m] Anonymous. Statistical abstracts. Journal of the American Statistical Association, 50(271):968–973, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281179.

Anonymous:1955:U

[Ano55n] Anonymous. [untitled]. Journal of the American Statistical Association, 50(269):257, March 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281111.

Anonymous:1955:VI

[Ano55o] Anonymous. Volume information. Journal of the American Statistical Association, 50(272):1427–1436, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281198.

Anonymous:1956:BMa

[Ano56a] Anonymous. Back matter. Journal of the American Statistical Association, 51(273):208, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280838.

Anonymous:1956:BMb

[Ano56b] Anonymous. Back matter. Journal of the American Statistical Association, 51(274):423–428, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281391.

Anonymous:1956:BMc

[Ano56c] Anonymous. Back matter. Journal of the American Statistical Association, 51(275):564, September 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281465.

Anonymous:1956:FMa

[Ano56d] Anonymous. Front matter. Journal of the American Statistical Association, 51(273):??, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280799.

Anonymous:1956:FMb

[Ano56e] Anonymous. Front matter. Journal of the American Statistical Association, 51(274):??, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281341.

Anonymous:1956:FMc

[Ano56f] Anonymous. Front matter. Journal of the American Statistical Association, 51(275):??, September 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281430.

Anonymous:1956:FMd

[Ano56g] Anonymous. Front matter. Journal of the American Statistical Association, 51(276):??, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281467.

Anonymous:1956:PRa

[Ano56h] Anonymous. Publications received. Journal of the American Statistical Association, 51(273):206–207, March 1956. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280837.

Anonymous:1956:PRb

[Ano56i] Anonymous. Publications received. Journal of the American Statistical Association, 51(274):421–422, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281390.

Anonymous:1956:PRc

[Ano56j] Anonymous. Publications received. Journal of the American Statistical Association, 51(275):562–563, September 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281464.

Anonymous:1956:PRd

[Ano56k] Anonymous. Publications received. Journal of the American Statistical Association, 51(276):711–712, December 1956. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281530.

Anonymous:1956:SAa

[Ano56l] Anonymous. Statistical abstracts. Journal of the American Statistical Association, 51(273):159–162, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280811.

Anonymous:1956:SAb

[Ano56m] Anonymous. Statistical abstracts. Journal of the American Statistical Association, 51(274):370–374, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281352.

Anonymous:1956:SAc

[Ano56n] Anonymous. Statistical abstracts. Journal of the American Statistical Association, 51(275):528–534, September 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281440.

Anonymous:1956:SAd

[Ano56o] Anonymous. Statistical abstracts. Journal of the American Statistical Association, 51(276):653–656, December 1956. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281494.

Anonymous:1956:U

[Ano56p] Anonymous. [untitled]. Journal of the American Statistical Association, 51(274):407, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281378.

Anonymous:1956:VI

[Ano56q] Anonymous. Volume information. Journal of the American Statistical Association, 51(276):i–721, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281466.

Anonymous:1957:BMa

[Ano57a] Anonymous. Back matter. Journal of the American Statistical Association, 52(277):132, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281429.

Anonymous:1957:BMb

[Ano57b] Anonymous. Back matter. Journal of the American Statistical Association, 52(278):292–300, June 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280873.

Anonymous:1957:BMc

[Ano57c] Anonymous. Back matter. Journal of the American Statistical Association, 52(279):414, September 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280928.

Anonymous:1957:BMd

[Ano57d] Anonymous. Back matter. Journal of the American Statistical Association, 52(280):ii–v, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281737.

Anonymous:1957:FMa

[Ano57e] Anonymous. Front matter. Journal of the American Statistical Association, 52(277):??, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281392.

Anonymous:1957:FMb

[Ano57f] Anonymous. Front matter. Journal of the American Statistical Association, 52(278):??, June 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280839.

Anonymous:1957:FMc

[Ano57g] Anonymous. Front matter. Journal of the American Statistical Association, 52(279):??, September 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280900.

Anonymous:1957:FMd

[Ano57h] Anonymous. Front matter. Journal of the American Statistical Association, 52(280):iii–iv, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281691.

Anonymous:1957:PRa

[Ano57i] Anonymous. Publications received. Journal of the American Statistical Association, 52(277):130–131, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281428.

Anonymous:1957:PRb

[Ano57j] Anonymous. Publications received. Journal of the American Statistical Association, 52(278):290–291, June 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280872.

Anonymous:1957:PRc

[Ano57k] Anonymous. Publications received. Journal of the American Statistical Association, 52(279):412–413, September 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280927.

Anonymous:1957:PRd

[Ano57l] Anonymous. Publications received. Journal of the American Statistical Association, 52(280):615–616, December 1957. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281736.

Anonymous:1957:SAa

[Ano57m] Anonymous. Statistical abstracts. Journal of the American Statistical Association, 52(277):95–99, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281403.

Anonymous:1957:SAb

[Ano57n] Anonymous. Statistical abstracts. Journal of the American Statistical Association, 52(278):263–264, June 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280852.

Anonymous:1957:VI

[Ano57o] Anonymous. Volume information. Journal of the American Statistical Association, 52(280):617–i, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281690.

Anonymous:1958:BM

[Ano58a] Anonymous. Back matter. Journal of the American Statistical Association, 53(282):629–634, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281904.

Anonymous:1958:FM

[Ano58b] Anonymous. Front matter. Journal of the American Statistical Association, 53(284):iii–vii, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281950.

Anonymous:1958:NAAa

[Ano58c] Anonymous. Notes about authors. Journal of the American Statistical Association, 53(281):200–207, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282580.

Anonymous:1958:NAAb

[Ano58d] Anonymous. Notes about authors. Journal of the American Statistical Association, 53(282):558–564, June 1958. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281876.

Anonymous:1958:NAAc

[Ano58e] Anonymous. Notes about authors. Journal of the American Statistical Association, 53(283):748–750, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282053.

Anonymous:1958:NAAd

[Ano58f] Anonymous. Notes about authors. Journal of the American Statistical Association, 53(284):1026–1029, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281964.

Anonymous:1958:PRa

[Ano58g] Anonymous. Publications received. Journal of the American Statistical Association, 53(281):257–258, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282616.

Anonymous:1958:PRb

[Ano58h] Anonymous. Publications received. Journal of the American Statistical Association, 53(282):628, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281903.

Anonymous:1958:PRc

[Ano58i] Anonymous. Publications received. Journal of the American Statistical Association, 53(283):775–776, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282074.

Anonymous:1958:PRd

[Ano58j] Anonymous. Publications received. Journal of the American Statistical Association, 53(284):1055–1056, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281991.

Anonymous:1958:VI

[Ano58k] Anonymous. Volume information. Journal of the American Statistical Association, 53(284):i–1062, December 1958. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281949.

Anonymous:1959:BMa

[Ano59a] Anonymous. Back matter. Journal of the American Statistical Association, 54(286):528–534, June 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281801.

Anonymous:1959:BMb

[Ano59b] Anonymous. Back matter. Journal of the American Statistical Association, 54(287):716, September 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282559.

Anonymous:1959:BMc

[Ano59c] Anonymous. Back matter. Journal of the American Statistical Association, 54(288):??, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282532.

Anonymous:1959:FMa

[Ano59d] Anonymous. Front matter. Journal of the American Statistical Association, 54(285):iii–iv, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282134.

Anonymous:1959:FMb

[Ano59e] Anonymous. Front matter. Journal of the American Statistical Association, 54(286):iii–iv, June 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281775.

Anonymous:1959:FMc

[Ano59f] Anonymous. Front matter. Journal of the American Statistical Association, 54(287):iii–iv, September 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282533.

Anonymous:1959:FMd

[Ano59g] Anonymous. Front matter. Journal of the American Statistical Association, 54(288):iii–iv, December 1959. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282494.

Anonymous:1959:NAAa

[Ano59h] Anonymous. Notes about authors. Journal of the American Statistical Association, 54(285):310–314, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282154.

Anonymous:1959:NAAb

[Ano59i] Anonymous. Notes about authors. Journal of the American Statistical Association, 54(286):480–483, June 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281786.

Anonymous:1959:NAAc

[Ano59j] Anonymous. Notes about authors. Journal of the American Statistical Association, 54(287):695–698, September 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282548.

Anonymous:1959:NAAd

[Ano59k] Anonymous. Notes about authors. Journal of the American Statistical Association, 54(288):807–809, December 1959. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282505.

Anonymous:1959:PRa

[Ano59l] Anonymous. Publications received. Journal of the American Statistical Association, 54(285):334, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282174.

Anonymous:1959:PRb

[Ano59m] Anonymous. Publications received. Journal of the American Statistical Association, 54(286):526–527, June 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281800.

Anonymous:1959:PRc

[Ano59n] Anonymous. Publications received. Journal of the American Statistical Association, 54(287):712–715, September 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282558.

Anonymous:1959:PRd

[Ano59o] Anonymous. Publications received. Journal of the American Statistical Association, 54(288):829–830, December 1959. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282531.

Anonymous:1959:VI

[Ano59p] Anonymous. Volume information. Journal of the American Statistical Association, 54(288):i–834, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282493.

Anscombe:1956:BRB

[Ans56] F. J. Anscombe. Book review: [untitled]. Journal of the American Statistical Association, 51(276):657–659, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281495.

Anscombe:1958:RIC

[Ans58] F. J. Anscombe. Rectifying inspection of a continuous output. Journal of the American Statistical Association, 53(283):702–719, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2282048.

Anscombe:1959:BRB

[Ans59a] F. J. Anscombe. Book review: [untitled]. Journal of the American Statistical Association, 54(288):813–814, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282516.

Anscombe:1959:CRI

[Ans59b] F. J. Anscombe. Corrigenda: Rectifying inspection of a continuous output. Journal of the American Statistical Association, 54 (288):810, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282507.

Arant:1952:BRB

[Ara52] Willard D. Arant. Book review: [untitled]. Journal of the American Statistical Association, 47(257):103–104, March 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279989.

Arkin:1956:BRB

[Ark56] Herbert Arkin. Book review: [untitled]. Journal of the American Statistical Association, 51(276):663–664, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281501.

Arnold:1952:BRB

[Arn52] Kenneth J. Arnold. Book review: [untitled]. Journal of the American Statistical Association, 47(259):553–554, September 1952. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281325.

Arrow:1954:BRBa

[Arr54a] Kenneth J. Arrow. Book review: [untitled]. Journal of the American Statistical Association, 49(266):393–395, June 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280947.

Arrow:1954:BRBb

[Arr54b] Kenneth J. Arrow. Book review: [untitled]. Journal of the American Statistical Association, 49(267):660–661, September 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281141.

Arrow:1958:BRBa

[Arr58a] Kenneth J. Arrow. Book review: [untitled]. Journal of the American Statistical Association, 53(281):223–224, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282592.

Arrow:1958:BRBb

[Arr58b] Kenneth J. Arrow. Book review: [untitled]. Journal of the American Statistical Association, 53(281):230, March 1958. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282598.

Arrow:1958:TEP

[Arr58c] Kenneth J. Arrow. Tinbergen on economic policy. Journal of the American Statistical Association, 53(281):89–97, March 1958. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282568.

Askovitz:1957:SCG

[Ask57] S. I. Askovitz. A short-cut graphic method for fitting the best straight line to a series of points according to the criterion of least squares. Journal of the American Statistical Association, 52(277):13–17, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281394.

Askovitz:1959:GMB

[Ask59] S. I. Askovitz. Graphic methods based upon properties of advancing centroids. Journal of the American Statistical Association, 54 (287):668–673, September 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282544.

Atkinson:1955:BRB

[Atk55] Thomas R. Atkinson. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1417–1419, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281275.

B:1956:BRB

[B.56] Jorge Arias B. Book review: [untitled]. Journal of the American Statistical Association, 51(276):662–663, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281500.

B:1957:BRBa

[B.57a] Jorge Arias B. Book review: [untitled]. Journal of the American Statistical Association, 52(279):387–388, September 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280911.

B:1957:BRBb

[B.57b] Jorge Arias B. Book review: [untitled]. Journal of the American Statistical Association, 52(280):581–582, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281712.

B:1958:BRB

[B.58] Jorge Arias B. Book review: [untitled]. Journal of the American Statistical Association, 53(281):229–230, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282597.

Babbar:1955:DSS

[Bab55] M. M. Babbar. Distributions of solutions of a set of linear equations (with an application to linear programming). Journal of the American Statistical Association, 50(271):854–869, September 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281170.

Backman:1953:BRB

[Bac53] Jules Backman. Book review: [untitled]. Journal of the American Statistical Association, 48(261):156–157, March 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280892.

Backman:1958:BRB

[Bac58] Jules Backman. Book review: [untitled]. Journal of the American Statistical Association, 53(281):241–242, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282606.

Bahadur:1957:BRB

[Bah57] R. R. Bahadur. Book review: [untitled]. Journal of the American Statistical Association, 52(278):270–271, June 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280856.

Bailey:1958:BRB

[Bai58] Martin J. Bailey. Book review: [untitled]. Journal of the American Statistical Association, 53(281):231–233, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282600.

Baker:1956:BRB

[Bak56] G. A. Baker. Book review: [untitled]. Journal of the American Statistical Association, 51(274):387–388, June 1956. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281363.

Baker:1958:EIT

[Bak58] G. A. Baker. Empiric investigation of a test of homogeneity for populations composed of normal distributions. Journal of the American Statistical Association, 53(282):551–557, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281875.

Bales:1951:SSP

[Bal51] Robert F. Bales. Some statistical problems in small group research. Journal of the American Statistical Association, 46(255):311–322, September 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280509.

Balderston:1956:BRB

[Bal56] F. E. Balderston. Book review: [untitled]. Journal of the American Statistical Association, 51(273):180–182, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280821.

Bancroft:1950:PVC

[Ban50] T. A. Bancroft. Probability values for the common tests of hypotheses. Journal of the American Statistical Association, 45(250):211– 217, June 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280678.

Bancroft:1958:BRB

[Ban58] T. A. Bancroft. Book review: [untitled]. Journal of the American Statistical Association, 53(282):598–599, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281879.

Bainbridge:1956:CTA

[BAR56] J. R. Bainbridge, M. Alison, and U. Radok. Corrigenda: Tabular analysis of factorial experiments and the use of punch cards. Journal of the American Statistical Association, 51(276):650, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2281479.

Bartholomew:1957:PLT

[Bar57] D. J. Bartholomew. A problem in life testing. Journal of the American Statistical Association, 52(279):350–355, September 1957. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280905.

Bargmann:1958:BRB

[Bar58] Rolf E. Bargmann. Book review: [untitled]. Journal of the American Statistical Association, 53(283):753–755, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282056.

Bush:1956:BRBc

[BC56] Robert R. Bush and Bernard P. Cohen. Book review: [untitled]. Journal of the American Statistical Association, 51(276):702–704, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281524.

Baldwin:1956:OSB

[BCMM56] Roger R. Baldwin, Wilbert E. Cantey, Herbert Maisel, and James P. McDermott. The optimum strategy in blackjack. Journal of the American Statistical Association, 51(275):429–439, September 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2281431.

Baldwin:1959:COS

[BCMM59] Roger R. Baldwin, Wilbert E. Cantey, Herbert Maisel, and James P. McDermott. Corrigenda: The optimum strategy in blackjack. Journal of the American Statistical Association, 54(288):810, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2282508.

Box:1959:BSR

[BD59] G. E. P. Box and Norman R. Draper. A basis for the selection of a response surface design. Journal of the American Statistical Association, 54(287):622–654, September 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282542.

Bauer:1954:DPL

[BDL^{+}54] Theodore J. Bauer, James F. Donohue, Vincent Larsen, Albert P. Iskrant, and Quentin R. Remein. Do persons lost to long term observation have the same experience as persons observed?: Evaluation of antisyphilitic therapy. Journal of the American Statistical Association, 49(265):36–50, March 1954. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281033.

Bean:1950:BRB

[Bea50] Louis H. Bean. Book review: [untitled]. Journal of the American Statistical Association, 45(251):461–464, September 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280305.

Beall:1954:BRB

[Bea54] Geoffrey Beall. Book review: [untitled]. Journal of the American Statistical Association, 49(266):401–402, June 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280952.

Beach:1957:BRB

[Bea57] E. F. Beach. Book review: [untitled]. Journal of the American Statistical Association, 52(277):102–103, March 1957. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281406.

Beattie:1959:SSC

[Bea59] Ronald H. Beattie. Sources of statistics on crime and correction. Journal of the American Statistical Association, 54(287):582–592, September 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2282538.

Belloc:1950:LFP

[Bel50] Nedra Bartlett Belloc. Labor-force participation and employment opportunities for women. Journal of the American Statistical Association, 45(251):400–410, September 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280295.

Belloc:1954:VMS

[Bel54] Nedra B. Belloc. Validation of morbidity survey data by comparison with hospital records. Journal of the American Statistical Association, 49(268):832–846, December 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281543.

Bellman:1955:BRB

[Bel55] Richard Bellman. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1357–1359, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281240.

Benjamin:1950:BRB

[Ben50a] Harold Benjamin. Book review: [untitled]. Journal of the American Statistical Association, 45(251):470, September 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280310.

Bennett:1950:BRB

[Ben50b] Rollin F. Bennett. Book review: [untitled]. Journal of the American Statistical Association, 45(251):469, September 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280309.

Bennett:1955:BRB

[Ben55] J. H. Bennett. Book review: [untitled]. Journal of the American Statistical Association, 50(271):979–980, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281182.

Benewitz:1957:BRB

[Ben57] Maurice C. Benewitz. Book review: [untitled]. Journal of the American Statistical Association, 52(279):405–406, September 1957. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280922.

Berkson:1950:TTR

[Ber50] Joseph Berkson. Are there two regressions? Journal of the American Statistical Association, 45(250):164–180, June 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280676.

Berkson:1953:SPR

[Ber53a] Joseph Berkson. A statistically precise and relatively simple method of estimating the bioassay with quantal response, based on the logistic function. Journal of the American Statistical Association, 48 (263):565–599, September 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor. org/stable/2281010. See comments [Ber61].

Berle:1953:WCM

[Ber53b] Adolf A. Berle, Jr. Wesley Clair Mitchell: The economic scientist. Journal of the American Statistical Association, 48(262):169–175, June 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2281281.

Berkson:1955:EIN

[Ber55a] Joseph Berkson. Estimate of the integrated normal curve by minimum normit chi-square with particular reference to bio-assay. Journal of the American Statistical Association, 50(270):529–549, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280974.

Berkson:1955:MLM

[Ber55b] Joseph Berkson. Maximum likelihood and minimum χ^{2} estimates of the logistic function. Journal of the American Statistical Association, 50(269):130–162, March 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281102.

Bernstein:1956:BRB

[Ber56] Peter L. Bernstein. Book review: [untitled]. Journal of the American Statistical Association, 51(275):558–560, September 1956. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281461.

Berndt:1957:RUP

[Ber57] Gerald D. Berndt. The regions of unimodality and positivity in the abbreviated Edgeworth and Gram–Charlier series. Journal of the American Statistical Association, 52(278):253–256, June 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280850.

Berkson:1958:SLC

[Ber58] Joseph Berkson. Smoking and lung cancer: Some observations on two recent reports. Journal of the American Statistical Association, 53(281):28–38, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282563.

Berkson:1961:OSL

[Ber61] Joseph Berkson. The other side of the lower bound. A note with a correction. Journal of the American Statistical Association, 56 (295):670–674, September 1961. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282087. See [Ber53a].

Beyer:1952:BRB

[Bey52] Glenn H. Beyer. Book review: [untitled]. Journal of the American Statistical Association, 47(259):571–573, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281335.

Brunk:1953:EDP

[BF53] Max E. Brunk and Walter T. Federer. Experimental designs and probability sampling in marketing research. Journal of the American Statistical Association, 48(263):440–452, September 1953. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281000.

Berkson:1952:SCC

[BG52] Joseph Berkson and Robert P. Gage. Survival curve for cancer patients following treatment. Journal of the American Statistical Association, 47(259):501–515, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281318.

Bainbridge:1956:TAF

[BGR56] J. R. Bainbridge, Alison M. Grant, and U. Radok. Tabular analysis of factorial experiments and the use of punch cards. Journal of the American Statistical Association, 51(273):149–158, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280810.

Belz:1954:ADR

[BH54] Maurice H. Belz and Robert Hooke. Approximate distribution of the range in the neighborhood of low percentage points. Journal of the American Statistical Association, 49(267):620–636, September 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2281133.

Brown:1953:ECE

[BHP53a] J. A. C. Brown, H. S. Houthakker, and S. J. Prais. Electronic computation in economic statistics. Journal of the American Statistical Association, 48(263):414–428, September 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280998.

Brown:1953:EEC

[BHP53b] J. A. C. Brown, H. S. Houthakker, and S. J. Prais. Errata: Electronic computation in economic statistics. Journal of the American Statistical Association, 48(264):912, December 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281086.

Brownlee:1953:MSS

[BHR53] K. A. Brownlee, J. L. Hodges, Jr., and Murray Rosenblatt. The up-and-down method with small samples. Journal of the American Statistical Association, 48(262):262–277, June 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281287.

Birnbaum:1952:NTD

[Bir52] Z. W. Birnbaum. Numerical tabulation of the distribution of Kolmogorov’s statistic for finite sample size. Journal of the American Statistical Association, 47(259):425–441, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281313.

Birnbaum:1954:CIT

[Bir54a] Allan Birnbaum. Combining independent tests of significance. Journal of the American Statistical Association, 49(267):559–574, September 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281130.

Birnbaum:1954:SMP

[Bir54b] Allan Birnbaum. Statistical methods for Poisson processes and exponential populations. Journal of the American Statistical Association, 49(266):254–266, June 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor. org/stable/2280934. See corrigenda [Bir66].

Birnbaum:1955:BRB

[Bir55] Z. W. Birnbaum. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1379–1382, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281254.

Birnbaum:1956:BRB

[Bir56] Allan Birnbaum. Book review: [untitled]. Journal of the American Statistical Association, 51(274):385–386, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281361.

Birnbaum:1966:CSM

[Bir66] Allan Birnbaum. Corrigenda: Statistical Methods for Poisson Processes and Exponential Populations. Journal of the American Statistical Association, 61(316):1246, December 1966. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2283225. See [Bir54b].

Barger:1954:QMU

[BK54] Harold Barger and Lawrence R. Klein. A quarterly model for the United States economy. Journal of the American Statistical Association, 49(267):413–437, September 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281121.

Bross:1957:RAT

[BK57] Irwin D. J. Bross and Ethel L. Kasten. Rapid analysis of 2 —times 2 tables. Journal of the American Statistical Association, 52(277):18– 28, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281395.

Blackwell:1951:BRB

[Bla51] David Blackwell. Book review: [untitled]. Journal of the American Statistical Association, 46(255):381–382, September 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280517.

Blank:1954:RBI

[Bla54] David M. Blank. Relationship between an index of house prices and building costs. Journal of the American Statistical Association, 49(265):67–78, March 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281035.

Blank:1956:BRB

[Bla56] David M. Blank. Book review: [untitled]. Journal of the American Statistical Association, 51(276):688–689, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281515.

Blalock:1958:CPI

[Bla58a] H. M. Blalock, Jr. Corrigenda: Probabilistic interpretations for the mean square contigency. Journal of the American Statistical Association, 53(284):1030, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281965.

Blalock:1958:PIM

[Bla58b] H. M. Blalock, Jr. Probabilistic interpretations for the mean square contingency. Journal of the American Statistical Association, 53 (281):102–105, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282570.

Blanch:1958:BRB

[Bla58c] Gertrude Blanch. Book review: [untitled]. Journal of the American Statistical Association, 53(281):218–219, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282587.

Blenko:1958:BRB

[Ble58] Walter J. Blenko. Book review: [untitled]. Journal of the American Statistical Association, 53(284):1037–1038, December 1958. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281977.

Blumen:1954:BRB

[Blu54] Isadore Blumen. Book review: [untitled]. Journal of the American Statistical Association, 49(265):196–197, March 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281053.

Blumberg:1958:BRB

[Blu58a] Mark Blumberg. Book review: [untitled]. Journal of the American Statistical Association, 53(284):1038, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281978.

Blumen:1958:NBS

[Blu58b] Isadore Blumen. A new bivariate sign test. Journal of the American Statistical Association, 53(282):448–456, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281867.

Bain:1953:BRB

[BO53] J. C. Bain and Paul S. Olmstead. Book review: [untitled]. Journal of the American Statistical Association, 48(263):659–661, September 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2281021.

Boddy:1957:BRB

[Bod57] Francis M. Boddy. Book review: [untitled]. Journal of the American Statistical Association, 52(278):280–281, June 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280865.

Boddy:1959:BRB

[Bod59] Francis M. Boddy. Book review: [untitled]. Journal of the American Statistical Association, 54(286):515–516, June 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281792.

Bogue:1950:TME

[Bog50] Donald J. Bogue. A technique for making extensive population estimates. Journal of the American Statistical Association, 45 (250):149–163, June 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280675.

Bohr:1955:BRB

[Boh55] Kenneth A. Bohr. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1400–1401, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281266.

Borenstein:1957:BRB

[Bor57] Israel Borenstein. Book review: [untitled]. Journal of the American Statistical Association, 52(279):401–405, September 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280921.

Boschan:1958:BRB

[Bos58] Paul Boschan. Book review: [untitled]. Journal of the American Statistical Association, 53(281):235–237, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282602.

Boulding:1959:BRB

[Bou59a] Kenneth E. Boulding. Book review: [untitled]. Journal of the American Statistical Association, 54(287):704–705, September 1959. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282552.

Bourque:1959:BRB

[Bou59b] Philip Bourque. Book review: [untitled]. Journal of the American Statistical Association, 54(285):318–319, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282159.

Bowker:1950:BRB

[Bow50a] Albert H. Bowker. Book review: [untitled]. Journal of the American Statistical Association, 45(252):574–575, December 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280730.

Bowman:1950:BRB

[Bow50b] Jean Claire Bowman. Book review: [untitled]. Journal of the American Statistical Association, 45(251):471–473, September 1950. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280311.

Bowman:1956:BRB

[Bow56] Ward S. Bowman, Jr. Book review: [untitled]. Journal of the American Statistical Association, 51(273):184–188, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280824.

Box:1955:BRB

[Box55] George E. P. Box. Book review: [untitled]. Journal of the American Statistical Association, 50(271):974–975, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281180.

Barnes:1955:AVG

[BPR55] Benjamin A. Barnes, Elinor Pearson, and Eric Reiss. The analysis of variance: a graphical representation of a statistical concept. Journal of the American Statistical Association, 50(272):1064–1072, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281205.

Brady:1951:BRBa

[Bra51a] Dorothy S. Brady. Book review: [untitled]. Journal of the American Statistical Association, 46(254):267–269, June 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279933.

Brady:1951:BRBb

[Bra51b] Dorothy S. Brady. Book review: [untitled]. Journal of the American Statistical Association, 46(254):269–270, June 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279934.

Brady:1954:BRB

[Bra54a] Dorothy S. Brady. Book review: [untitled]. Journal of the American Statistical Association, 49(266):395–399, June 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280948.

Brady:1954:KRF

[Bra54b] Dorothy S. Brady. The Kinsey Report on Females. Journal of the American Statistical Association, 49(268):696–705, December 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281534.

Brayfield:1955:BRB

[Bra55] Arthur H. Brayfield. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1377–1379, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281253.

Bradt:1956:BRB

[Bra56a] R. N. Bradt. Book review: [untitled]. Journal of the American Statistical Association, 51(273):173–175, March 1956. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2280817.

Braga:1956:BRB

[Bra56b] Fred W. Braga. Book review: [untitled]. Journal of the American Statistical Association, 51(276):685–686, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281513.

Bradley:1958:CCI

[Bra58a] James V. Bradley. Complete counterbalancing of immediate sequential effects in a Latin square design. Journal of the American Statistical Association, 53(282):525–528, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281872.

Bradley:1958:CCC

[Bra58b] James V. Bradley. Corrigenda: Complete counterbalancing of immediate sequential effects in a Latin square design. Journal of the American Statistical Association, 53(284):1030–1031, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2281966.

Bratt:1958:BRB

[Bra58c] Elmer C. Bratt. Book review: [untitled]. Journal of the American Statistical Association, 53(281):230–231, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282599.

Bratt:1959:BRBa

[Bra59a] Elmer C. Bratt. Book review: [untitled]. Journal of the American Statistical Association, 54(288):823, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282525.

Bratt:1959:BRBb

[Bra59b] Elmer C. Bratt. Book review: [untitled]. Journal of the American Statistical Association, 54(288):824–825, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282527.

Breakwell:1956:EOA

[Bre56] John V. Breakwell. Economically optimum acceptance tests. Journal of the American Statistical Association, 51(274):243–256, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2281344.

Bremer:1958:BRB

[Bre58] Hans O. Bremer. Book review: [untitled]. Journal of the American Statistical Association, 53(282):600, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281881.

Brickman:1950:BRB

[Bri50] Myrtle Brickman. Book review: [untitled]. Journal of the American Statistical Association, 45(250):287–288, June 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280687.

Brinegar:1956:BRB

[Bri56] Claude S. Brinegar. Book review: [untitled]. Journal of the American Statistical Association, 51(273):166–169, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280814.

Brimmer:1959:BRB

[Bri59] Andrew F. Brimmer. Book review: [untitled]. Journal of the American Statistical Association, 54(288):823–824, December 1959. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282526.

Bross:1950:TCS

[Bro50] Irwin Bross. Two-choice selection. Journal of the American Statistical Association, 45(252):530–540, December 1950. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2280721.

Brownlee:1952:BRBb

[Bro52a] K. A. Brownlee. Book review: Probit Analysis: A Statistical Treatment of the Sigmoid Response Curve, by D. J. Finney; F.

Tattersfield. Journal of the American Statistical Association, 47 (260):687–691, December 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280787.

Brownlee:1952:BRBa

[Bro52b] K. A. Brownlee. Book review: [untitled]. Journal of the American Statistical Association, 47(258):321–325, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280759.

Broida:1953:BRB

[Bro53a] Arthur L. Broida. Book review: [untitled]. Journal of the American Statistical Association, 48(264):924–928, December 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281092.

Bross:1953:BRB

[Bro53b] Irwin Bross. Book review: [untitled]. Journal of the American Statistical Association, 48(263):655–657, September 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281019.

Brooks:1955:EOS

[Bro55a] Samuel H. Brooks. The estimation of an optimum subsampling number. Journal of the American Statistical Association, 50(270): 398–415, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2280969.

Brown:1955:TEP

[Bro55b] Hugh H. Brown. A technique for estimating the population of counties. Journal of the American Statistical Association, 50(270):323– 343, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280964.

Brownlee:1955:SPV

[Bro55c] K. A. Brownlee. Statistics of the 1954 polio vaccine trials. Journal of the American Statistical Association, 50(272):1005–1013, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2281200.

Brower:1956:BRB

[Bro56a] E. J. Brower. Book review: [untitled]. Journal of the American Statistical Association, 51(274):412–413, June 1956. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281383.

Brown:1956:BRB

[Bro56b] George H. Brown. Book review: [untitled]. Journal of the American Statistical Association, 51(274):402–404, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281374.

Brownlee:1957:BRB

[Bro57a] K. A. Brownlee. Book review: [untitled]. Journal of the American Statistical Association, 52(277):100–102, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281404.

Brownlee:1957:NEN

[Bro57b] K. A. Brownlee. A note on the effects of nonresponse on surveys. Journal of the American Statistical Association, 52(277):29– 32, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281396.

Brodbeck:1958:BRB

[Bro58a] May Brodbeck. Book review: [untitled]. Journal of the American Statistical Association, 53(282):600–603, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281884.

Bronfenbrenner:1958:BRB

[Bro58b] M. Bronfenbrenner. Book review: [untitled]. Journal of the American Statistical Association, 53(282):603–604, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281885.

Brownlee:1958:BRB

[Bro58c] K. A. Brownlee. Book review: [untitled]. Journal of the American Statistical Association, 53(281):222–223, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282591.

Brooks:1959:BRB

[Bro59] Samuel H. Brooks. Book review: [untitled]. Journal of the American Statistical Association, 54(287):710, September 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282555.

Brunsman:1955:BRB

[Bru55] Howard G. Brunsman. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1385–1386, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2281257.

Bry:1957:BRB

[Bry57] Gerhard Bry. Book review: [untitled]. Journal of the American Statistical Association, 52(280):597–598, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281726.

Birnbaum:1950:BDN

[BS50] Z. W. Birnbaum and Monroe G. Sirken. Bias due to non-availability in sampling surveys. Journal of the American Statistical Association, 45(249):98–111, March 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280431.

Bose:1952:CAP

[BS52] R. C. Bose and T. Shimamoto. Classification and analysis of partially balanced incomplete block designs with two associate classes. Journal of the American Statistical Association, 47(258):151–184, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280741.

Birnbaum:1958:SML

[BS58] Z. W. Birnbaum and S. C. Saunders. A statistical model for lifelength of materials. Journal of the American Statistical Association, 53(281):151–160, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282575.

Buckland:1954:BRB

[Buc54] William R. Buckland. Book review: [untitled]. Journal of the American Statistical Association, 49(265):186–189, March 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281047.

Budd:1956:BRB

[Bud56] Edward C. Budd. Book review: [untitled]. Journal of the American Statistical Association, 51(276):676–677, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281510.

Buechley:1956:BRB

[Bue56] Robert W. Buechley. Book review: Social Area Analysis, by Eshref Shevky; Wendell Bell. Journal of the American Statistical Association, 51(273):195–197, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280830.

Buehler:1957:CIP

[Bue57] Robert J. Buehler. Confidence intervals for the product of two binomial parameters. Journal of the American Statistical Association, 52(280):482–493, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281697.

Burkhead:1953:CFD

[Bur53] Jesse Burkhead. Changes in the functional distribution of income. Journal of the American Statistical Association, 48(262):192–219, June 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2281283.

Burr:1954:UET

[Bur54] Irving W. Burr. Use of experiments in teaching engineering statistics. Journal of the American Statistical Association, 49(268): 897–905, December 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281548.

Burgess:1955:BRB

[Bur55] Ernest W. Burgess. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1376–1377, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281252.

Burk:1958:SAI

[Bur58] Marguerite C. Burk. Some analyses of income-food relationships. Journal of the American Statistical Association, 53(284):905–927, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281960.

Bush:1956:BRBa

[Bus56a] Robert R. Bush. Book review: [untitled]. Journal of the American Statistical Association, 51(274):414–416, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281385.

Bush:1956:BRBb

[Bus56b] Robert R. Bush. Book review: [untitled]. Journal of the American Statistical Association, 51(276):661–662, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281499.

Cameron:1957:BRB

[Cam57] Joseph M. Cameron. Book review: [untitled]. Journal of the American Statistical Association, 52(279):389–390, September 1957. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280913.

Canning:1950:BRB

[Can50] John B. Canning. Book review: [untitled]. Journal of the American Statistical Association, 45(252):581–582, December 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280734.

Carter:1953:INM

[Car53] Hugh Carter. Improving national marriage and divorce statistics. Journal of the American Statistical Association, 48(263):453–461, September 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281001.

Chung:1958:RTM

[CF58] J. H. Chung and D. A. S. Fraser. Randomization tests for a multivariate two-sample problem. Journal of the American Statistical Association, 53(283):729–735, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282050.

Coulter:1959:AVS

[CG59] Elizabeth J. Coulter and Lillian Guralnick. Analysis of vital statistics by census tract. Journal of the American Statistical Association, 54(288):730–740, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282496.

Chapin:1951:BRB

[Cha51] F. Stuart Chapin. Book review: [untitled]. Journal of the American Statistical Association, 46(256):530–532, December 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280414.

Chapman:1952:BRB

[Cha52] Douglas G. Chapman. Book review: [untitled]. Journal of the American Statistical Association, 47(260):686–687, December 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280786.

Chernick:1955:BRB

[Che55a] Jack Chernick. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1388–1389, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281259.

Chernoff:1955:BRB

[Che55b] Herman Chernoff. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1334–1335, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281228.

Chernoff:1956:CET

[Che56] Herman Chernoff. Corrigenda: Exercises in theoretical statistics. Journal of the American Statistical Association, 51(276):650, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281480.

Chipman:1953:BRB

[Chi53] John S. Chipman. Book review: [untitled]. Journal of the American Statistical Association, 48(264):917–923, December 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281090.

Chin:1954:BRB

[Chi54] Rockwood Chin. Book review: [untitled]. Journal of the American Statistical Association, 49(267):677–678, September 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281152.

Chinitz:1959:BRB

[Chi59] Benjamin Chinitz. Book review: [untitled]. Journal of the American Statistical Association, 54(286):521–522, June 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281796.

Choynowski:1959:MBP

[Cho59] Mieczyslaw Choynowski. Maps based on probabilities. Journal of the American Statistical Association, 54(286):385–388, June 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281779.

Chung:1950:SSF

[Chu50] J. H. Chung. Sequential sampling from finite lots when the proportion defective is small. Journal of the American Statistical Association, 45(252):557–569, December 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280724.

Churchman:1951:BRB

[Chu51] C. West Churchman. Book review: [untitled]. Journal of the American Statistical Association, 46(254):261, June 1951. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2279930.

Chung:1957:BRB

[Chu57] K. L. Chung. Book review: [untitled]. Journal of the American Statistical Association, 52(280):589–590, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281718.

Chernoff:1954:UNP

[CL54a] Herman Chernoff and Gerald J. Lieberman. Use of normal probability paper. Journal of the American Statistical Association, 49 (268):778–785, December 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281539.

Cohen:1954:REC

[CL54b] Samuel E. Cohen and Benjamin Lipstein. Response errors in the collection of wage statistics by mail questionnaire. Journal of the American Statistical Association, 49(266):240–250, June 1954. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280932.

Clark:1950:BRB

[Cla50] Andrew G. Clark. Book review: [untitled]. Journal of the American Statistical Association, 45(252):570–571, December 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280726.

Clark:1952:BRB

[Cla52] Kenneth E. Clark. Book review: [untitled]. Journal of the American Statistical Association, 47(258):325–327, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280760.

Clark:1953:BRB

[Cla53a] Colin Clark. Book review: [untitled]. Journal of the American Statistical Association, 48(262):374–376, June 1953. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281302.

Clark:1953:PPI

[Cla53b] Robert E. Clark. Percentage points of the incomplete beta function. Journal of the American Statistical Association, 48(264):831–843, December 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281077.

Clark:1957:TMR

[Cla57a] Frank Eugene Clark. Truncation to meet requirements on means. Journal of the American Statistical Association, 52(280):527–536, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281702.

Clark:1957:BRB

[Cla57b] Paul G. Clark. Book review: [untitled]. Journal of the American Statistical Association, 52(279):406–407, September 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280923.

Clawson:1959:SDA

[Cla59] Marion Clawson. Statistical data available for economic research on certain types of recreation. Journal of the American Statistical Association, 54(285):281–309, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282153.

Cornfield:1950:CSN

[CM50a] Jerome Cornfield and Nathan Mantel. Correction to “Some New Aspects of the Application of Maximum Likelihood to the Calculation of the Dosage Response Curve”. Journal of the American Statistical Association, 45(252):569, December 1950. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2280725.

Cornfield:1950:SNA

[CM50b] Jerome Cornfield and Nathan Mantel. Some new aspects of the application of maximum likelihood to the calculation of the dosage response curve. Journal of the American Statistical Association, 45(250):181–210, June 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280677.

Cochran:1953:SPK

[CMT53] William G. Cochran, Frederick Mosteller, and John W. Tukey. Statistical problems of the Kinsey report. Journal of the American Statistical Association, 48(264):673–716, December 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281066.

Cochran:1954:PS

[CMT54] William G. Cochran, Frederick Mosteller, and John W. Tukey. Principles of sampling. Journal of the American Statistical Association, 49(265):13–35, March 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281032.

Cochran:1955:CMT

[CMT55] W. G. Cochran, F. Mosteller, and John W. Tukey. The Cochran– Mosteller–Tukey report on the Kinsey Study: A symposium (pp. 811–829). Journal of the American Statistical Association, 50(271): 829, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Coale:1955:BRB

[Coa55a] Ansley J. Coale. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1372–1373, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281249.

Coale:1955:CPU

[Coa55b] Ansley J. Coale. Corrigenda: The population of the United States in 1950 classified by age, sex, and color — a revision of census figures. Journal of the American Statistical Association, 50(272):1331, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281219.

Coale:1955:PUS

[Coa55c] Ansley J. Coale. The population of the United States in 1950 classified by age, sex, and color — a revision of census figures. Journal of the American Statistical Association, 50(269):16–54, March 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281097.

Cochran:1952:BRB

[Coc52] William G. Cochran. Book review: [untitled]. Journal of the American Statistical Association, 47(257):93–94, March 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279982.

Cochran:1954:BRB

[Coc54a] William G. Cochran. Book review: [untitled]. Journal of the American Statistical Association, 49(265):200–201, March 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281055.

Cochran:1954:PSA

[Coc54b] William G. Cochran. The present structure of the association. Journal of the American Statistical Association, 49(265):1–12, March 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2281031.

Cochran:1955:TLF

[Coc55] William G. Cochran. A test of a linear function of the deviations between observed and expected numbers. Journal of the American Statistical Association, 50(270):377–397, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280968.

Cochran:1956:BRB

[Coc56] William G. Cochran. Book review: [untitled]. Journal of the American Statistical Association, 51(276):664–666, December 1956. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281502.

Cochran:1957:BRB

[Coc57] William G. Cochran. Book review: [untitled]. Journal of the American Statistical Association, 52(279):386, September 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280909.

Coe:1955:CNP

[Coe55a] Paul F. Coe. Corrigenda: Nonwhite population increases in metropolitan areas. Journal of the American Statistical Association, 50(272):1331, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281220.

Coe:1955:NPI

[Coe55b] Paul F. Coe. Nonwhite population increases in metropolitan areas. Journal of the American Statistical Association, 50(270):283–308, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280962.

Cohen:1950:EPP

[Coh50] A. C. Cohen, Jr. Estimating parameters of Pearson Type III populations from truncated samples. Journal of the American Statistical Association, 45(251):411–423, September 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280296.

Cohen:1951:EPL

[Coh51] A. C. Cohen, Jr. Estimating parameters of logarithmic-normal distributions by maximum likelihood. Journal of the American Statistical Association, 46(254):206–212, June 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279920.

Cohen:1954:EPP

[Coh54] A. C. Cohen, Jr. Estimation of the Poisson parameter from truncated samples and from consored samples. Journal of the American Statistical Association, 49(265):158–168, March 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281042.

Cohen:1955:MLE

[Coh55a] A. Clifford Cohen, Jr. Maximum likelihood estimation of the dispersion parameter of a chi-distributed radial error from truncated and censored samples with applications to target analysis. Journal of the American Statistical Association, 50(272):1122–1135, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2281209.

Cohen:1955:RSS

[Coh55b] A. Clifford Cohen, Jr. Restriction and selection in samples from bivariate normal distributions. Journal of the American Statistical Association, 50(271):884–893, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281173.

Cohen:1956:BRB

[Coh56] Jerome B. Cohen. Book review: [untitled]. Journal of the American Statistical Association, 51(274):375, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281353.

Cohen:1957:BRB

[Coh57] Jerome B. Cohen. Book review: [untitled]. Journal of the American Statistical Association, 52(280):603–604, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281731.

Cohen:1959:TST

[Coh59a] Arthur Cohen. Tables for the sign test when observations are estimates of binomial parameters. Journal of the American Statistical Association, 54(288):784–793, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282502.

Cohn:1959:PEF

[Coh59b] Samuel M. Cohn. Problems in estimating Federal Government expenditures. Journal of the American Statistical Association, 54 (288):717–729, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282495.

Coleman:1956:BRB

[Col56] James S. Coleman. Book review: [untitled]. Journal of the American Statistical Association, 51(276):701–702, December 1956. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281523.

Colean:1957:BRB

[Col57] Miles L. Colean. Book review: [untitled]. Journal of the American Statistical Association, 52(280):594–595, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281724.

Combs:1959:BRB

[Com59] Jerry W. Combs, Jr. Book review: [untitled]. Journal of the American Statistical Association, 54(285):327–329, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282167.

Connor:1952:BRB

[Con52] W. S. Connor, Jr. Book review: [untitled]. Journal of the American Statistical Association, 47(257):108–109, March 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279992.

Connor:1956:BRB

[Con56] W. S. Connor. Book review: [untitled]. Journal of the American Statistical Association, 51(274):386–387, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281362.

Coombs:1951:MMP

[Coo51a] C. H. Coombs. Mathematical models in psychological scaling. Journal of the American Statistical Association, 46(256):480–489, December 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280397.

Cooper:1951:BRB

[Coo51b] M. R. Cooper. Book review: [untitled]. Journal of the American Statistical Association, 46(254):273–274, June 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279937.

Cooper:1958:BRB

[Coo58] W. W. Cooper. Book review: [untitled]. Journal of the American Statistical Association, 53(282):621–624, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281898.

Copeland:1955:SOE

[Cop55] Morris A. Copeland. Statistics and objective economics. Journal of the American Statistical Association, 50(271):639–659, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2281156.

Cornell:1950:BRB

[Cor50] F. G. Cornell. Book review: [untitled]. Journal of the American Statistical Association, 45(252):583, December 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280735.

Cornfield:1951:BRB

[Cor51] Jerome Cornfield. Book review: [untitled]. Journal of the American Statistical Association, 46(256):525–526, December 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280410.

Cornfield:1955:BRB

[Cor55] Jerome Cornfield. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1368–1371, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281247.

Cornfield:1958:BRB

[Cor58] Jerome Cornfield. Book review: [untitled]. Journal of the American Statistical Association, 53(281):253–254, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282614.

Cover:1951:BRB

[Cov51] John H. Cover. Book review: [untitled]. Journal of the American Statistical Association, 46(253):136–137, March 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280110.

Cowden:1952:MPC

[Cow52] Dudley J. Cowden. The multiple-partial correlation coefficient. Journal of the American Statistical Association, 47(259):442–456, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281314.

Cowden:1958:PCR

[Cow58] Dudley J. Cowden. A procedure for computing regression coefficients. Journal of the American Statistical Association, 53(281): 144–150, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2282574.

Cowden:1959:CPC

[Cow59] Dudley J. Cowden. Corrigenda: a procedure for computing regression coefficients. Journal of the American Statistical Association, 54(288):811, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282509.

Cox:1957:NG

[Cox57a] D. R. Cox. Note on grouping. Journal of the American Statistical Association, 52(280):543–547, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281704.

Cox:1957:SF

[Cox57b] Gertrude M. Cox. Statistical frontiers. Journal of the American Statistical Association, 52(277):1–12, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281393.

Cox:1959:BRB

[Cox59] D. R. Cox. Book review: [untitled]. Journal of the American Statistical Association, 54(285):332–333, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282172.

Carter:1956:TAP

[CR56] Charles F. Carter and Mary Robson. A test of the accuracy of a production index. Journal of the American Statistical Association, 51(273):17–23, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280802.

Craig:1950:BRB

[Cra50] C. C. Craig. Book review: [untitled]. Journal of the American Statistical Association, 45(250):303–307, June 1950. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2280697.

Craig:1953:BRB

[Cra53a] C. C. Craig. Book review: [untitled]. Journal of the American Statistical Association, 48(262):363–364, June 1953. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281296.

Craig:1953:CNC

[Cra53b] C. C. Craig. Combination of neighboring cells in contingency tables. Journal of the American Statistical Association, 48(261):104–112, March 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280883.

Crawford:1953:BRB

[Cra53c] John Crawford. Book review: [untitled]. Journal of the American Statistical Association, 48(261):157–159, March 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280893.

Crow:1950:BRB

[Cro50] James F. Crow. Book review: [untitled]. Journal of the American Statistical Association, 45(252):571–573, December 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280728.

Crow:1952:SCW

[Cro52] Edwin L. Crow. Some cases in which Yates’ correction should not be applied. Journal of the American Statistical Association, 47 (258):303–304, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280753.

Crow:1955:GCI

[Cro55] Edwin L. Crow. Generality of confidence intervals for a regression function. Journal of the American Statistical Association, 50 (271):850–853, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281169.

Crow:1956:CGC

[Cro56] Edwin L. Crow. Corrigenda: Generality of confidence intervals for a regression function. Journal of the American Statistical Association, 51(276):650, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281481.

Cromarty:1959:EMU

[Cro59] William A. Cromarty. An econometric model for United States agriculture. Journal of the American Statistical Association, 54 (287):556–574, September 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282535.

Crum:1956:FYR

[Cru56] W. L. Crum. Fiscal-year reporting for corporate income tax. Journal of the American Statistical Association, 51(274):304–369, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2281351.

Crosetti:1956:MEI

[CS56] Albert H. Crosetti and Robert C. Schmitt. A method of estimating the intercensal population of counties. Journal of the American Statistical Association, 51(276):587–590, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281470.

Curtiss:1954:BRB

[Cur54] J. H. Curtiss. Book review: [untitled]. Journal of the American Statistical Association, 49(266):400–401, June 1954. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2280951.

Cutler:1955:RSE

[Cut55] Sidney J. Cutler. A review of the statistical evidence on the association between smoking and lung cancer. Journal of the American Statistical Association, 50(270):267–282, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280961.

Dalenius:1954:BRB

[Dal54] Tore Dalenius. Book review: [untitled]. Journal of the American Statistical Association, 49(267):650–653, September 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281138.

Daly:1958:DFP

[Dal58] Rex F. Daly. Demand for farm products at retail and the farm level: Some empirical measurements and related problems. Journal of the American Statistical Association, 53(283):656–668, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2282044.

Daniel:1953:SCE

[Dan53] C. Daniel. Statistics in chemical experimentation. Journal of the American Statistical Association, 48(263):476–485, September 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2281003.

Darley:1951:BRB

[Dar51] John G. Darley. Book review: [untitled]. Journal of the American Statistical Association, 46(254):277, June 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279941.

Darling:1955:BRB

[Dar55] D. A. Darling. Book review: [untitled]. Journal of the American Statistical Association, 50(270):617–619, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280987.

Davis:1952:ASF

[Dav52] D. J. Davis. An analysis of some failure data. Journal of the American Statistical Association, 47(258):113–150, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280740.

David:1956:RVN

[Dav56] H. A. David. The ranking of variances in normal populations. Journal of the American Statistical Association, 51(276):621–626, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281475.

David:1959:BRB

[Dav59a] H. A. David. Book review: [untitled]. Journal of the American Statistical Association, 54(288):814–815, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282517.

Davis:1959:BRB

[Dav59b] Lane Davis. Book review: [untitled]. Journal of the American Statistical Association, 54(286):513–515, June 1959. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281791.

Day:1951:BRB

[Day51] Besse B. Day. Book review: [untitled]. Journal of the American Statistical Association, 46(255):390–393, September 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280522.

Day:1959:BRB

[Day59] Lincoln H. Day. Book review: [untitled]. Journal of the American Statistical Association, 54(286):522–523, June 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281797.

deCarvalho:1956:BRB

[dC56] Alceu Vicente de Carvalho. Book review: [untitled]. Journal of the American Statistical Association, 51(275):543–544, September 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2281450.

DeFinetti:1951:BRB

[De 51] Bruno De Finetti. Book review: [untitled]. Journal of the American Statistical Association, 46(256):532–534, December 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280415.

Deardorff:1952:PMC

[Dea52] Neva R. Deardorff. Prepaid medical care as a source of morbidity data. Journal of the American Statistical Association, 47(258):281– 302, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280752.

Deemer:1954:CLF

[Dee54] Walter L. Deemer, Jr. Cargo loss in ferrying operations. Journal of the American Statistical Association, 49(268):876–889, December 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2281546.

Deming:1953:PMA

[Dem53a] W. Edwards Deming. On a probability mechanism to attain an economic balance between the resultant error of response and the bias of nonresponse. Journal of the American Statistical Association, 48 (264):743–772, December 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281069.

Deming:1953:DBE

[Dem53b] W. Edwards Deming. On the distinction between enumerative and analytic surveys. Journal of the American Statistical Association, 48(262):244–255, June 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281285.

Deming:1954:PRS

[Dem54] W. Edwards Deming. On the presentation of the results of sample surveys as legal evidence. Journal of the American Statistical Association, 49(268):814–825, December 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281541.

Deming:1955:BRB

[Dem55] W. Edwards Deming. Book review: [untitled]. Journal of the American Statistical Association, 50(270):615–617, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280985.

Deming:1956:SSD

[Dem56] W. Edwards Deming. On simplifications of sampling design through replication with equal probabilities and without stages. Journal of the American Statistical Association, 51(273):24–53, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280803.

Deming:1958:BRB

[Dem58] W. Edwards Deming. Book review: [untitled]. Journal of the American Statistical Association, 53(281):228–229, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282596.

Denison:1955:BRB

[Den55] Edward F. Denison. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1394–1396, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281263.

Derrick:1956:BRB

[Der56] Lucile Derrick. Book review: [untitled]. Journal of the American Statistical Association, 51(274):380, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281357.

Desind:1950:BRB

[Des50] Philip Desind. Book review: [untitled]. Journal of the American Statistical Association, 45(249):144–145, March 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280440.

Deming:1959:PML

[DG59] W. Edwards Deming and Gerald J. Glasser. On the problem of matching lists by samples. Journal of the American Statistical Association, 54(286):403–415, June 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281781.

Daniel:1950:DEM

[DH50] Cuthbert Daniel and Nicholas Heerema. Design of experiments for most precise slope estimation or linear extrapolation. Journal of the American Statistical Association, 45(252):546–556, December 1950.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280723.

Dalenius:1959:MVS

[DH59] Tore Dalenius and Joseph L. Hodges, Jr. Minimum variance stratification. Journal of the American Statistical Association, 54(285):88– 101, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2282141.

See corrigenda [DH63].

Dalenius:1963:CMV

[DH63] Tore Dalenius and Joseph L. Hodges, Jr. Corrigenda: Minimum Variance Stratification. Journal of the American Statistical Association, 58(304):1161, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283341. See [DH59].

Digby:1957:BRB

[Dig57] James F. Digby. Book review: [untitled]. Journal of the American Statistical Association, 52(280):591–592, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281721.

deJanosi:1959:NRB

[dJ59] Peter E. de Janosi. A note on the relationship between earning expectations and new car purchases. Journal of the American Statistical Association, 54(287):575–577, September 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282536.

delaGarza:1956:QER

[dlG56] A. de la Garza. Quadratic extrapolation and a related test of hypotheses. Journal of the American Statistical Association, 51 (276):644–649, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281478.

delaGarza:1955:SMC

[dlGHN55] A. de la Garza, D. S. Hawxhurst, and L. T. Newman. Some minimum cost experimental procedures in quadratic regression. Journal of the American Statistical Association, 50(269):178–184, March 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2281104.

DeGroot:1958:SAU

[DN58] M. H. DeGroot and Jack Nadler. Some aspects of the use of the sequential probability ratio test. Journal of the American Statistical Association, 53(281):187–199, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282579.

Donnelly:1956:BRB

[Don56] T. G. Donnelly. Book review: [untitled]. Journal of the American Statistical Association, 51(275):540–541, September 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281448.

Dorn:1950:BRB

[Dor50a] Harold F. Dorn. Book review: [untitled]. Journal of the American Statistical Association, 45(250):289–290, June 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280689.

Dorn:1950:PPF

[Dor50b] Harold F. Dorn. Pitfalls in population forecasts and projections. Journal of the American Statistical Association, 45(251):311–334, September 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280290.

Dorfman:1954:BRB

[Dor54] Robert Dorfman. Book review: [untitled]. Journal of the American Statistical Association, 49(267):667–670, September 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281146.

Dornbusch:1955:BRB

[Dor55] Sanford M. Dornbusch. Book review: [untitled]. Journal of the American Statistical Association, 50(270):619–620, June 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280989.

Dorn:1956:BRBa

[Dor56a] Harold F. Dorn. Book review: [untitled]. Journal of the American Statistical Association, 51(275):539–540, September 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281447.

Dorn:1956:BRBb

[Dor56b] Harold F. Dorn. Book review: [untitled]. Journal of the American Statistical Association, 51(276):666–667, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281503.

Dorn:1956:BRBc

[Dor56c] Harold F. Dorn. Book review: [untitled]. Journal of the American Statistical Association, 51(276):709–710, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281529.

Dowd:1954:BRB

[Dow54] Douglas F. Dowd. Book review: [untitled]. Journal of the American Statistical Association, 49(267):674–675, September 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281150.

Dunsing:1958:EVD

[DR58] Marilyn Dunsing and Margaret G. Reid. Effect of varying degrees of transitory income on income elasticity of expenditures. Journal of the American Statistical Association, 53(282):348–359, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281860.

Dresch:1958:BRB

[Dre58] Francis W. Dresch. Book review: [untitled]. Journal of the American Statistical Association, 53(283):761, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282063.

Dublin:1950:AJL

[Dub50] Louis I. Dublin. Alfred James Lotka, 1880–1949. Journal of the American Statistical Association, 45(249):138–139, March 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280437.

Dudding:1952:BRB

[Dud52] Bernard P. Dudding. Book review: [untitled]. Journal of the American Statistical Association, 47(260):691–692, December 1952. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280788.

Dunn:1952:EMP

[Dun52] Halbert L. Dunn. Evolving mechanisms for the production of international health statistics. Journal of the American Statistical Association, 47(260):653–662, December 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280783.

Duncan:1954:BRB

[Dun54] Acheson J. Duncan. Book review: [untitled]. Journal of the American Statistical Association, 49(267):654–657, September 1954. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281139.

Dunnett:1955:MCP

[Dun55] Charles W. Dunnett. A multiple comparison procedure for comparing several treatments with a control. Journal of the American Statistical Association, 50(272):1096–1121, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281208.

Duncan:1956:EDB

[Dun56a] Acheson J. Duncan. The economic design of —barX charts used to maintain current control of a process. Journal of the American Statistical Association, 51(274):228–242, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281343.

Duncan:1956:RMP

[Dun56b] Otis Dudley Duncan. Research on metropolitan population: Evaluation of data. Journal of the American Statistical Association, 51 (276):591–596, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281471.

Duncan:1957:BRB

[Dun57a] Acheson J. Duncan. Book review: [untitled]. Journal of the American Statistical Association, 52(280):583–586, December 1957. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281714.

Duncan:1957:CPO

[Dun57b] Acheson J. Duncan. Charts of the 10% and 50% points of the operating characteristic curves for fixed effects analysis of variance F tests, α = 0.01 and 0.05. Journal of the American Statistical Association, 52(279):345–349, September 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280904.

Duncan:1958:DOD

[Dun58] Acheson J. Duncan. Design and operation of a double-limit variables sampling plan. Journal of the American Statistical Association, 53(282):543–550, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281874.

Dunn:1959:BRB

[Dun59a] Olive Jean Dunn. Book review: [untitled]. Journal of the American Statistical Association, 54(288):816–817, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282519.

Dunn:1959:CIM

[Dun59b] Olive Jean Dunn. Confidence intervals for the means of dependent, normally distributed variables. Journal of the American Statistical Association, 54(287):613–621, September 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282541.

Durand:1951:BRB

[Dur51] David Durand. Book review: [untitled]. Journal of the American Statistical Association, 46(256):529–530, December 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280413.

Durbin:1953:NRW

[Dur53] J. Durbin. A note on regression when there is extraneous information about one of the coefficients. Journal of the American Statistical Association, 48(264):799–808, December 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281073.

Durand:1954:JCR

[Dur54] David Durand. Joint confidence regions for multiple regression coefficients. Journal of the American Statistical Association, 49 (265):130–146, March 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281040.

Durand:1956:NMI

[Dur56] David Durand. A note on matrix inversion by the square root method. Journal of the American Statistical Association, 51(274): 288–292, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2281349.

Dwyer:1953:EEM

[DW53a] Paul S. Dwyer and Frederick V. Waugh. Errata: On errors in matrix inversion. Journal of the American Statistical Association, 48 (264):911–912, December 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281085.

Dwyer:1953:EMI

[DW53b] Paul S. Dwyer and Frederick V. Waugh. On errors in matrix inversion. Journal of the American Statistical Association, 48(262):289– 319, June 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281289.

Dwass:1954:BRB

[Dwa54] Meyer Dwass. Book review: [untitled]. Journal of the American Statistical Association, 49(268):925–926, December 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281558.

Dwyer:1951:BRB

[Dwy51] Paul S. Dwyer. Book review: [untitled]. Journal of the American Statistical Association, 46(254):258–259, June 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279928.

Dwyer:1959:BRB

[Dwy59] Paul S. Dwyer. Book review: [untitled]. Journal of the American Statistical Association, 54(287):701–702, September 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282550.

Dykstra:1958:FES

[Dyk58] Otto Dykstra, Jr. Factorial experimentation in Scheffe’s analysis of variance for paired comparisons. Journal of the American Statistical Association, 53(282):529–542, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281873.

El-Badry:1956:SPM

[EB56] M. A. El-Badry. A sampling procedure for mailed questionnaires. Journal of the American Statistical Association, 51(274):209–227, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2281342.

Eckler:1955:CEL

[EBP55] A. Ross Eckler, Gertrude Bancroft, and Robert Pearl. Concepts employed in labor force measurements and uses of labor force data. Journal of the American Statistical Association, 50(271):677–688, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281159.

El-Badry:1955:AST

[EBS55] M. A. El-Badry and F. F. Stephan. On adjusting sample tabulations to census counts. Journal of the American Statistical Association, 50(271):738–762, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281163.

Eckstein:1958:BRB

[Eck58] Alexander Eckstein. Book review: [untitled]. Journal of the American Statistical Association, 53(281):238–240, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282604.

Edgett:1956:MRM

[Edg56] George L. Edgett. Multiple regression with missing observations among the independent variables. Journal of the American Statistical Association, 51(273):122–131, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280808.

Edwards:1958:BRB

[Edw58] Edgar O. Edwards. Book review: [untitled]. Journal of the American Statistical Association, 53(282):615–617, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281894.

Eisenhart:1953:BRB

[Eis53] Churchill Eisenhart. Book review: [untitled]. Journal of the American Statistical Association, 48(262):359–361, June 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281294.

Eisenhart:1955:BRB

[Eis55] Churchill Eisenhart. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1335–1337, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281229.

Eisenpress:1956:RTA

[Eis56] Harry Eisenpress. Regression techniques applied to seasonal corrections and adjustments for calendar shifts. Journal of the American Statistical Association, 51(276):615–620, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281474.

Eisemann:1958:MIC

[Eis58] Doris M. Eisemann. Manufacturers’ inventory cycles and monetary policy. Journal of the American Statistical Association, 53 (283):680–688, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282046.

Eisenpress:1959:BRB

[Eis59] Harry Eisenpress. Book review: [untitled]. Journal of the American Statistical Association, 54(285):319–320, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282161.

Elkin:1953:ERB

[Elk53] Jack M. Elkin. Estimating the ratio between the proportions of two classes when one is a sub-class of the other. Journal of the American Statistical Association, 48(261):128–130, March 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280885.

Elveback:1958:BRB

[Elv58a] Lila Elveback. Book review: [untitled]. Journal of the American Statistical Association, 53(284):1038–1039, December 1958. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281979.

Elveback:1958:ESC

[Elv58b] Lila Elveback. Estimation of survivorship in chronic disease: The “actuarial” method. Journal of the American Statistical Association, 53(282):420–440, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281865.

Enke:1956:BRB

[Enk56] Stephen Enke. Book review: [untitled]. Journal of the American Statistical Association, 51(275):551–553, September 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281457.

Epstein:1954:BRB

[Eps54] Benjamin Epstein. Book review: [untitled]. Journal of the American Statistical Association, 49(267):665–667, September 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281144.

Epstein:1955:CSN

[Eps55] Benjamin Epstein. Comparison of some non-parametric tests against normal alternatives with an application to life testing. Journal of the American Statistical Association, 50(271):894–900, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281174.

Ernst:1958:IMP

[Ern58] Harry Benjamin Ernst. An index of manufacturing production in New England. Journal of the American Statistical Association, 53 (282):336–347, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281859.

Epstein:1953:LT

[ES53] Benjamin Epstein and Milton Sobel. Life testing. Journal of the American Statistical Association, 48(263):486–502, September 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2281004.

Evans:1951:SOA

[Eva51a] W. Duane Evans. On stratification and optimum allocations. Journal of the American Statistical Association, 46(253):95–104, March 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280098.

Evans:1951:VES

[Eva51b] W. Duane Evans. On the variance of estimates of the standard deviation and variance. Journal of the American Statistical Association, 46(254):220–224, June 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279922.

Evans:1957:BRB

[Eva57] W. Duane Evans. Book review: [untitled]. Journal of the American Statistical Association, 52(280):592–593, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281722.

F:1951:BRBa

[F.51a] S. F. Book review: [untitled]. Journal of the American Statistical Association, 46(255):406, September 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280532.

F:1951:BRBb

[F.51b] S. F. Book review: [untitled]. Journal of the American Statistical Association, 46(255):407, September 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280533.

F:1951:BRBc

[F.51c] S. F. Book review: [untitled]. Journal of the American Statistical Association, 46(256):522–523, December 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280406.

F:1951:BRBd

[F.51d] S. F. Book review: [untitled]. Journal of the American Statistical Association, 46(256):523, December 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280407.

F:1956:BRBa

[F.56a] D. D. F. Book review: [untitled]. Journal of the American Statistical Association, 51(276):659–660, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281496.

F:1956:BRBb

[F.56b] D. D. F. Book review: [untitled]. Journal of the American Statistical Association, 51(276):676, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281509.

F:1957:BRBa

[F.57a] D. D. F. Book review: [untitled]. Journal of the American Statistical Association, 52(277):102, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281405.

F:1957:BRBb

[F.57b] D. D. F. Book review: [untitled]. Journal of the American Statistical Association, 52(278):272, June 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280857.

F:1957:BRBc

[F.57c] D. D. F. Book review: [untitled]. Journal of the American Statistical Association, 52(279):393, September 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280916.

F:1958:BRBa

[F.58a] D. D. F. Book review: [untitled]. Journal of the American Statistical Association, 53(282):620–621, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281897.

F:1958:BRBb

[F.58b] D. D. F. Book review: [untitled]. Journal of the American Statistical Association, 53(283):757–758, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282060.

F:1958:BRBc

[F.58c] D. D. F. Book review: [untitled]. Journal of the American Statistical Association, 53(283):758, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282061.

F:1959:BRB

[F.59] R. F. Book review: [untitled]. Journal of the American Statistical Association, 54(288):820, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282522.

Fabricant:1952:FAS

[Fab52] Solomon Fabricant. Factors in the accumulation of social statistics. Journal of the American Statistical Association, 47(258):255–262, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280749.

Fabricant:1954:CBP

[Fab54] Solomon Fabricant. Cycles in the balance of payments. Journal of the American Statistical Association, 49(265):79–87, March 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281036.

Fand:1958:BRB

[Fan58] David I. Fand. Book review: [untitled]. Journal of the American Statistical Association, 53(281):242–245, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282607.

Federighi:1959:ETP

[Fed59] Enrico T. Federighi. Extended tables of the percentage points of Student’s t-distribution. Journal of the American Statistical Association, 54(287):683–688, September 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282546.

Fels:1957:BRB

[Fel57] Eberhard M. Fels. Book review: [untitled]. Journal of the American Statistical Association, 52(278):283–285, June 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280868.

Fels:1959:SSS

[Fel59] Eberhard M. Fels. Some Soviet statistical books of 1957. Journal of the American Statistical Association, 54(285):12–29, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282136.

Fertig:1950:BRB

[Fer50] John W. Fertig. Book review: [untitled]. Journal of the American Statistical Association, 45(250):291–292, June 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280690.

Ferber:1953:MAS

[Fer53] Robert Ferber. Measuring the accuracy and structure of Businessmen’s expectations. Journal of the American Statistical Association, 48(263):385–413, September 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280997.

Ferber:1955:BRB

[Fer55a] Robert Ferber. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1414–1415, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281273.

Ferber:1955:RRS

[Fer55b] Robert Ferber. On the reliability of responses secured in sample surveys. Journal of the American Statistical Association, 50 (271):788–810, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281166.

Ferber:1956:ERI

[Fer56] Robert Ferber. The effect of respondent ignorance on survey results. Journal of the American Statistical Association, 51(276):576–586, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281469.

Ferber:1958:ASI

[Fer58a] Robert Ferber. The accuracy and structure of industry expectations in relation to those of individual firms. Journal of the American Statistical Association, 53(282):317–335, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281858.

Ferris:1958:BRB

[Fer58b] George E. Ferris. Book review: [untitled]. Journal of the American Statistical Association, 53(281):252, March 1958. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282613.

Feyerherm:1959:BRB

[Fey59] A. M. Feyerherm. Book review: [untitled]. Journal of the American Statistical Association, 54(286):512, June 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281790.

Freudenthal:1954:MLF

[FG54] A. M. Freudenthal and E. J. Gumbel. Minimum life in fatigue. Journal of the American Statistical Association, 49(267):575–597, September 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281131.

Ficken:1955:BRB

[Fic55] F. A. Ficken. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1359–1360, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281241.

Finney:1957:BRB

[Fin57] D. J. Finney. Book review: [untitled]. Journal of the American Statistical Association, 52(279):388–389, September 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280912.

Firey:1951:BRB

[Fir51] Walter Firey. Book review: [untitled]. Journal of the American Statistical Association, 46(253):140–141, March 1951. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2280115.

Fisher:1951:BRB

[Fis51] Walter D. Fisher. Book review: [untitled]. Journal of the American Statistical Association, 46(253):130–132, March 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280107.

Fisher:1952:HEC

[Fis52] Gene H. Fisher. Hicks’ “elementary case” economic model for the United States, 1929–1941. Journal of the American Statistical Association, 47(259):541–549, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281321.

Fisher:1953:DME

[Fis53] Jacob Fisher. Data for measuring the effectiveness of public incomemaintenance programs. Journal of the American Statistical Association, 48(261):15–27, March 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280876.

Fisher:1954:BRB

[Fis54] Walter D. Fisher. Book review: [untitled]. Journal of the American Statistical Association, 49(268):923–925, December 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281557.

Fisher:1958:GMH

[Fis58] Walter D. Fisher. On grouping for maximum homogeneity. Journal of the American Statistical Association, 53(284):789–798, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2281952.

FitzPatrick:1957:LAS

[Fit57] Paul J. FitzPatrick. Leading American statisticians in the nineteenth century. Journal of the American Statistical Association, 52 (279):301–321, September 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280901.

FitzPatrick:1958:LAS

[Fit58] Paul J. FitzPatrick. Leading American statisticians of the nineteenth century II. Journal of the American Statistical Association, 53(283):689–701, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282047.

Foley:1953:CTU

[Fol53] Donald L. Foley. Census tracts and urban research. Journal of the American Statistical Association, 48(264):733–742, December 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281068.

Folger:1954:BRB

[Fol54] John Folger. Book review: [untitled]. Journal of the American Statistical Association, 49(266):402, June 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280953.

Foote:1953:MBB

[Foo53] Richard J. Foote. The mathematical basis for the bean method of graphic multiple correlation. Journal of the American Statistical Association, 48(264):778–788, December 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281071.

Foote:1958:MDA

[Foo58] Richard J. Foote. A modified Doolittle approach for multiple and partial correlation and regression. Journal of the American Statistical Association, 53(281):133–143, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282573.

Foote:1959:BRB

[Foo59] Richard J. Foote. Book review: [untitled]. Journal of the American Statistical Association, 54(288):822–823, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282524.

Fox:1951:RBP

[Fox51a] Karl A. Fox. Relations between prices, consumption, and production. Journal of the American Statistical Association, 46(255): 323–333, September 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280510.

Fox:1951:BRB

[Fox51b] Philip G. Fox. Book review: [untitled]. Journal of the American Statistical Association, 46(253):130, March 1951. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280106.

Fox:1953:BRB

[Fox53] Karl A. Fox. Book review: [untitled]. Journal of the American Statistical Association, 48(263):665–666, September 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281024.

Frankel:1951:BRB

[Fra51] Lester R. Frankel. Book review: [untitled]. Journal of the American Statistical Association, 46(253):127–129, March 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280103.

Frankel:1954:BRB

[Fra54] Lester R. Frankel. Book review: [untitled]. Journal of the American Statistical Association, 49(265):189–190, March 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281049.

Freeman:1950:BRB

[Fre50a] H. A. Freeman. Book review: [untitled]. Journal of the American Statistical Association, 45(249):145–147, March 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280441.

Freund:1950:BRB

[Fre50b] John E. Freund. Book review: [untitled]. Journal of the American Statistical Association, 45(251):459–460, September 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280304.

Freund:1950:DS

[Fre50c] John E. Freund. The degree of stereotypy. Journal of the American Statistical Association, 45(250):265–269, June 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280685.

Freeman:1951:BRB

[Fre51] Harold A. Freeman. Book review: [untitled]. Journal of the American Statistical Association, 46(255):409–410, September 1951. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280535.

Freeman:1955:BRB

[Fre55] Harold A. Freeman. Book review: [untitled]. Journal of the American Statistical Association, 50(270):610–611, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280981.

French:1959:BRB

[Fre59a] Ben C. French. Book review: [untitled]. Journal of the American Statistical Association, 54(285):317–318, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282158.

Freund:1959:BRB

[Fre59b] Rudolf J. Freund. Book review: [untitled]. Journal of the American Statistical Association, 54(285):315, March 1959. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282155.

Fry:1951:BRB

[Fry51] Thornton C. Fry. Book review: [untitled]. Journal of the American Statistical Association, 46(253):132–134, March 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280108.

Fry:1956:ACI

[Fry56] Thornton C. Fry. The automatic computer in industry. Journal of the American Statistical Association, 51(276):565–575, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2281468.

Fuller:1955:BRB

[Ful55] Varden Fuller. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1406–1408, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281269.

G:1953:BRB

[G.53] U. G. Book review: [untitled]. Journal of the American Statistical Association, 48(261):155, March 1953. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280890.

Garvy:1952:BRB

[Gar52] George Garvy. Book review: [untitled]. Journal of the American Statistical Association, 47(260):694–697, December 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280790.

Garvy:1953:VTD

[Gar53] George Garvy. The velocity of time deposits. Journal of the American Statistical Association, 48(262):176–191, June 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281282.

Garvy:1955:BRB

[Gar55] George Garvy. Book review: [untitled]. Journal of the American Statistical Association, 50(270):625–627, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280992.

Garner:1956:OCC

[Gar56a] Norman R. Garner. The operating characteristic curve for sequential sampling by variables when the Producer’s and Consumer’s risks are equal. Journal of the American Statistical Association, 51(273):108–110, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280806.

Garvy:1956:BRB

[Gar56b] George Garvy. Book review: [untitled]. Journal of the American Statistical Association, 51(276):677–682, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281511.

Garner:1958:CSV

[Gar58] Norman R. Garner. Curtailed sampling for variables. Journal of the American Statistical Association, 53(284):862–867, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281955.

Gaylor:1956:ETE

[Gay56] D. W. Gaylor. Equivalence of two estimates of product variance. Journal of the American Statistical Association, 51(275):451–453, September 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281433.

Geary:1953:NLF

[Gea53] R. C. Geary. Non-linear functional relationship between two variables when one variable is controlled. Journal of the American Statistical Association, 48(261):94–103, March 1953. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2280882.

Gershenson:1953:BRB

[Ger53] M. I. Gershenson. Book review: [untitled]. Journal of the American Statistical Association, 48(264):915–916, December 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281088.

Gerschenkron:1957:PML

[Ger57] Alexander Gerschenkron. Problems in measuring long term growth in income and wealth. Journal of the American Statistical Association, 52(280):450–457, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281693.

Gates:1959:ARA

[GF59] C. R. Gates and J. P. Fearey. Accuracy requirements for acceptance testing of complex systems. Journal of the American Statistical Association, 54(286):447–464, June 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281783.

Golub:1956:ASE

[GG56a] Abraham Golub and Frank E. Grubbs. Analysis of sensitivity experiments when the levels of stimulus cannot be controlled. Journal of the American Statistical Association, 51(274):257–265, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281345.

Golub:1956:CAS

[GG56b] Abraham Golub and Frank E. Grubbs. Corrigenda: Analysis of Sensitivity Experiments when the Levels of Stimulas Cannot be Controlled. Journal of the American Statistical Association, 51 (276):650–651, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281482. See [GG67].

Golub:1967:CAS

[GG67] Abraham Golub and Frank E. Grubbs. Corrigenda: Analysis of Sensitivity Experiments when the Levels of Stimulus Cannot be Controlled. Journal of the American Statistical Association, 62 (320):1519, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2283814. See [GG56b].

Gumbel:1953:CND

[GGD53] E. J. Gumbel, J. Arthur Greenwood, and David Durand. The circular normal distribution: Theory and tables. Journal of the American Statistical Association, 48(261):131–152, March 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280886.

Goodman:1958:PUR

[GH58] Leo A. Goodman and H. O. Hartley. The precision of unbiased ratio-type estimators. Journal of the American Statistical Association, 53(282):491–508, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor. org/stable/2281870. See corrigenda [GH63].

Goodman:1963:CPU

[GH63] Leo A. Goodman and H. O. Hartley. Corrigenda: The Precision of Unbiased Ratio-Type Estimators. Journal of the American Statistical Association, 58(304):1162, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2283344. See [GH58].

Gillman:1950:OAT

[Gil50] Leonard Gillman. Operations analysis and the theory of games: An advertising example. Journal of the American Statistical Association, 45(252):541–545, December 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280722.

Gilbert:1957:BRB

[Gil57a] Milton Gilbert. Book review: [untitled]. Journal of the American Statistical Association, 52(277):113–114, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281418.

Gillim:1957:BRB

[Gil57b] Marion Hamilton Gillim. Book review: [untitled]. Journal of the American Statistical Association, 52(277):112–113, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281415.

Gilbert:1959:BRB

[Gil59] John P. Gilbert. Book review: [untitled]. Journal of the American Statistical Association, 54(286):510–511, June 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281788.

Ginzberg:1952:BRB

[Gin52] Eli Ginzberg. Book review: [untitled]. Journal of the American Statistical Association, 47(259):561–565, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281331.

Girshick:1953:BRB

[Gir53] M. A. Girshick. Book review: [untitled]. Journal of the American Statistical Association, 48(263):645–647, September 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281014.

Goodman:1950:CST

[GK50] Roe Goodman and Leslie Kish. Controlled selection — a technique in probability sampling. Journal of the American Statistical Association, 45(251):350–372, September 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280293.

Goodman:1954:MAC

[GK54] Leo A. Goodman and William H. Kruskal. Measures of association for cross classifications. Journal of the American Statistical Association, 49(268):732–764, December 1954. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281536.

Goodman:1957:CMA

[GK57] Leo A. Goodman and William H. Kruskal. Corrigenda: Measures of association for cross classifications. Journal of the American Statistical Association, 52(280):578, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281708.

Goodman:1958:CMA

[GK58] Leo A. Goodman and William H. Kruskal. Corrigenda: Measures of association for cross classifications. Journal of the American Statistical Association, 53(284):1031, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281967.

Goodman:1959:MAC

[GK59] Leo A. Goodman and William H. Kruskal. Measures of association for cross classifications. II: Further discussion and references. Journal of the American Statistical Association, 54(285):123–163, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2282143.

Glover:1956:BRB

[Glo56] K. F. Glover. Book review: [untitled]. Journal of the American Statistical Association, 51(274):396–398, June 1956. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281371.

Godwin:1955:GTI

[God55] H. J. Godwin. On generalizations of Tchebychef’s inequality. Journal of the American Statistical Association, 50(271):923–945, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281177.

Gogek:1954:BRB

[Gog54] Sophia Gogek. Book review: [untitled]. Journal of the American Statistical Association, 49(266):400, June 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280950.

Golub:1953:DSS

[Gol53] Abraham Golub. Designing single-sampling inspection plans when the sample size is fixed. Journal of the American Statistical Association, 48(262):278–288, June 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281288.

Goldhamer:1955:BRB

[Gol55a] Herbert Goldhamer. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1384–1385, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281256.

Goldsmith:1955:BRB

[Gol55b] Selma F. Goldsmith. Book review: [untitled]. Journal of the American Statistical Association, 50(271):992–994, September 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281190.

Goldberger:1958:BRB

[Gol58] Arthur S. Goldberger. Book review: [untitled]. Journal of the American Statistical Association, 53(284):1050–1052, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281988.

Goodman:1952:SNA

[Goo52] Leo A. Goodman. Serial number analysis. Journal of the American Statistical Association, 47(260):622–634, December 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280780.

Goodman:1953:BRB

[Goo53a] Leo A. Goodman. Book review: [untitled]. Journal of the American Statistical Association, 48(261):153, March 1953. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2280887.

Goodman:1953:MMU

[Goo53b] Leo A. Goodman. Methods of measuring useful life of equipment under operational conditions. Journal of the American Statistical Association, 48(263):503–530, September 1953. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281005.

Goodman:1954:SPT

[Goo54] Leo A. Goodman. Some practical techniques in serial number analysis. Journal of the American Statistical Association, 49(265):97– 112, March 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281038.

Good:1956:BRB

[Goo56] I. J. Good. Book review: [untitled]. Journal of the American Statistical Association, 51(274):388–390, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281364.

Good:1958:STP

[Goo58a] I. J. Good. Significance tests in parallel and in series. Journal of the American Statistical Association, 53(284):799–813, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281953.

Goodman:1958:BRB

[Goo58b] Leo A. Goodman. Book review: [untitled]. Journal of the American Statistical Association, 53(281):246–247, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282609.

Gordon:1956:BRBa

[Gor56a] Margaret S. Gordon. Book review: [untitled]. Journal of the American Statistical Association, 51(273):189–190, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280826.

Gordon:1956:BRBb

[Gor56b] R. A. Gordon. Book review: [untitled]. Journal of the American Statistical Association, 51(274):394–395, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281369.

Gordon:1958:BRB

[Gor58] Margaret S. Gordon. Book review: [untitled]. Journal of the American Statistical Association, 53(282):606, June 1958. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281887.

Gould:1952:BRB

[Gou52] J. M. Gould. Book review: [untitled]. Journal of the American Statistical Association, 47(258):329–330, June 1952. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2280762.

Gray:1955:MFS

[Gra55] Percy G. Gray. The memory factor in social surveys. Journal of the American Statistical Association, 50(270):344–363, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280965.

Gray:1956:CMF

[Gra56] Percy G. Gray. Corrigenda: The memory factor in social surveys. Journal of the American Statistical Association, 51(276):651, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281483.

Graves:1958:BRB

[Gra58] Robert L. Graves. Book review: [untitled]. Journal of the American Statistical Association, 53(284):1043–1044, December 1958. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281982.

Grabill:1959:BRB

[Gra59] Wilson H. Grabill. Book review: [untitled]. Journal of the American Statistical Association, 54(287):705–707, September 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282553.

Green:1952:BRB

[Gre52a] Howard Whipple Green. Book review: [untitled]. Journal of the American Statistical Association, 47(259):560–561, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2281330.

Green:1952:LSA

[Gre52b] Bert F. Green, Jr. Latent structure analysis and its relation to factor analysis. Journal of the American Statistical Association, 47(257):71–76, March 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279978.

Greenberg:1952:BRB

[Gre52c] B. G. Greenberg. Book review: [untitled]. Journal of the American Statistical Association, 47(260):706–708, December 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280796.

Greville:1952:BRB

[Gre52d] Thomas N. E. Greville. Book review: [untitled]. Journal of the American Statistical Association, 47(259):558–560, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2281329.

Green:1953:BRB

[Gre53a] Bert F. Green. Book review: [untitled]. Journal of the American Statistical Association, 48(263):657–659, September 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281020.

Greenwood:1953:PCS

[Gre53b] Robert E. Greenwood. Probabilities of certain solitaire card games. Journal of the American Statistical Association, 48(261):88–93, March 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280881.

Grenander:1954:BRB

[Gre54] Ulf Grenander. Book review: [untitled]. Journal of the American Statistical Association, 49(267):665, September 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281143.

Gregory:1955:BRB

[Gre55] Geoffrey Gregory. Book review: [untitled]. Journal of the American Statistical Association, 50(270):612–613, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280983.

Greenberg:1956:BRB

[Gre56a] B. G. Greenberg. Book review: [untitled]. Journal of the American Statistical Association, 51(273):170–173, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280816.

Gretton:1956:BRB

[Gre56b] Owen C. Gretton. Book review: [untitled]. Journal of the American Statistical Association, 51(276):674–675, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281508.

Greenwald:1957:BRB

[Gre57] William I. Greenwald. Book review: [untitled]. Journal of the American Statistical Association, 52(278):278–279, June 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280863.

Greenwald:1958:BRB

[Gre58a] William I. Greenwald. Book review: [untitled]. Journal of the American Statistical Association, 53(284):1032–1033, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281971.

Gregory:1958:BRB

[Gre58b] G. R. Gregory. Book review: [untitled]. Journal of the American Statistical Association, 53(282):614–615, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281893.

Gringorten:1951:VSW

[Gri51] Iriving I. Gringorten. The verification and scoring of weather forecasts. Journal of the American Statistical Association, 46(255): 279–296, September 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280507.

Griffin:1957:BRBa

[Gri57a] John I. Griffin. Book review: [untitled]. Journal of the American Statistical Association, 52(277):104–105, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281407.

Griffin:1957:BRBb

[Gri57b] John I. Griffin. Book review: [untitled]. Journal of the American Statistical Association, 52(280):582–583, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281713.

Griffin:1958:GCT

[Gri58] Harold D. Griffin. Graphic computation of tau as a coefficient of disarray. Journal of the American Statistical Association, 53(282):441– 447, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281866.

Gridgeman:1959:LTT

[Gri59a] N. T. Gridgeman. The lady tasting tea, and allied topics. Journal of the American Statistical Association, 54(288):776–783, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2282501.

Griliches:1959:DFI

[Gri59b] Zvi Griliches. The demand for fertilizer in 1954: An inter-state study. Journal of the American Statistical Association, 54(286): 377–384, June 1959. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2281778.

Greenwood:1950:SSR

[GS50] Joseph A. Greenwood and Marion M. Sandomire. Sample size required for estimating the standard deviation as a per cent of its true value. Journal of the American Statistical Association, 45(250):257– 260, June 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280683.

Grab:1954:ETE

[GS54a] Edwin L. Grab and I. Richard Savage. Errata: Tables of the expected value of 1/X for positive Bernoulli and Poisson variables. Journal of the American Statistical Association, 49(268):906, December 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281549.

Grab:1954:TEV

[GS54b] Edwin L. Grab and I. Richard Savage. Tables of the expected value of 1/X for positive Bernoulli and Poisson variables. Journal of the American Statistical Association, 49(265):169–177, March 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281043.

Greenberg:1959:MII

[GS59] G. Greenberg and A. E. Sarhan. Matrix inversion, its interest and application in analysis of data. Journal of the American Statistical Association, 54(288):755–766, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282499.

Gordon:1956:CLS

[GT56] Myron J. Gordon and William J. Taylor. The condition for lot size production. Journal of the American Statistical Association, 51 (276):627–636, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281476.

Gumbel:1954:ACN

[Gum54] E. J. Gumbel. Applications of the circular normal distribution. Journal of the American Statistical Association, 49(266):267–297, June 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280935.

Gurland:1954:BRB

[Gur54] John Gurland. Book review: [untitled]. Journal of the American Statistical Association, 49(265):197–200, March 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281054.

Guttman:1956:BRB

[Gut56] Louis Guttman. Book review: [untitled]. Journal of the American Statistical Association, 51(275):560–561, September 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281463.

Graybill:1956:NUB

[GW56] Franklin A. Graybill and A. W. Wortha. A note on uniformly best unbiased estimators for variance components. Journal of the American Statistical Association, 51(274):266–268, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281346.

Greenberg:1950:TAS

[GWS50] B. G. Greenberg, John J. Wright, and Cecil G. Sheps. A technique for analyzing some factors affecting the incidence of syphilis.

Journal of the American Statistical Association, 45(251):373–399, September 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280294.

Hader:1955:BRB

[Had55] Robert J. Hader. Book review: [untitled]. Journal of the American Statistical Association, 50(271):989–990, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281188.

Haemer:1955:BRBa

[Hae55a] Kenneth W. Haemer. Book review: [untitled]. Journal of the American Statistical Association, 50(269):260–261, March 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281114.

Haemer:1955:BRBb

[Hae55b] Kenneth W. Haemer. Book review: [untitled]. Journal of the American Statistical Association, 50(269):261–262, March 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281115.

Haemer:1955:BRBc

[Hae55c] Kenneth W. Haemer. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1341–1343, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281233.

Haemer:1957:BRBa

[Hae57a] Kenneth W. Haemer. Book review: [untitled]. Journal of the American Statistical Association, 52(277):107–108, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281410.

Haemer:1957:BRBb

[Hae57b] Kenneth W. Haemer. Book review: [untitled]. Journal of the American Statistical Association, 52(277):108–110, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281411.

Haemer:1957:BRBc

[Hae57c] Kenneth W. Haemer. Book review: [untitled]. Journal of the American Statistical Association, 52(277):110–111, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281412.

Hagedorn:1952:BRB

[Hag52] George G. Hagedorn. Book review: [untitled]. Journal of the American Statistical Association, 47(259):569–571, September 1952. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281334.

Hajnal:1955:PPF

[Haj55] John Hajnal. The prospects for population forecasts. Journal of the American Statistical Association, 50(270):309–322, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280963.

Hall:1950:BRB

[Hal50] Ray Ovid Hall. Book review: [untitled]. Journal of the American Statistical Association, 45(250):292–294, June 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280691.

Halperin:1952:ETN

[Hal52] Max Halperin. Estimation in the truncated normal distribution. Journal of the American Statistical Association, 47(259):457–465, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281315.

Halmos:1953:BRB

[Hal53] Paul R. Halmos. Book review: [untitled]. Journal of the American Statistical Association, 48(261):154–155, March 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280889.

Hall:1955:BRB

[Hal55a] Oswald Hall. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1391, December 1955. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281261.

Halmos:1955:BRB

[Hal55b] Paul R. Halmos. Book review: [untitled]. Journal of the American Statistical Association, 50(271):1002, September 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281195.

Hammer:1950:ICP

[Ham50] Preston C. Hammer. Interference with a controlled process. Journal of the American Statistical Association, 45(250):249–256, June 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280682.

Hamilton:1952:BRB

[Ham52] C. Horace Hamilton. Book review: [untitled]. Journal of the American Statistical Association, 47(260):702–704, December 1952. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280794.

Hamaker:1954:BRB

[Ham54a] H. C. Hamaker. Book review: [untitled]. Journal of the American Statistical Association, 49(266):386–389, June 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280944.

Hammer:1954:BRB

[Ham54b] P. C. Hammer. Book review: [untitled]. Journal of the American Statistical Association, 49(265):195, March 1954. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281052.

Hammond:1959:IRT

[Ham59] E. Cuyler Hammond. Inhalation in relation to type and amount of smoking. Journal of the American Statistical Association, 54 (285):35–51, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282138.

Handler:1956:BRB

[Han56] A. B. Handler. Book review: [untitled]. Journal of the American Statistical Association, 51(273):182–183, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280822.

Harberger:1951:BRB

[Har51a] Arnold C. Harberger. Book review: [untitled]. Journal of the American Statistical Association, 46(255):395–396, September 1951. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280526.

Hartkemeier:1951:BRB

[Har51b] Harry P. Hartkemeier. Book review: [untitled]. Journal of the American Statistical Association, 46(256):523–524, December 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280408.

Harshbarger:1952:BRB

[Har52] Boyd Harshbarger. Book review: [untitled]. Journal of the American Statistical Association, 47(257):109–110, March 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279993.

Hartkemeier:1953:BRB

[Har53] Harry P. Hartkemeier. Book review: [untitled]. Journal of the American Statistical Association, 48(262):372–374, June 1953. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281301.

Harrington:1955:BRB

[Har55] Edwin C. Harrington, Jr. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1361, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281242.

Harberger:1956:BRB

[Har56a] Arnold C. Harberger. Book review: [untitled]. Journal of the American Statistical Association, 51(275):550–551, September 1956. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281456.

Harberger:1956:EPE

[Har56b] Arnold C. Harberger. The economics of the President’s economic reports. Journal of the American Statistical Association, 51(275): 454–460, September 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281434.

Harvey:1956:BRB

[Har56c] Robert O. Harvey. Book review: [untitled]. Journal of the American Statistical Association, 51(276):686–688, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281514.

Harrison:1959:BRB

[Har59] Stephen Harrison. Book review: [untitled]. Journal of the American Statistical Association, 54(285):320–322, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282162.

Hastay:1951:BRB

[Has51] Millard Hastay. Book review: [untitled]. Journal of the American Statistical Association, 46(255):388–390, September 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280521.

Hastay:1953:BRB

[Has53] Millard Hastay. Book review: [untitled]. Journal of the American Statistical Association, 48(261):162–163, March 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280895.

Hassler:1956:BRB

[Has56] J. B. Hassler. Book review: [untitled]. Journal of the American Statistical Association, 51(276):700–701, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281522.

Hauser:1956:BRB

[Hau56] Philip M. Hauser. Book review: [untitled]. Journal of the American Statistical Association, 51(273):201–202, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280833.

Healy:1953:HTS

[HD53] M. J. R. Healy and G. V. Dyke. A Hollerith technique for the solution of normal equations. Journal of the American Statistical Association, 48(264):809–815, December 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281074.

Heflebower:1951:EAP

[Hef51] Richard B. Heflebower. An economic appraisal of price measures. Journal of the American Statistical Association, 46(256):461–479, December 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280396.

Heinze:1956:BRB

[Hei56] Shirley J. Heinze. Book review: [untitled]. Journal of the American Statistical Association, 51(274):400–402, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281373.

Hendricks:1956:BRB

[Hen56] Walter A. Hendricks. Book review: [untitled]. Journal of the American Statistical Association, 51(274):405, June 1956. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281376.

Henry:1957:BRB

[Hen57] Edwin R. Henry. Book review: [untitled]. Journal of the American Statistical Association, 52(279):409–410, September 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280925.

Herfindahl:1957:BRB

[Her57] Orris C. Herfindahl. Book review: [untitled]. Journal of the American Statistical Association, 52(277):119–122, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281422.

Halperin:1955:TPP

[HGCZ55] Max Halperin, Samuel W. Greenhouse, Jerome Cornfield, and Julia Zalokar. Tables of percentage points for the Studentized maximum absolute deviate in normal samples. Journal of the American Statistical Association, 50(269):185–195, March 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281105.

Hansen:1951:RES

[HHMM51] Morris H. Hansen, William N. Hurwitz, Eli S. Marks, and W. Parker Mauldin. Response errors in surveys. Journal of the American Statistical Association, 46(254):147–190, June 1951. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2279918.

Hansen:1955:RCC

[HHNS55] Morris H. Hansen, William N. Hurwitz, Harold Nisselson, and Joseph Steinberg. The redesign of the census current population survey. Journal of the American Statistical Association, 50 (271):701–719, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281161.

Hickman:1955:BRB

[Hic55] Bert G. Hickman. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1404–1406, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281268.

Hickman:1958:EWI

[Hic58] Bert G. Hickman. An experiment with weighted indexes of cyclical diffusion. Journal of the American Statistical Association, 53 (281):39–53, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282564.

Hildreth:1954:PEO

[Hil54] Clifford Hildreth. Point estimates of ordinates of concave functions. Journal of the American Statistical Association, 49(267):598–619, September 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281132.

Hildebrand:1955:BRB

[Hil55a] George H. Hildebrand. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1397–1399, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2281265.

Hildreth:1955:CPE

[Hil55b] Clifford Hildreth. Corrigenda: Point estimates of ordinates of concave functions. Journal of the American Statistical Association, 50 (272):1331–1332, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281222.

Himsworth:1955:BRB

[Him55] F. R. Himsworth. Book review: [untitled]. Journal of the American Statistical Association, 50(271):980–984, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281183.

Hirsch:1955:BRB

[Hir55] Werner Z. Hirsch. Book review: [untitled]. Journal of the American Statistical Association, 50(271):984–986, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281184.

Hirsch:1956:SPM

[Hir56] Werner Z. Hirsch. Statistician and policy maker: a partnership in the making. Journal of the American Statistical Association, 51(273):12–16, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280801.

Hanson:1958:IIA

[HM58] Robert H. Hanson and Eli S. Marks. Influence of the interviewer on the accuracy of survey results. Journal of the American Statistical Association, 53(283):635–655, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282043.

Hoadley:1950:BRB

[Hoa50] Walter E. Hoadley, Jr. Book review: [untitled]. Journal of the American Statistical Association, 45(252):579–581, December 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280733.

Hoadley:1957:BRB

[Hoa57] Walter E. Hoadley, Jr. Book review: [untitled]. Journal of the American Statistical Association, 52(277):123–124, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281424.

Hoadley:1959:STT

[Hoa59] Walter E. Hoadley, Jr. Statisticians — today and tomorrow: Presidential address. Journal of the American Statistical Association, 54(285):1–11, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282135.

Hodges:1952:BRB

[Hod52] J. L. Hodges. Book review: [untitled]. Journal of the American Statistical Association, 47(258):315–317, June 1952. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2280756.

Hodges:1955:BRB

[Hod55] J. L. Hodges, Jr. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1365–1367, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281245.

Hoeffding:1955:BRB

[Hoe55] Wassily Hoeffding. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1343–1345, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281234.

Holland:1951:BRB

[Hol51] Daniel M. Holland. Book review: [untitled]. Journal of the American Statistical Association, 46(255):407–409, September 1951. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280534.

Holloway:1957:SMF

[Hol57] Clark Holloway, Jr. A systematic method of finding defining contrasts. Journal of the American Statistical Association, 52(277):46– 52, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281399.

Holt:1958:BRB

[Hol58] Charles C. Holt. Book review: [untitled]. Journal of the American Statistical Association, 53(282):624, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281899.

Hood:1957:BRB

[Hoo57] Wm. C. Hood. Book review: [untitled]. Journal of the American Statistical Association, 52(277):115–117, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281420.

Hooke:1958:URI

[Hoo58] Robert Hooke. Use of randomization in the investigation of unknown functions. Journal of the American Statistical Association, 53(281):176–186, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282578.

Horwich:1955:BRB

[Hor55] George Horwich. Book review: [untitled]. Journal of the American Statistical Association, 50(269):258–260, March 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281113.

Horton:1957:BRB

[Hor57] H. Burke Horton. Book review: [untitled]. Journal of the American Statistical Association, 52(279):391, September 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280914.

Hoselitz:1956:BRB

[Hos56] Bert F. Hoselitz. Book review: [untitled]. Journal of the American Statistical Association, 51(276):696–698, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281519.

Hotelling:1951:IRF

[Hot51] Harold Hotelling. The impact of R. A. Fisher on statistics. Journal of the American Statistical Association, 46(253):35–46, March 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280091.

Householder:1953:BRD

[Hou53] A. S. Householder. Book review: Design for a Brain, by W. Ross Ashby. Journal of the American Statistical Association, 48 (263):669–671, September 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281028.

Houthakker:1954:BRB

[Hou54a] H. S. Houthakker. Book review: [untitled]. Journal of the American Statistical Association, 49(267):670–672, September 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281148.

Houthakker:1954:DA

[Hou54b] H. S. Houthakker. Demand analysis. Journal of the American Statistical Association, 49(265):88–96, March 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281037.

Houseman:1956:BRB

[Hou56] Earl E. Houseman. Book review: [untitled]. Journal of the American Statistical Association, 51(275):545–546, September 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281452.

Howard:1957:BRB

[How57] John A. Howard. Book review: [untitled]. Journal of the American Statistical Association, 52(280):605–607, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281732.

Howes:1958:BRB

[How58] David R. Howes. Book review: [untitled]. Journal of the American Statistical Association, 53(282):599–600, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281880.

Hauser:1950:WU

[HP50] Philip M. Hauser and Robert B. Pearl. Who are the unemployed? Journal of the American Statistical Association, 45(252):479–500, December 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280718.

Hughes:1958:FBS

[HR58] J. R. T. Hughes and Stanley Reiter. The first 1,945 British steamships. Journal of the American Statistical Association, 53 (282):360–381, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281861.

Hagood:1952:PPS

[HS52] Margaret Jarman Hagood and Jacob S. Siegel. Population projections for sales forecasting. Journal of the American Statistical Association, 47(259):524–540, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281320.

Horvitz:1952:GSR

[HT52] D. G. Horvitz and D. J. Thompson. A generalization of sampling without replacement from a finite universe. Journal of the American Statistical Association, 47(260):663–685, December 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280784.

Hughes:1952:BRB

[Hug52] Harry M. Hughes. Book review: [untitled]. Journal of the American Statistical Association, 47(260):692–694, December 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280789.

Hultgren:1956:BRB

[Hul56] Thor Hultgren. Book review: [untitled]. Journal of the American Statistical Association, 51(276):672–674, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281507.

Hutchinson:1958:NIS

[Hut58] E. P. Hutchinson. Notes on immigration statistics of the United States. Journal of the American Statistical Association, 53(284): 963–1025, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281963.

Hamaker:1955:EDS

[HV55] H. C. Hamaker and R. Van Strik. The efficiency of double sampling for attributes. Journal of the American Statistical Association, 50 (271):830–849, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281168.

Hyman:1957:BRB

[Hym57] Ray Hyman. Book review: [untitled]. Journal of the American Statistical Association, 52(280):607–610, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281733.

Hyman:1958:BRB

[Hym58] Ray Hyman. Book review: [untitled]. Journal of the American Statistical Association, 53(281):251–252, March 1958. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282612.

Hyrenius:1953:EUR

[Hyr53a] Hannes Hyrenius. Errata: On the use of ranges, cross-ranges and extremes in comparing small samples. Journal of the American Statistical Association, 48(264):907, December 1953. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281081.

Hyrenius:1953:URC

[Hyr53b] Hannes Hyrenius. On the use of ranges, cross-ranges and extremes in comparing small samples. Journal of the American Statistical Association, 48(263):534–545, September 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281007.

Ireson:1952:BRB

[Ire52] W. Grant Ireson. Book review: [untitled]. Journal of the American Statistical Association, 47(258):317–319, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280757.

Ireson:1958:BRB

[Ire58] W. Grant Ireson. Book review: [untitled]. Journal of the American Statistical Association, 53(282):625–626, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281901.

Isaacs:1957:BRB

[Isa57] Reginald R. Isaacs. Book review: [untitled]. Journal of the American Statistical Association, 52(278):287–289, June 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280871.

Jacobson:1953:BRB

[Jac53] Paul H. Jacobson. Book review: [untitled]. Journal of the American Statistical Association, 48(263):668–669, September 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281027.

Jacoby:1956:BRB

[Jac56] Neil H. Jacoby. Book review: [untitled]. Journal of the American Statistical Association, 51(276):690–693, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281516.

Jaffe:1952:BRB

[Jaf52] A. J. Jaffe. Book review: [untitled]. Journal of the American Statistical Association, 47(258):348–349, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280771.

Jaffe:1956:BRB

[Jaf56] A. J. Jaffe. Book review: [untitled]. Journal of the American Statistical Association, 51(273):190–192, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280827.

Jahoda:1954:BRB

[Jah54] Marie Jahoda. Book review: [untitled]. Journal of the American Statistical Association, 49(265):189, March 1954. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281048.

Jahn:1958:BRB

[Jah58] Julius A. Jahn. Book review: [untitled]. Journal of the American Statistical Association, 53(281):247–249, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282610.

James:1950:BRB

[Jam50] William C. James. Book review: [untitled]. Journal of the American Statistical Association, 45(251):456, September 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280302.

Jarrett:1956:BRB

[Jar56] R. F. Jarrett. Book review: [untitled]. Journal of the American Statistical Association, 51(275):560, September 1956. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281462.

Jardine:1958:RMM

[Jar58] R. Jardine. Ranking methods and the measurement of attitudes. Journal of the American Statistical Association, 53(283):720–728, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2282049.

Jasny:1950:IOS

[Jas50] Naum Jasny. International organizations and Soviet statistics. Journal of the American Statistical Association, 45(249):48–64, March 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280427.

Jastram:1951:WPP

[Jas51a] Roy W. Jastram. Willard Phillips, a predecessor of Paasche in index number formation. Journal of the American Statistical Association, 46(253):124–126, March 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280102.

Jaszi:1951:NIS

[Jas51b] George Jaszi. National income: Status and prospects as seen by an estimator. Journal of the American Statistical Association, 46 (255):345–357, September 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280512.

Jaspen:1956:MCH

[Jas56] Nathan Jaspen. Machine computation of higher moments. Journal of the American Statistical Association, 51(275):489–500, September 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2281438.

Jebe:1951:BRB

[Jeb51] Emil H. Jebe. Book review: [untitled]. Journal of the American Statistical Association, 46(254):259–261, June 1951. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2279929.

Jebe:1952:ESS

[Jeb52] Emil H. Jebe. Estimation for sub-sampling designs employing the county as a primary sampling unit. Journal of the American Statistical Association, 47(257):49–70, March 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279977.

Jackson:1957:AMQ

[JM57] J. Edward Jackson and Robert H. Morris. An application of multivariate quality control to photographic processing. Journal of the American Statistical Association, 52(278):186–199, June 1957. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280844.

Johnson:1951:CAF

[JN51] D. Gale Johnson and Marilyn Corn Nottenburg. A critical analysis of farm employment estimates. Journal of the American Statistical Association, 46(254):191–205, June 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279919.

Johnson:1950:BRB

[Joh50a] Palmer O. Johnson. Book review: [untitled]. Journal of the American Statistical Association, 45(251):454–455, September 1950. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280301.

Johnson:1950:QQD

[Joh50b] Palmer O. Johnson. The quantification of qualitative data in discriminant analysis. Journal of the American Statistical Association, 45(249):65–76, March 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280428.

Johnson:1954:BRB

[Joh54] Palmer O. Johnson. Book review: [untitled]. Journal of the American Statistical Association, 49(266):379–383, June 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280941.

Johnson:1956:BRBb

[Joh56a] D. Gale Johnson. Book review: [untitled]. Journal of the American Statistical Association, 51(276):698, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281520.

Johnson:1956:BRBc

[Joh56b] Howard W. Johnson. Book review: [untitled]. Journal of the American Statistical Association, 51(276):699–700, December 1956. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281521.

Johnson:1956:BRBa

[Joh56c] M. Clemens Johnson. Book review: [untitled]. Journal of the American Statistical Association, 51(274):413–414, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281384.

Jones:1951:BRB

[Jon51] Howard L. Jones. Book review: [untitled]. Journal of the American Statistical Association, 46(254):262–265, June 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279931.

Jones:1952:BRB

[Jon52] Lyle V. Jones. Book review: [untitled]. Journal of the American Statistical Association, 47(258):327–328, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280761.

Jones:1953:AMW

[Jon53] Howard L. Jones. Approximating the mode from weighted sample values. Journal of the American Statistical Association, 48 (261):113–127, March 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280884.

Jones:1955:ASP

[Jon55a] Howard L. Jones. The application of sampling procedures to business operations. Journal of the American Statistical Association, 50(271):763–774, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281164.

Jones:1955:CSA

[Jon55b] William O. Jones. Colonial social accounting. Journal of the American Statistical Association, 50(271):665–676, September 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281158.

Jones:1956:BRBb

[Jon56a] Frank E. Jones. Book review: [untitled]. Journal of the American Statistical Association, 51(275):541–543, September 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281449.

Jones:1956:IPS

[Jon56b] Howard L. Jones. Investigating the properties of a sample mean by employing random subsample means. Journal of the American Statistical Association, 51(273):54–83, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280804.

Jones:1956:BRBa

[Jon56c] Lyle V. Jones. Book review: [untitled]. Journal of the American Statistical Association, 51(273):169–170, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280815.

Jones:1957:BRB

[Jon57] Howard L. Jones. Book review: [untitled]. Journal of the American Statistical Association, 52(278):268–270, June 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280855.

Jones:1958:BRB

[Jon58a] Alice Hanson Jones. Book review: [untitled]. Journal of the American Statistical Association, 53(282):611–613, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281891.

Jones:1958:ISC

[Jon58b] Howard L. Jones. Inadmissible samples and confidence limits. Journal of the American Statistical Association, 53(282):482–490, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2281869.

Jones:1959:HMG

[Jon59] Howard L. Jones. How many of a group of random numbers will be usable in selecting a particular sample? Journal of the American Statistical Association, 54(285):102–122, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282142.

Joskow:1958:BRB

[Jos58] Jules Joskow. Book review: [untitled]. Journal of the American Statistical Association, 53(284):1033–1034, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281972.

Jackson:1955:CET

[JR55a] J. Edward Jackson and Eleanor L. Ross. Corrigenda: Extended tables for use with the “G” test for means. Journal of the American Statistical Association, 50(272):1332, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281223.

Jackson:1955:ETU

[JR55b] J. Edward Jackson and Eleanor L. Ross. Extended tables for use with the “G” test for means. Journal of the American Statistical Association, 50(270):416–433, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280970.

Kahn:1956:BRB

[Kah56] Robert L. Kahn. Book review: [untitled]. Journal of the American Statistical Association, 51(274):418–420, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281389.

Kaiser:1959:BRB

[Kai59a] Henry F. Kaiser. Book review: [untitled]. Journal of the American Statistical Association, 54(287):699–700, September 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282549.

Kaitz:1959:CGE

[Kai59b] Hyman B. Kaitz. A check on gross errors in certain variance computations. Journal of the American Statistical Association, 54 (288):741–743, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282497.

Kallianpur:1951:BRB

[Kal51] Gopinath Kallianpur. Book review: [untitled]. Journal of the American Statistical Association, 46(256):535–536, December 1951. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280417.

Kaplan:1959:BRB

[Kap59a] David L. Kaplan. Book review: [untitled]. Journal of the American Statistical Association, 54(285):329–330, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282169.

Kaplan:1959:SMN

[Kap59b] Norman M. Kaplan. Some methodological notes on the deflation of construction. Journal of the American Statistical Association, 54 (287):535–555, September 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282534.

Karmel:1950:NPK

[Kar50] P. H. Karmel. A note on P. K. Whelpton’s calculation of parity adjusted reproduction rates. Journal of the American Statistical Association, 45(249):119–124, March 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280434.

Karlin:1956:BRB

[Kar56] Samuel Karlin. Book review: [untitled]. Journal of the American Statistical Association, 51(274):391–392, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281366.

Karlin:1958:BRB

[Kar58a] Samuel Karlin. Book review: [untitled]. Journal of the American Statistical Association, 53(283):761–762, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282064.

Karpinos:1958:WHS

[Kar58b] Bernard D. Karpinos. Weight-height standards based on world war II experience. Journal of the American Statistical Association, 53(282):408–419, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281864.

Karst:1958:LCF

[Kar58c] Otto J. Karst. Linear curve fitting using least deviations. Journal of the American Statistical Association, 53(281):118–132, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282572.

Katz:1953:CIN

[Kat53] Leo Katz. Confidence intervals for the number showing a certain characteristic in a population when sampling is without replacement. Journal of the American Statistical Association, 48(262):256– 261, June 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281286.

Katz:1955:BRB

[Kat55] Leo Katz. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1345–1346, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281235.

Katz:1956:BRBa

[Kat56a] Leo Katz. Book review: [untitled]. Journal of the American Statistical Association, 51(273):165–166, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280813.

Katz:1956:BRBb

[Kat56b] Leo Katz. Book review: [untitled]. Journal of the American Statistical Association, 51(274):380–381, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281358.

Katz:1956:BRBc

[Kat56c] Leo Katz. Book review: [untitled]. Journal of the American Statistical Association, 51(274):416, June 1956. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281386.

Kempthorne:1953:PER

[KB53] O. Kempthorne and W. D. Barclay. The partition of error in randomized blocks. Journal of the American Statistical Association, 48(263):610–614, September 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281012.

Kempthorne:1954:BRB

[Kem54] Oscar Kempthorne. Book review: [untitled]. Journal of the American Statistical Association, 49(266):383–385, June 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280942.

Kempthorne:1955:RTE

[Kem55] Oscar Kempthorne. The randomization theory of experimental inference. Journal of the American Statistical Association, 50(271): 946–967, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281178.

kempthorne:1956:CRT

[kem56] Oscar kempthorne. Corrigenda: The randomization theory of experimentals inference. Journal of the American Statistical Association, 51(276):651, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281484.

Kendall:1952:BRB

[Ken52] M. G. Kendall. Book review: [untitled]. Journal of the American Statistical Association, 47(257):99–100, March 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279986.

Kendrick:1956:BRBa

[Ken56a] John W. Kendrick. Book review: [untitled]. Journal of the American Statistical Association, 51(273):178–180, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280820.

Kendrick:1956:BRBb

[Ken56b] John W. Kendrick. Book review: [untitled]. Journal of the American Statistical Association, 51(275):546–548, September 1956. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281453.

Kent:1957:BRBa

[Ken57a] Leonard Kent. Book review: [untitled]. Journal of the American Statistical Association, 52(278):273–276, June 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280861.

Kent:1957:BRBb

[Ken57b] Leonard Kent. Book review: [untitled]. Journal of the American Statistical Association, 52(280):612–614, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281735.

Kenen:1958:BRB

[Ken58] Peter B. Kenen. Book review: [untitled]. Journal of the American Statistical Association, 53(282):613–614, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281892.

Kershaw:1956:BRB

[Ker56] Joseph A. Kershaw. Book review: [untitled]. Journal of the American Statistical Association, 51(274):407–409, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281379.

Kester:1956:BRB

[Kes56] Henry I. Kester. Book review: [untitled]. Journal of the American Statistical Association, 51(273):183–184, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280823.

Keyfitz:1951:SPP

[Key51] Nathan Keyfitz. Sampling with probabilities proportional to size: Adjustment for changes in the probabilities. Journal of the American Statistical Association, 46(253):105–109, March 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280099.

Keyfitz:1957:ESV

[Key57] Nathan Keyfitz. Estimates of sampling variance where two units are selected from each stratum. Journal of the American Statistical Association, 52(280):503–510, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281699.

Kish:1958:NSD

[KH58] Leslie Kish and Irene Hess. On noncoverage of sample dwellings. Journal of the American Statistical Association, 53(282):509–524, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2281871.

Kish:1959:VRT

[KH59] Leslie Kish and Irene Hess. On variances of ratios and their differences in multi-stage samples. Journal of the American Statistical Association, 54(286):416–446, June 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281782. See corrigenda [KH63].

Kish:1963:CVR

[KH63] Leslie Kish and Irene Hess. Corrigenda: On Variances of Ratios and Their Differences in Multi-Stage Samples. Journal of the American Statistical Association, 58(304):1162, December 1963. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2283347. See [KH59].

Kinsey:1955:CMT

[KHS^{+}55] Alfred C. Kinsey, Herbert Hyman, Paul B. Sheatsley, A. H. Hobbs, R. D. Lambert, Nicholas Pastore, Jacob Goldstein, Lewis M. Terman, Paul Wallin, W. Allen Wallis, William G. Cochran, Frederick Mosteller, and John W. Tukey. The Cochran–Mosteller–Tukey Report on the Kinsey Study: A Symposium. Journal of the American Statistical Association, 50(271):811–829, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281167.

Kimball:1953:MGL

[Kim53] Bradford F. Kimball. A multiple group least squares’ problem and the significance of the associated orthogonal polynomials. Journal of the American Statistical Association, 48(262):320–335, June 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2281290.

Kimball:1955:CPA

[Kim55a] Bradford F. Kimball. Corrigenda: Practical applications of the theory of extreme values. Journal of the American Statistical Association, 50(272):1332, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281224.

Kimball:1955:PAT

[Kim55b] Bradford F. Kimball. Practical applications of the theory of extreme values. Journal of the American Statistical Association, 50 (270):517–528, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280973.

Kimball:1956:BRB

[Kim56] Bradford F. Kimball. Book review: [untitled]. Journal of the American Statistical Association, 51(273):176–177, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280819.

Kimball:1957:ETK

[Kim57] A. W. Kimball. Errors of the third kind in statistical consulting. Journal of the American Statistical Association, 52(278):133–142, June 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280840.

King:1953:SPR

[Kin53] E. P. King. On some procedures for the rejection of suspected data. Journal of the American Statistical Association, 48(263):531–533, September 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281006.

King:1954:OGO

[Kin54] E. P. King. Optimum grouping in one-criterion variance components analysis. Journal of the American Statistical Association, 49 (267):637–639, September 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281134.

Kiser:1952:FTD

[Kis52] Clyde V. Kiser. Fertility trends and differentials in the United States. Journal of the American Statistical Association, 47(257):25– 48, March 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2279976.

Kitagawa:1955:CDB

[Kit55] Evelyn M. Kitagawa. Components of a difference between two rates. Journal of the American Statistical Association, 50(272):1168–1194, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281213.

Kitagawa:1956:CCD

[Kit56] Evelyn M. Kitagawa. Corrigenda: Components of a difference between two rates. Journal of the American Statistical Association, 51(276):651, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281485.

Kish:1954:REE

[KL54] Leslie Kish and John B. Lansing. Response errors in estimating the value of homes. Journal of the American Statistical Association, 49 (267):520–538, September 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281128.

Klaman:1958:BRB

[Kla58] Saul B. Klaman. Book review: [untitled]. Journal of the American Statistical Association, 53(282):606–609, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281888.

Klein:1952:BRB

[Kle52] L. R. Klein. Book review: [untitled]. Journal of the American Statistical Association, 47(258):332–334, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280764.

Klebaner:1957:BRB

[Kle57] Benjamin J. Klebaner. Book review: [untitled]. Journal of the American Statistical Association, 52(280):602–603, December 1957.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281730.

Klein:1951:RAS

[KM51] L. R. Klein and J. N. Morgan. Results of alternative statistical treatments of sample survey data. Journal of the American Statistical Association, 46(256):442–460, December 1951. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2280395.

Kaplan:1958:NEI

[KM58] E. L. Kaplan and Paul Meier. Nonparametric estimation from incomplete observations. Journal of the American Statistical Association, 53(282):457–481, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281868.

Knight:1950:BRB

[Kni50] M. M. Knight. Book review: [untitled]. Journal of the American Statistical Association, 45(251):464–465, September 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280306.

Knowler:1954:BRB

[Kno54] Lloyd A. Knowler. Book review: [untitled]. Journal of the American Statistical Association, 49(268):916–919, December 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281554.

Koch:1952:BRB

[Koc52] Albert R. Koch. Book review: [untitled]. Journal of the American Statistical Association, 47(260):700–702, December 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280793.

Kohn:1956:BRB

[Koh56] Robert Kohn. Book review: [untitled]. Journal of the American Statistical Association, 51(273):203–204, March 1956. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2280834.

Konijn:1952:BRB

[Kon52] H. S. Konijn. Book review: [untitled]. Journal of the American Statistical Association, 47(259):575–579, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281338.

Kossack:1953:BRB

[Kos53] Carl F. Kossack. Book review: [untitled]. Journal of the American Statistical Association, 48(261):153–154, March 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280888.

Kozelka:1955:BRB

[Koz55] Robert M. Kozelka. Book review: [untitled]. Journal of the American Statistical Association, 50(270):607–608, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280979.

Karlin:1956:DPM

[KR56] Samuel Karlin and H. Rubin. Distributions possessing a monotone likelihood ratio. Journal of the American Statistical Association, 51(276):637–643, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281477.

Kramer:1959:BRB

[Kra59] Clyde Y. Kramer. Book review: [untitled]. Journal of the American Statistical Association, 54(288):816, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282518.

Kruskal:1951:BRB

[Kru51] William Kruskal. Book review: [untitled]. Journal of the American Statistical Association, 46(255):410–412, September 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280536.

Krumbein:1954:ASM

[Kru54] W. C. Krumbein. Applications of statistical methods to sedimentary rocks. Journal of the American Statistical Association, 49 (265):51–66, March 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281034.

Kruskal:1957:HNW

[Kru57] William H. Kruskal. Historical notes on the Wilcoxon unpaired two-sample test. Journal of the American Statistical Association, 52(279):356–360, September 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280906.

Kruskal:1958:BRB

[Kru58a] William H. Kruskal. Book review: [untitled]. Journal of the American Statistical Association, 53(282):595–598, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281878.

Kruskal:1958:OMA

[Kru58b] William H. Kruskal. Ordinal measures of association. Journal of the American Statistical Association, 53(284):814–861, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281954.

Krutilla:1959:BRB

[Kru59] John V. Krutilla. Book review: [untitled]. Journal of the American Statistical Association, 54(285):325–327, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282166.

Kullback:1958:BRB

[Kul58] S. Kullback. Book review: [untitled]. Journal of the American Statistical Association, 53(284):1034–1035, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281973.

Kurnow:1951:BRB

[Kur51] Ernest Kurnow. Book review: [untitled]. Journal of the American Statistical Association, 46(256):524–525, December 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280409.

Kuznets:1950:BRB

[Kuz50a] Simon Kuznets. Book review: [untitled]. Journal of the American Statistical Association, 45(252):576–579, December 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280732.

Kuznets:1950:CSR

[Kuz50b] Simon Kuznets. Conditions of statistical research. Journal of the American Statistical Association, 45(249):1–14, March 1950. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280424.

Kruskal:1952:URO

[KW52] William H. Kruskal and W. Allen Wallis. Use of ranks in onecriterion variance analysis. Journal of the American Statistical Association, 47(260):583–621, December 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280779.

Kruskal:1953:EUR

[KW53] William H. Kruskal and W. Allen Wallis. Errata: Use of ranks in one-criterion variance analysis. Journal of the American Statistical Association, 48(264):907–911, December 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281082.

L:1951:BRBa

[L.51a] M. A. L. Book review: [untitled]. Journal of the American Statistical Association, 46(256):536, December 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280418.

L:1951:BRBb

[L.51b] M. A. L. Book review: [untitled]. Journal of the American Statistical Association, 46(256):536, December 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280419.

L:1951:BRBc

[L.51c] M. A. L. Book review: [untitled]. Journal of the American Statistical Association, 46(256):537, December 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280420.

L:1954:BRBa

[L.54a] M. A. L. Book review: [untitled]. Journal of the American Statistical Association, 49(265):202, March 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281058.

L:1954:BRBb

[L.54b] M. A. L. Book review: [untitled]. Journal of the American Statistical Association, 49(265):203, March 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281060.

L:1955:BRB

[L.55] M. A. L. Book review: [untitled]. Journal of the American Statistical Association, 50(269):264, March 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281117.

Lacey:1956:BRB

[Lac56] Oliver L. Lacey. Book review: [untitled]. Journal of the American Statistical Association, 51(274):377–379, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281356.

Lah:1956:AGF

[Lah56] Ivo Lah. Analytical graduation of fertility rates. Journal of the American Statistical Association, 51(275):461–466, September 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2281435.

Lang:1956:BRB

[Lan56] Kurt Lang. Book review: [untitled]. Journal of the American Statistical Association, 51(273):197–199, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280831.

Landis:1959:CGL

[Lan59a] Benson Y. Landis. Corrigenda: “A Guide to the Literature on Statistics of Religious Affiliations with References to Related Social Studies”. Journal of the American Statistical Association, 54 (288):811, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282511. See [Lan59b].

Landis:1959:GLS

[Lan59b] Benson Y. Landis. A guide to the literature on statistics of religious affiliation with references to related social studies. Journal of the American Statistical Association, 54(286):335–357, June 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2281776. See cor-

rigenda [Lan59a].

Laurent:1955:BRBa

[Lau55a] Andre G. Laurent. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1338, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281231.

Laurent:1955:BRBb

[Lau55b] Andre G. Laurent. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1361–1364, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281243.

Lansing:1958:CSA

[LB58] John B. Lansing and Dwight M. Blood. A cross-section analysis of non-business air travel. Journal of the American Statistical Association, 53(284):928–947, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281961.

Lebergott:1954:MEM

[Leb54] Stanley Lebergott. Measurement for economic models. Journal of the American Statistical Association, 49(266):209–226, June 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280930.

Lee:1951:BRB

[Lee51] Ivan M. Lee. Book review: [untitled]. Journal of the American Statistical Association, 46(255):398–402, September 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280528.

Lee:1952:CEA

[Lee52] Ivan M. Lee. A critical evaluation of available agricultural statistics. Journal of the American Statistical Association, 47(258):267–280, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280751.

Lee:1953:BRB

[Lee53] Ivan M. Lee. Book review: [untitled]. Journal of the American Statistical Association, 48(263):663–665, September 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281023.

Lee:1956:BRBb

[Lee56a] Ivan M. Lee. Book review: [untitled]. Journal of the American Statistical Association, 51(274):405–407, June 1956. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281377.

Lee:1956:BRBa

[Lee56b] Maurice W. Lee. Book review: [untitled]. Journal of the American Statistical Association, 51(274):393, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281368.

Lee:1959:BRB

[Lee59] Ivan M. Lee. Book review: [untitled]. Journal of the American Statistical Association, 54(286):516–518, June 1959. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281793.

Lehmann:1953:BRB

[Leh53] E. L. Lehmann. Book review: [untitled]. Journal of the American Statistical Association, 48(262):358–359, June 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281293.

LeNeveu:1956:BRB

[LeN56] A. H. LeNeveu. Book review: [untitled]. Journal of the American Statistical Association, 51(274):410–411, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281381.

Leong:1950:IPV

[Leo50] Y. S. Leong. Index of the physical volume production of minerals, 1880–1948. Journal of the American Statistical Association, 45(249):15–29, March 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280425.

Leonard:1958:OR

[Leo58] William R. Leonard. An outlook report. Journal of the American Statistical Association, 53(281):1–10, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282560.

Lerner:1957:BRB

[Ler57] William Lerner. Book review: [untitled]. Journal of the American Statistical Association, 52(280):580–581, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281711.

Levy:1957:BRB

[Lev57] S. Jay Levy. Book review: [untitled]. Journal of the American Statistical Association, 52(279):393–396, September 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280917.

Lewis-Faning:1954:BRB

[LF54] E. Lewis-Faning. Book review: [untitled]. Journal of the American Statistical Association, 49(265):190–193, March 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281050.

Li:1955:DRS

[Li55] C. C. Li. A diagrammatic representation of the sum of squares and products. Journal of the American Statistical Association, 50 (272):1056–1063, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281204.

Lieberman:1952:MIS

[Lie52] Gerald J. Lieberman. Multistation inspection schemes. Journal of the American Statistical Association, 47(259):466–474, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2281316.

Lieberman:1953:BRB

[Lie53] G. J. Lieberman. Book review: [untitled]. Journal of the American Statistical Association, 48(262):365–366, June 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281297.

Lieberman:1955:BRB

[Lie55] Gerald J. Lieberman. Book review: [untitled]. Journal of the American Statistical Association, 50(270):611–612, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280982.

Lieberman:1958:BRB

[Lie58a] Gerald J. Lieberman. Book review: [untitled]. Journal of the American Statistical Association, 53(284):1035, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281974.

Lieberman:1958:PSP

[Lie58b] Milton D. Lieberman. Philippine statistical program development and the survey of households. Journal of the American Statistical Association, 53(281):78–88, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282567.

Lindley:1954:BRB

[Lin54] D. V. Lindley. Book review: [untitled]. Journal of the American Statistical Association, 49(267):647–649, September 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281136.

Lindley:1958:BRB

[Lin58] Dennis V. Lindley. Book review: [untitled]. Journal of the American Statistical Association, 53(284):1049–1050, December 1958. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281987.

Lindley:1959:BRB

[Lin59] D. V. Lindley. Book review: [untitled]. Journal of the American Statistical Association, 54(288):825–827, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282528.

Laderman:1953:IP

[LL53] J. Laderman and S. B. Littauer. The inventory problem. Journal of the American Statistical Association, 48(264):717–732, December 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2281067.

Laderman:1954:EIP

[LLT54] J. Laderman, S. B. Littauer, and John W. Tukey. Errata: The inventory problem. Journal of the American Statistical Association, 49(268):906, December 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281550.

Locks:1954:BRB

[Loc54] Mitchell O. Locks. Book review: [untitled]. Journal of the American Statistical Association, 49(265):193–195, March 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281051.

Locks:1959:APA

[Loc59] Mitchell O. Locks. Automatic programming for automatic computers. Journal of the American Statistical Association, 54(288): 744–754, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282498.

Loeve:1951:BRB

[Loe51] M. Loeve. Book review: [untitled]. Journal of the American Statistical Association, 46(255):384–385, September 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280519.

Loftus:1950:BRB

[Lof50] P. J. Loftus. Book review: [untitled]. Journal of the American Statistical Association, 45(251):465–466, September 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280307.

Lomax:1954:BFA

[Lom54] K. S. Lomax. Business failures: Another example of the analysis of failure data. Journal of the American Statistical Association, 49 (268):847–852, December 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281544.

Long:1956:BRB

[Lon56] Clarence D. Long. Book review: [untitled]. Journal of the American Statistical Association, 51(273):192–194, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280828.

Lorimer:1952:BRBa

[Lor52a] Frank Lorimer. Book review: [untitled]. Journal of the American Statistical Association, 47(258):345–346, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280769.

Lorimer:1952:BRBb

[Lor52b] Frank Lorimer. Book review: [untitled]. Journal of the American Statistical Association, 47(258):349–350, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280772.

Lord:1953:BRB

[Lor53] Frederic M. Lord. Book review: [untitled]. Journal of the American Statistical Association, 48(264):913–915, December 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281087.

Lord:1955:EPI

[Lor55a] Frederic M. Lord. Estimation of parameters from incomplete data. Journal of the American Statistical Association, 50(271):870–876, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281171.

Lord:1955:NCM

[Lor55b] Frederic M. Lord. Nomograph for computing multiple correlation coefficients. Journal of the American Statistical Association, 50 (272):1073–1077, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281206.

Lord:1959:PMT

[Lor59] Frederic M. Lord. Problems in mental test theory arising from errors of measurement. Journal of the American Statistical Association, 54(286):472–479, June 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281785.

Lowry:1958:BRB

[Low58] Dorothy C. Lowry. Book review: [untitled]. Journal of the American Statistical Association, 53(283):752–753, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282055.

Lieberman:1955:CSP

[LR55a] Gerald J. Lieberman and George J. Resnikoff. Corrigenda: Sampling plans for inspection by variables. Journal of the American Statistical Association, 50(272):1333, December 1955. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281227.

Lieberman:1955:SPI

[LR55b] Gerald J. Lieberman and George J. Resnikoff. Sampling plans for inspection by variables. Journal of the American Statistical Association, 50(270):457–516, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280972.

M:1954:BRB

[M.54] F. M. Book review: [untitled]. Journal of the American Statistical Association, 49(268):926–927, December 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281559.

M:1959:BRB

[M.59] W. G. M. Book review: [untitled]. Journal of the American Statistical Association, 54(288):820, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282521.

Maccoby:1952:BRB

[Mac52] Nathan Maccoby. Book review: [untitled]. Journal of the American Statistical Association, 47(260):705–706, December 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280795.

Macura:1955:BRB

[Mac55] Milos Macura. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1409–1411, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281271.

Mack:1956:BRB

[Mac56] Ruth P. Mack. Book review: [untitled]. Journal of the American Statistical Association, 51(274):395–396, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281370.

Maccoby:1959:BRB

[Mac59a] Eleanor E. Maccoby. Book review: [untitled]. Journal of the American Statistical Association, 54(287):702–704, September 1959. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282551.

MacKinnon:1959:CTT

[Mac59b] William J. MacKinnon. Compact table of twelve probability levels of the symmetric binomial cumulative distribution for sample sizes to 1,000. Journal of the American Statistical Association, 54(285):164–172, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor. org/stable/2282144. See corrigenda [Mac59c].

MacKinnon:1959:CCT

[Mac59c] William J. MacKinnon. Corrigenda: “Compact Table of Twelve Probability Levels of the Symmetric Binomial Comulative Distribution for Sample Sizes to 1,000”. Journal of the American Statistical Association, 54(288):811, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282512. See [Mac59b].

Madow:1950:UCP

[Mad50] Lillian H. Madow. On the use of the county as the primary sampling unit for state estimates. Journal of the American Statistical Association, 45(249):30–47, March 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280426.

Madow:1951:MPA

[Mad51] William G. Madow. The measurement and prediction of attitudes. Journal of the American Statistical Association, 46(256):508–514, December 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280400.

Madow:1953:BRB

[Mad53] William G. Madow. Book review: [untitled]. Journal of the American Statistical Association, 48(263):654–655, September 1953. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281018.

Madow:1955:BRB

[Mad55] William G. Madow. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1346–1348, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281236.

Madow:1958:BRBa

[Mad58a] William G. Madow. Book review: [untitled]. Journal of the American Statistical Association, 53(281):208–212, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282581.

Madow:1958:BRBb

[Mad58b] William G. Madow. Book review: [untitled]. Journal of the American Statistical Association, 53(283):774, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282073.

Madow:1958:BRBc

[Mad58c] William G. Madow. Book review: [untitled]. Journal of the American Statistical Association, 53(284):1036–1037, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281976.

Madansky:1959:CFS

[Mad59a] Albert Madansky. Corrigenda: “The Fitting of Straight Lines When Both Variables are Subject to Error”. Journal of the American Statistical Association, 54(288):812, December 1959. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282513. See [Mad59b].

Madansky:1959:FSL

[Mad59b] Albert Madansky. The fitting of straight lines when both variables are subject to error. Journal of the American Statistical Association, 54(285):173–205, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor. org/stable/2282145. See corrigenda [Mad59a].

Maisel:1954:BRB

[Mai54] Sherman J. Maisel. Book review: [untitled]. Journal of the American Statistical Association, 49(266):399, June 1954. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2280949.

Mair:1956:BRBa

[Mai56a] George F. Mair. Book review: [untitled]. Journal of the American Statistical Association, 51(273):199–201, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280832.

Mair:1956:BRBb

[Mai56b] George F. Mair. Book review: [untitled]. Journal of the American Statistical Association, 51(276):706–708, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281527.

Mainland:1958:BRB

[Mai58] Donald Mainland. Book review: [untitled]. Journal of the American Statistical Association, 53(283):772–773, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282072.

Malzberg:1950:HMP

[Mal50] Benjamin Malzberg. Horatio Milo Pollock, 1868–1950. Journal of the American Statistical Association, 45(251):452–453, September 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280300.

Malinowski:1953:BRB

[Mal53] Z. S. Malinowski. Book review: [untitled]. Journal of the American Statistical Association, 48(263):647–650, September 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281015.

Mann:1952:BRB

[Man52] H. B. Mann. Book review: [untitled]. Journal of the American Statistical Association, 47(258):351, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280774.

Mandel:1953:SFO

[Man53] B. J. Mandel. Sampling the Federal old-age and survivors insurance records. Journal of the American Statistical Association, 48 (263):462–475, September 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281002.

Mandel:1957:FSL

[Man57] John Mandel. Fitting a straight line to certain types of cumulative data. Journal of the American Statistical Association, 52 (280):552–566, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281706.

Mann:1958:BRB

[Man58] Irwin Mann. Book review: [untitled]. Journal of the American Statistical Association, 53(283):758–761, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282062.

Marshall:1951:LST

[Mar51] Andrew W. Marshall. A large-sample test of the hypothesis that one of two random variables is stochastically larger than the other. Journal of the American Statistical Association, 46(255):366–374, September 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280514.

Markowitz:1952:BRB

[Mar52] Harry Markowitz. Book review: [untitled]. Journal of the American Statistical Association, 47(259):557–558, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281328.

Marshall:1953:BRB

[Mar53] A. W. Marshall. Book review: [untitled]. Journal of the American Statistical Association, 48(262):362–363, June 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281295.

Marshall:1955:CNN

[Mar55] Herbert Marshall. Canada — northern neighbor. Journal of the American Statistical Association, 50(269):1–15, March 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281096.

Marshall:1956:BRB

[Mar56] A. W. Marshall. Book review: [untitled]. Journal of the American Statistical Association, 51(276):708–709, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281528.

Markham:1957:BRB

[Mar57] Jesse W. Markham. Book review: [untitled]. Journal of the American Statistical Association, 52(278):285–286, June 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280869.

Margolis:1958:BRB

[Mar58a] Julius Margolis. Book review: [untitled]. Journal of the American Statistical Association, 53(284):1045–1046, December 1958. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281984.

Marshall:1958:BRB

[Mar58b] A. W. Marshall. Book review: [untitled]. Journal of the American Statistical Association, 53(284):1044–1045, December 1958. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281983.

Massey:1951:KST

[Mas51] Frank J. Massey, Jr. The Kolmogorov–Smirnov test for goodness of fit. Journal of the American Statistical Association, 46(253):68–78, March 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280095.

Maslowski:1959:BRB

[Mas59] James J. Maslowski. Book review: [untitled]. Journal of the American Statistical Association, 54(286):520–521, June 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281795.

Mather:1951:RFS

[Mat51] Kenneth Mather. R. A. Fisher’s statistical methods for research workers: An appreciation. Journal of the American Statistical Association, 46(253):51–54, March 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280093.

Matchett:1957:BRB

[Mat57] Gerald J. Matchett. Book review: [untitled]. Journal of the American Statistical Association, 52(279):397–399, September 1957. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280919.

Mauldin:1958:BRB

[Mau58] W. Parker Mauldin. Book review: [untitled]. Journal of the American Statistical Association, 53(282):618–620, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281896.

Mayne:1950:BRB

[May50] Alvin Mayne. Book review: [untitled]. Journal of the American Statistical Association, 45(251):467–468, September 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280308.

Mayer:1955:BRB

[May55] Kurt Mayer. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1371–1372, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281248.

Mayer:1958:BRB

[May58] Harold M. Mayer. Book review: [untitled]. Journal of the American Statistical Association, 53(284):1053–1054, December 1958. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281990.

McAllister:1959:BRB

[McA59] Harry E. McAllister. Book review: [untitled]. Journal of the American Statistical Association, 54(287):707–709, September 1959. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282554.

McCarthy:1957:BRB

[McC57] Philip J. McCarthy. Book review: [untitled]. Journal of the American Statistical Association, 52(280):599, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281728.

McDaniel:1952:BRB

[McD52] Henry B. McDaniel. Book review: [untitled]. Journal of the American Statistical Association, 47(257):106–108, March 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279991.

McGee:1958:BRB

[McG58] John S. McGee. Book review: [untitled]. Journal of the American Statistical Association, 53(282):604–606, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281886.

McIvor:1955:BRB

[McI55] R. Craig McIvor. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1401–1403, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281267.

McIntyre:1957:BRB

[McI57] Francis McIntyre. Book review: [untitled]. Journal of the American Statistical Association, 52(278):265–267, June 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280853.

McLaughlin:1951:SME

[McL51] Glenn E. McLaughlin. Statistical measurement and economic mobilization planning. Journal of the American Statistical Association, 46(255):358–365, September 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280513.

McPherson:1958:BRB

[McP58] W. W. McPherson. Book review: [untitled]. Journal of the American Statistical Association, 53(281):240–241, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282605.

Meier:1954:ASL

[Mei54] Paul Meier. Analysis of simple lattice designs with unequal sets of replications. Journal of the American Statistical Association, 49 (268):786–813, December 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281540.

Meiselman:1958:BRB

[Mei58] David I. Meiselman. Book review: [untitled]. Journal of the American Statistical Association, 53(281):233–235, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282601.

Mendelson:1955:SM

[Men55] Morris Mendelson. A structure of moneyflows. Journal of the American Statistical Association, 50(269):72–92, March 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281099.

Merrell:1950:BRB

[Mer50] Margaret Merrell. Book review: [untitled]. Journal of the American Statistical Association, 45(250):294–295, June 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280692.

Merrell:1951:BRB

[Mer51] Margaret Merrell. Book review: [untitled]. Journal of the American Statistical Association, 46(256):528–529, December 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280412.

Metropolis:1950:BRB

[Met50a] N. Metropolis. Book review: Giant Brains or Machines That Think, by Edmund C. Berkeley. Journal of the American Statistical Association, 45(252):573–574, December 1950. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://links.jstor.org/sici?sici=0162-1459%
28195012%2945%3A252%3C573%3AGBOMTT%3E2.0.CO%3B2-U;   http:
//www.jstor.org/stable/2280729.

Metzner:1950:ASQ

[Met50b] Charles A. Metzner. An application of scaling to questionnaire construction. Journal of the American Statistical Association, 45 (249):112–118, March 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280433.

Meyer:1959:BRB

[Mey59] John Meyer. Book review: [untitled]. Journal of the American Statistical Association, 54(285):316–317, March 1959. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282157.

Marshall:1955:AMP

[MG55] Andrew W. Marshall and Herbert Goldhamer. An application of Markov processes to the study of the epidemiology of mental disease. Journal of the American Statistical Association, 50(269):99– 129, March 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281101.

Michael:1959:BRB

[Mic59a] Walther P. Michael. Book review: [untitled]. Journal of the American Statistical Association, 54(288):820–822, December 1959. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282523.

Mickey:1959:SFP

[Mic59b] M. R. Mickey. Some finite population unbiased ratio and regression estimators. Journal of the American Statistical Association, 54(287):594–612, September 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor. org/stable/2282540. See corrigenda [Mic60].

Mickey:1960:CSF

[Mic60] M. R. Mickey. Corrigenda: Some finite population unbiased ratio and regression estimators. Journal of the American Statistical Association, 55(292):755, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281602. See [Mic59b].

Miller:1951:CID

[Mil51] Herman P. Miller. Changes in income distribution in the United States. Journal of the American Statistical Association, 46(256):

438–441, December 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280394.

Miller:1953:ACI

[Mil53] Herman P. Miller. An appraisal of the 1950 census income data. Journal of the American Statistical Association, 48(261):28–43, March 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280877.

Miller:1955:ESS

[Mil55] Herman P. Miller. Elements of symmetry in the skewed income curve. Journal of the American Statistical Association, 50(269):55– 71, March 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281098.

Miller:1956:TPP

[Mil56a] Leslie H. Miller. Table of percentage points of Kolmogorov statistics. Journal of the American Statistical Association, 51(273):111– 121, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280807.

Miller:1956:BRB

[Mil56b] M. H. Miller. Book review: [untitled]. Journal of the American Statistical Association, 51(275):557–558, September 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281460.

Mills:1956:BRBa

[Mil56c] Edwin S. Mills. Book review: [untitled]. Journal of the American Statistical Association, 51(274):376–377, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281354.

Mills:1956:BRBb

[Mil56d] Edwin S. Mills. Book review: [untitled]. Journal of the American Statistical Association, 51(276):660, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281497.

Mincer:1957:ANG

[Min57a] Jacob Mincer. Applications of a new graphic method in statistical measurement. Journal of the American Statistical Association, 52 (280):472–478, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281695.

Mincer:1957:BRB

[Min57b] Jacob Mincer. Book review: [untitled]. Journal of the American Statistical Association, 52(278):267–268, June 1957. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2280854.

Mitten:1951:BRB

[Mit51] Loring G. Mitten. Book review: [untitled]. Journal of the American Statistical Association, 46(256):521–522, December 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280405.

Mitra:1957:TTL

[Mit57a] Sujit Kumar Mitra. Tables for tolerance limits for a normal population based on sample mean and range or mean range. Journal of the American Statistical Association, 52(277):88–94, March 1957. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281402.

Mitten:1957:BRB

[Mit57b] L. G. Mitten. Book review: [untitled]. Journal of the American Statistical Association, 52(279):396–397, September 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280918.

Marks:1950:REC

[MM50] Eli S. Marks and W. Parker Mauldin. Response errors in census research. Journal of the American Statistical Association, 45 (251):424–438, September 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280297.

Marks:1953:PES

[MMN53] Eli S. Marks, W. Parker Mauldin, and Harold Nisselson. The postenumeration survey of the 1950 Census: a case history in survey design. Journal of the American Statistical Association, 48(262):220– 243, June 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281284.

Mode:1956:BRB

[Mod56] Elmer B. Mode. Book review: [untitled]. Journal of the American Statistical Association, 51(274):377, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281355.

Mood:1951:EDS

[Moo51] A. M. Mood. Erratum: On the determination of sample sizes in designing experiments. Journal of the American Statistical Association, 46(256):515, December 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280401.

Moore:1955:PMS

[Moo55] P. G. Moore. The properties of the mean square successive difference in samples from various populations. Journal of the American Statistical Association, 50(270):434–456, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280971.

Moore:1956:CPM

[Moo56] P. G. Moore. Corrigenda: The properties of the mean square successive difference in samples from various populations. Journal of the American Statistical Association, 51(276):651, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281486.

Moonan:1957:LTS

[Moo57a] William J. Moonan. Linear transformation to a set of stochastically dependent normal variables. Journal of the American Statistical Association, 52(278):247–252, June 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280849.

Moore:1957:TNU

[Moo57b] P. G. Moore. Transformations to normality using fractional powers of the variable. Journal of the American Statistical Association, 52(278):237–246, June 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280848.

Moore:1958:MR

[Moo58] Geoffrey H. Moore. Measuring recessions. Journal of the American Statistical Association, 53(282):259–316, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281857.

Moriyama:1951:RMT

[Mor51a] Iwao M. Moriyama. Recent mortality trends and differentials. Journal of the American Statistical Association, 46(254):213–219, June 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2279921.

Morton:1951:BRB

[Mor51b] J. E. Morton. Book review: [untitled]. Journal of the American Statistical Association, 46(255):386–387, September 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280520.

Moran:1954:BRB

[Mor54] P. A. P. Moran. Book review: [untitled]. Journal of the American Statistical Association, 49(267):661–664, September 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281142.

Morris:1955:BRB

[Mor55] Morris David Morris. Book review: [untitled]. Journal of the American Statistical Association, 50(269):257–258, March 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281112.

Moranda:1959:CEC

[Mor59] P. B. Moranda. Comparison of estimates of circular probable error. Journal of the American Statistical Association, 54(288):794–800, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2282503.

See corrigenda [Mor60].

Moranda:1960:CCE

[Mor60] P. B. Moranda. Corrigenda: Comparison of Estimates of Circular Probable Error. Journal of the American Statistical Association, 55(292):755–756, December 1960. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281603. See [Mor59].

Moses:1951:BRB

[Mos51] Lincoln E. Moses. Book review: [untitled]. Journal of the American Statistical Association, 46(256):527–528, December 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280411.

Mosteller:1952:WSC

[Mos52] Frederick Mosteller. The world series competition. Journal of the American Statistical Association, 47(259):355–380, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2281309.

Moshman:1953:CVL

[Mos53] Jack Moshman. Critical values of the log-normal distribution. Journal of the American Statistical Association, 48(263):600–609, September 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281011.

Moser:1955:RDS

[Mos55] C. A. Moser. Recent developments in the sampling of human populations in Great Britain. Journal of the American Statistical Association, 50(272):1195–1214, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281214.

Moses:1956:BRB

[Mos56a] Lincoln E. Moses. Book review: [untitled]. Journal of the American Statistical Association, 51(276):704–705, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281525.

Moses:1956:STA

[Mos56b] Lincoln E. Moses. Some theoretical aspects of the lot plot sampling inspection plan. Journal of the American Statistical Association, 51(273):84–107, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280805.

Mudgett:1955:MSM

[Mud55] Bruce D. Mudgett. The measurement of seasonal movements in price and quantity indexes. Journal of the American Statistical Association, 50(269):93–98, March 1955. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281100.

Murphy:1955:BRB

[Mur55a] Henry C. Murphy. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1396–1397, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281264.

Murty:1955:DQM

[Mur55b] V. N. Murty. The distribution of the quotient of maximum values in samples from a rectangular distribution. Journal of the American Statistical Association, 50(272):1136–1141, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281210.

Muth:1958:BRB

[Mut58] Richard F. Muth. Book review: [untitled]. Journal of the American Statistical Association, 53(283):765–767, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282067.

Myers:1954:AAR

[Mye54a] Robert J. Myers. Accuracy of age reporting in the 1950 United States census. Journal of the American Statistical Association, 49 (268):826–831, December 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281542.

Myers:1954:FIM

[Mye54b] Robert J. Myers. Factors in interpreting mortality after retirement. Journal of the American Statistical Association, 49(267):499–509, September 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281126.

Nathan:1957:BRB

[Nat57] Robert R. Nathan. Book review: [untitled]. Journal of the American Statistical Association, 52(278):279–280, June 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280864.

Notestein:1950:PR

[NCD50] Frank W. Notestein, P. V. Cardon, and Joseph S. Davis. Population and resources. Journal of the American Statistical Association, 45 (251):346–349, September 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280292.

Neter:1952:SAS

[Net52] John Neter. Some applications of statistics for auditing. Journal of the American Statistical Association, 47(257):6–24, March 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279975.

Neter:1958:BRB

[Net58] John Neter. Book review: [untitled]. Journal of the American Statistical Association, 53(281):212–214, March 1958. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282582.

Newcomb:1957:BRB

[New57] Robinson Newcomb. Book review: [untitled]. Journal of the American Statistical Association, 52(279):399–401, September 1957. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280920.

Nicholson:1951:BRB

[Nic51] George Nicholson. Book review: [untitled]. Journal of the American Statistical Association, 46(255):393–394, September 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280523.

Nicholson:1955:SA

[Nic55] George E. Nicholson, Jr. Statistical abstracts. Journal of the American Statistical Association, 50(270):600–606, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280978.

Nicholson:1957:EPI

[Nic57] George E. Nicholson, Jr. Estimation of parameters from incomplete multivariate samples. Journal of the American Statistical Association, 52(280):523–526, December 1957. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281701.

Nisselson:1954:BRB

[Nis54] Harold Nisselson. Book review: [untitled]. Journal of the American Statistical Association, 49(266):402–403, June 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280954.

Noether:1955:URI

[Noe55] Gottfried E. Noether. Use of the range instead of the standard deviation. Journal of the American Statistical Association, 50(272): 1040–1055, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281203.

Noether:1956:BRB

[Noe56a] Gottfried E. Noether. Book review: [untitled]. Journal of the American Statistical Association, 51(276):660–661, December 1956. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281498.

Noether:1956:CUR

[Noe56b] Gottfried E. Noether. Corrigenda: Use of the range instead of the standard deviation. Journal of the American Statistical Association, 51(276):651, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281487.

Noether:1956:TST

[Noe56c] Gottfried E. Noether. Two sequential tests against trend. Journal of the American Statistical Association, 51(275):440–450, September 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2281432.

Noether:1957:BRB

[Noe57a] Gottfried E. Noether. Book review: [untitled]. Journal of the American Statistical Association, 52(277):105, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281408.

Noether:1957:TCI

[Noe57b] Gottfried E. Noether. Two confidence intervals for the ratio of two probabilities and some measures of effectiveness. Journal of the American Statistical Association, 52(277):36–45, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281398.

Noether:1959:BRB

[Noe59] Gottfried E. Noether. Book review: [untitled]. Journal of the American Statistical Association, 54(288):827–828, December 1959. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282530.

Nordin:1952:BRB

[Nor52a] J. A. Nordin. Book review: [untitled]. Journal of the American Statistical Association, 47(259):565–568, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281332.

Nordquist:1952:MMT

[Nor52b] A. V. Nordquist. Measuring magnitudes and trends in the production of livestock and meat. Journal of the American Statistical Association, 47(257):77–90, March 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279979.

Norton:1954:BRBa

[Nor54a] H. W. Norton. Book review: [untitled]. Journal of the American Statistical Association, 49(266):389–391, June 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280945.

Norton:1954:BRBb

[Nor54b] L. J. Norton. Book review: [untitled]. Journal of the American Statistical Association, 49(266):391–392, June 1954. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2280946.

Nordbotten:1955:MEE

[Nor55a] Svein Nordbotten. Measuring the error of editing the questionnaires in a census. Journal of the American Statistical Association, 50(270):364–369, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280966.

Nordin:1955:BRB

[Nor55b] J. A. Nordin. Book review: [untitled]. Journal of the American Statistical Association, 50(271):996–1000, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281192.

Notestein:1950:PWY

[Not50] Frank W. Notestein. The population of the world in the year 2000. Journal of the American Statistical Association, 45(251):335–345, September 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280291.

Nutter:1954:GM

[Nut54] G. Warren Nutter. Growth by merger. Journal of the American Statistical Association, 49(267):448–466, September 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281123.

Nutter:1958:BRB

[Nut58] G. Warren Nutter. Book review: [untitled]. Journal of the American Statistical Association, 53(283):763–765, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282066.

Nutter:1959:BRB

[Nut59] Warren Nutter. Book review: [untitled]. Journal of the American Statistical Association, 54(285):319, March 1959. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282160.

Nisselson:1959:SPH

[NW59] Harold Nisselson and Theodore D. Woolsey. Some problems of the household interview design for the national health survey. Journal of the American Statistical Association, 54(285):69–87, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282140.

Neiswanger:1959:CPE

[NY59a] W. A. Neiswanger and T. A. Yancey. Corrigenda: Parameter estimates and autonomous growth. Journal of the American Statistical Association, 54(288):812, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282514. See [NY59b].

Neiswanger:1959:PEA

[NY59b] W. A. Neiswanger and T. A. Yancey. Parameter estimates and autonomous growth. Journal of the American Statistical Association, 54(286):389–402, June 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor. org/stable/2281780. See corrigenda [NY59a].

Old:1951:BRB

[Old51a] Bruce S. Old. Book review: [untitled]. Journal of the American Statistical Association, 46(255):405–406, September 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280531.

Olds:1951:SPI

[Old51b] Edwin G. Olds. Statistics in production and inspection. Journal of the American Statistical Association, 46(255):297–310, September 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280508.

Olds:1954:EAT

[Old54] Edwin G. Olds. The experimental approach in the teaching of statistics. Journal of the American Statistical Association, 49 (268):890–896, December 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281547.

Olds:1955:BRB

[Old55] Edwin G. Olds. Book review: [untitled]. Journal of the American Statistical Association, 50(270):614–615, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280984.

Olkin:1957:BRB

[Olk57] Ingram Olkin. Book review: [untitled]. Journal of the American Statistical Association, 52(280):587–588, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281716.

ORourke:1953:BRB

[O’R53] George E. O’Rourke. Book review: [untitled]. Journal of the American Statistical Association, 48(263):661–663, September 1953. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281022.

Orcutt:1955:BRB

[Orc55] Guy H. Orcutt. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1419–1420, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281276.

Ostle:1959:CBS

[OS59] Bernard Ostle and George P. Steck. Correlation between sample means and sample ranges. Journal of the American Statistical Association, 54(286):465–471, June 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281784.

Oshima:1957:NIS

[Osh57] Harry T. Oshima. National income statistics of underdeveloped countries. Journal of the American Statistical Association, 52 (278):162–174, June 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280842.

Ostle:1954:BRB

[Ost54] Bernard Ostle. Book review: [untitled]. Journal of the American Statistical Association, 49(267):649–650, September 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281137.

Ott:1957:BRB

[Ott57] Ellis R. Ott. Book review: [untitled]. Journal of the American Statistical Association, 52(280):586–587, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281715.

Pabst:1950:BRB

[Pab50] W. R. Pabst, Jr. Book review: [untitled]. Journal of the American Statistical Association, 45(250):288–289, June 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280688.

Pabst:1955:BRB

[Pab55] William R. Pabst, Jr. Book review: [untitled]. Journal of the American Statistical Association, 50(271):988–989, September 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281187.

Paden:1950:BRB

[Pad50] Donald W. Paden. Book review: [untitled]. Journal of the American Statistical Association, 45(252):575–576, December 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280731.

Page:1957:BRB

[Pag57] Thornton Page. Book review: [untitled]. Journal of the American Statistical Association, 52(280):590–591, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281720.

Palmer:1953:BRB

[Pal53] Gladys L. Palmer. Book review: [untitled]. Journal of the American Statistical Association, 48(263):667–668, September 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281026.

Parker:1956:BRB

[Par56] William N. Parker. Book review: [untitled]. Journal of the American Statistical Association, 51(276):693–695, December 1956. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281517.

Parzen:1959:BRB

[Par59] Emanuel Parzen. Book review: [untitled]. Journal of the American Statistical Association, 54(286):525, June 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281799.

Perry:1959:BRB

[Per59] C. L. Perry. Book review: [untitled]. Journal of the American Statistical Association, 54(287):710–711, September 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282556.

Petersen:1955:BRB

[Pet55] William Petersen. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1386–1388, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281258.

Petersen:1957:BRB

[Pet57a] William Petersen. Book review: [untitled]. Journal of the American Statistical Association, 52(277):126–128, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281426.

Peterson:1957:BRB

[Pet57b] John M. Peterson. Book review: [untitled]. Journal of the American Statistical Association, 52(278):282–283, June 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280867.

Pfouts:1955:BRB

[Pfo55] R. W. Pfouts. Book review: [untitled]. Journal of the American Statistical Association, 50(270):628–629, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280994.

Pitman:1957:SS

[Pit57] E. J. G. Pitman. Statistics and science. Journal of the American Statistical Association, 52(279):322–330, September 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280902.

Pennock:1957:CES

[PJ57a] Jean L. Pennock and Carol M. Jaeger. Corrigenda: Estimating the service life of household goods by actuarial methods. Journal of the American Statistical Association, 52(280):578, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281709.

Pennock:1957:ESL

[PJ57b] Jean L. Pennock and Carol M. Jaeger. Estimating the service life of household goods by actuarial methods. Journal of the American Statistical Association, 52(278):175–185, June 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280843.

Potter:1956:BRB

[Pot56] Robert G. Potter, Jr. Book review: [untitled]. Journal of the American Statistical Association, 51(274):409–410, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281380.

Pratt:1959:RZT

[Pra59] John W. Pratt. Remarks on zeros and ties in the Wilcoxon signed rank procedures. Journal of the American Statistical Association, 54(287):655–667, September 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282543.

Press:1955:TSP

[Pre55] Harry Press. Time series problems in aeronautics. Journal of the American Statistical Association, 50(272):1022–1039, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2281202.

Price:1953:BRB

[Pri53] Daniel O. Price. Book review: [untitled]. Journal of the American Statistical Association, 48(262):370–372, June 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281300.

Price:1954:BRBb

[Pri54a] Daniel O. Price. Book review: [untitled]. Journal of the American Statistical Association, 49(268):919–921, December 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281555.

Price:1954:BRBa

[Pri54b] G. Baley Price. Book review: [untitled]. Journal of the American Statistical Association, 49(267):657–660, September 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281140.

Price:1955:ETS

[Pri55] Daniel O. Price. Examination of two sources of error in the estimation of net internal migration. Journal of the American Statistical Association, 50(271):689–700, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281160.

Proschan:1953:CTI

[Pro53] Frank Proschan. Confidence and tolerance intervals for the normal distribution. Journal of the American Statistical Association, 48 (263):550–564, September 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281009.

Politz:1950:NAG

[PS50] Alfred Politz and Willard Simmons. Note on “an attempt to get the not-at-homes into the sample without callbacks”. Journal of the American Statistical Association, 45(249):136–137, March 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280436.

Parzen:1957:BRB

[PW57] Emanuel Parzen and W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 52(279):391–393, September 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280915.

Pyke:1957:BRB

[Pyk57] Ronald Pyke. Book review: [untitled]. Journal of the American Statistical Association, 52(279):384–386, September 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280908.

Quandt:1958:EPL

[Qua58] Richard E. Quandt. The estimation of the parameters of a linear regression system obeying two separate regimes. Journal of the American Statistical Association, 53(284):873–880, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281957.

R:1950:BN

[R.50] E. R. Briefer notices. Journal of the American Statistical Association, 45(249):147–148, March 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280442.

R:1951:BRBb

[R.51] H. V. R. Book review: [untitled]. Journal of the American Statistical Association, 46(255):413, September 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280537.

R:1953:BRB

[R.53] H. V. R. Book review: [untitled]. Journal of the American Statistical Association, 48(264):916–917, December 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281089.

Raff:1951:DBU

[Raf51] Morton S. Raff. The distribution of blocks in an uncongested stream of automobile traffic. Journal of the American Statistical Association, 46(253):114–123, March 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280101.

Raff:1956:CAP

[Raf56a] Morton S. Raff. Corrigenda: On approximating the point binomial. Journal of the American Statistical Association, 51(276):651, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281488.

Raff:1956:APB

[Raf56b] Morton S. Raff. On approximating the point binomial. Journal of the American Statistical Association, 51(274):293–303, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281350.

Raff:1957:BRB

[Raf57] Morton S. Raff. Book review: [untitled]. Journal of the American Statistical Association, 52(277):106–107, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281409.

Raiffa:1955:BRB

[Rai55] Howard Raiffa. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1339–1341, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281232.

Raj:1953:EPT

[Raj53] Des Raj. Estimation of the parameters of Type III populations from truncated samples. Journal of the American Statistical Association, 48(262):336–349, June 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281291.

Raj:1956:SES

[Raj56] Des Raj. Some estimators in sampling with varying probabilities without replacement. Journal of the American Statistical Association, 51(274):269–284, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281347.

Raj:1958:RAS

[Raj58] Des Raj. On the relative accuracy of some sampling techniques. Journal of the American Statistical Association, 53(281):98–101, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2282569.

Ramachandran:1958:SSC

[Ram58a] K. V. Ramachandran. On the Studentized smallest chi-square. Journal of the American Statistical Association, 53(284):868–872, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281956.

See corrigenda [Ram59].

Ramachandran:1958:TV

[Ram58b] K. V. Ramachandran. A test of variances. Journal of the American Statistical Association, 53(283):741–747, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282052.

Ramachandran:1959:CSS

[Ram59] K. V. Ramachandran. Corrigenda: On the Studentized smallest chi-square. Journal of the American Statistical Association, 54 (288):812, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282515. See [Ram58a].

Rao:1959:NMS

[Rao59] J. N. K. Rao. A note on mean square successive differences. Journal of the American Statistical Association, 54(288):801–806, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2282504.

Rapoport:1958:BRB

[Rap58] Anatol Rapoport. Book review: [untitled]. Journal of the American Statistical Association, 53(281):254–256, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282615.

Rashley:1956:BRB

[Ras56] F. J. Rashley. Book review: [untitled]. Journal of the American Statistical Association, 51(274):404–405, June 1956. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281375.

Raup:1958:BRB

[Rau58] Phillip M. Raup. Book review: [untitled]. Journal of the American Statistical Association, 53(284):1052–1053, December 1958. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281989.

Rosander:1951:SCI

[RBJ51] A. C. Rosander, R. H. Blythe, Jr., and D. W. Johnson. Sampling 1949 corporation income tax returns. Journal of the American Statistical Association, 46(254):233–241, June 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279924.

Reed:1952:MPA

[Ree52] Lowell J. Reed. Man as a planning animal. Journal of the American Statistical Association, 47(257):1–5, March 1952. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2279974.

Reid:1952:BRB

[Rei52] Margaret G. Reid. Book review: [untitled]. Journal of the American Statistical Association, 47(258):346–347, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280770.

Reid:1953:BRB

[Rei53] Margaret G. Reid. Book review: [untitled]. Journal of the American Statistical Association, 48(261):163–165, March 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280896.

Reiter:1956:EBR

[Rei56] Stanley Reiter. Estimates of bounded relative error for the ratio of variances of normal distributions. Journal of the American Statistical Association, 51(275):481–488, September 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281437.

Reid:1958:BRB

[Rei58] Margaret G. Reid. Book review: [untitled]. Journal of the American Statistical Association, 53(281):246, March 1958. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282608.

Reid:1959:IRD

[Rei59] Margaret G. Reid. Increase in rent of dwelling units from 1940 to 1950. Journal of the American Statistical Association, 54(286):358– 376, June 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281777.

Resnikoff:1958:BRB

[Res58] George J. Resnikoff. Book review: [untitled]. Journal of the American Statistical Association, 53(284):1035–1036, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281975.

Revens:1950:BRB

[Rev50] Reuben L. Revens. Book review: [untitled]. Journal of the American Statistical Association, 45(250):296–298, June 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280693.

Rosander:1958:CUR

[RGM58a] A. C. Rosander, H. E. Guterman, and A. J. McKeon. Corrigenda: The use of random work sampling for cost analysis and control. Journal of the American Statistical Association, 53(284):1031, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281968.

Rosander:1958:URW

[RGM58b] A. C. Rosander, H. E. Guterman, and A. J. McKeon. The use of random work sampling for cost analysis and control. Journal of the American Statistical Association, 53(282):382–397, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281862.

Rice:1954:PCU

[Ric54] Stuart A. Rice. Problems of coordinating the United States statistical system. Journal of the American Statistical Association, 49 (267):438–447, September 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281122.

Rice:1956:BRB

[Ric56] Emmett J. Rice. Book review: [untitled]. Journal of the American Statistical Association, 51(276):682–685, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281512.

Richard:1957:BRB

[Ric57] Charles L. Richard. Book review: [untitled]. Journal of the American Statistical Association, 52(277):113, March 1957. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281417.

Richard:1959:BRB

[Ric59] O. W. Richard. Book review: [untitled]. Journal of the American Statistical Association, 54(286):523–524, June 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281798.

Rider:1951:DQR

[Rid51a] Paul R. Rider. The distribution of the quotient of ranges in samples from a rectangular population. Journal of the American Statistical Association, 46(256):502–507, December 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280399.

Rider:1951:DRS

[Rid51b] Paul R. Rider. The distribution of the range in samples from a discrete rectangular population. Journal of the American Statistical Association, 46(255):375–378, September 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280515.

Rider:1953:DPR

[Rid53a] Paul R. Rider. The distribution of the product of ranges in samples from a rectangular population. Journal of the American Statistical Association, 48(263):546–549, September 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281008.

Rider:1953:EDP

[Rid53b] Paul R. Rider. Errata: The distribution of the product of ranges in samples from a rectangular population. Journal of the American Statistical Association, 48(264):911, December 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281083.

Rider:1953:TPD

[Rid53c] Paul R. Rider. Truncated Poisson distributions. Journal of the American Statistical Association, 48(264):826–830, December 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281076.

Rider:1955:CTB

[Rid55a] Paul R. Rider. Corrigenda: Truncated binomial and negative binomial distributions. Journal of the American Statistical Association, 50(272):1332, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281225.

Rider:1955:DPM

[Rid55b] Paul R. Rider. The distribution of the product of maximum values in samples from a rectangular distribution. Journal of the American Statistical Association, 50(272):1142–1143, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2281211. See

corrigend [Rid56, Rid59].

Rider:1955:TBN

[Rid55c] Paul R. Rider. Truncated binomial and negative binomial distributions. Journal of the American Statistical Association, 50 (271):877–883, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281172.

Rider:1956:CDP

[Rid56] Paul R. Rider. Corrigenda: The distribution of the product of maximum values in samples from a rectangular distribution. Journal of the American Statistical Association, 51(276):651, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2281489. See

[Rid55b, Rid59].

Rider:1957:MSE

[Rid57] Paul R. Rider. The midrange of a sample as an estimator of the population midrange. Journal of the American Statistical Association, 52(280):537–542, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281703.

Rider:1959:CDP

[Rid59] Paul R. Rider. Corrigenda: “The Distribution of the Product of Maximum Values in Samples from a Rectangular Distribution”. Journal of the American Statistical Association, 54(288):811, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2282510.

See [Rid55b, Rid56].

Rios:1956:BRB

[Rio56] Sixto Rios. Book review: [untitled]. Journal of the American Statistical Association, 51(273):204, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280835.

Robson:1952:MSA

[RK52] D. S. Robson and A. J. King. Multiple sampling of attributes. Journal of the American Statistical Association, 47(258):203–215, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280744.

Robson:1953:EMS

[RK53] D. S. Robson and A. J. King. Errata: Multiple sampling of attributes. Journal of the American Statistical Association, 48(264): 911, December 1953. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2281084.

Roberts:1955:BRB

[Rob55] David R. Roberts. Book review: [untitled]. Journal of the American Statistical Association, 50(270):620–623, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280990.

Roberts:1957:BRB

[Rob57a] Harry V. Roberts. Book review: [untitled]. Journal of the American Statistical Association, 52(277):122, March 1957. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281423.

Robson:1957:AMP

[Rob57b] D. S. Robson. Applications of multivariate polykays to the theory of unbiased ratio-type estimation. Journal of the American Statistical Association, 52(280):511–522, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281700.

Roberts:1958:BRB

[Rob58] Harry V. Roberts. Book review: [untitled]. Journal of the American Statistical Association, 53(281):215–217, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282585.

Rogoff:1956:BRB

[Rog56] Natalie Rogoff. Book review: [untitled]. Journal of the American Statistical Association, 51(273):188–189, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280825.

Roose:1951:BRB

[Roo51] Kenneth D. Roose. Book review: [untitled]. Journal of the American Statistical Association, 46(254):265–266, June 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279932.

Roose:1952:BRB

[Roo52] Kenneth D. Roose. Book review: [untitled]. Journal of the American Statistical Association, 47(257):104–106, March 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279990.

Rosenbluth:1952:BRB

[Ros52] Gideon Rosenbluth. Book review: [untitled]. Journal of the American Statistical Association, 47(258):342–344, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280767.

Roshwalb:1953:EWC

[Ros53] Irving Roshwalb. Effect of weighting by card-duplication on the efficiency of survey results. Journal of the American Statistical Association, 48(264):773–777, December 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281070.

Roshwalb:1954:EEW

[Ros54] Irving Roshwalb. Errata: Effect of weighting by card-duplication on efficiency of survey results. Journal of the American Statistical Association, 49(268):906, December 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281551.

Roseman:1957:BRB

[Ros57] Herman G. Roseman. Book review: [untitled]. Journal of the American Statistical Association, 52(278):276–278, June 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280862.

Rosenblatt:1959:BRB

[Ros59] Judah Rosenblatt. Book review: [untitled]. Journal of the American Statistical Association, 54(285):323–324, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282164.

Rottenberg:1952:PIS

[Rot52] Simon Rottenberg. Problems in inter-spatial comparisons of worker efficiency: Illustrations from Puerto Rico. Journal of the American Statistical Association, 47(260):647–652, December 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280782.

Rothwell:1955:BRB

[Rot55a] Doris P. Rothwell. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1411–1413, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281272.

Rottenberg:1955:BRB

[Rot55b] Simon Rottenberg. Book review: [untitled]. Journal of the American Statistical Association, 50(271):1000–1001, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281193.

Rothwell:1958:UVS

[Rot58] Doris P. Rothwell. Use of varying seasonal weights in price index construction. Journal of the American Statistical Association, 53(281):66–77, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282566.

Rubin:1958:BRB

[Rub58] Ernest Rubin. Book review: [untitled]. Journal of the American Statistical Association, 53(283):767–768, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282068.

Rudy:1953:BRB

[Rud53] Norman Rudy. Book review: [untitled]. Journal of the American Statistical Association, 48(262):367–368, June 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281298.

Russell:1957:PER

[Rus57] Arch Russell. The President’s economic report. Journal of the American Statistical Association, 52(278):257–262, June 1957. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280851.

S:1951:BRB

[S.51] L. J. S. Book review: [untitled]. Journal of the American Statistical Association, 46(256):534, December 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280416.

S:1958:BRB

[S.58] I. R. S. Book review: [untitled]. Journal of the American Statistical Association, 53(281):227, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282595.

Sen:1954:CST

[SAF54] A. R. Sen, R. L. Anderson, and A. L. Finkner. A comparison of stratified two-stage sampling systems. Journal of the American Statistical Association, 49(267):539–558, September 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281129.

Sarhan:1955:EPS

[Sar55] A. E. Sarhan. Estimation of the parameters of a skewed distribution by linear systematic statistics. Journal of the American Statistical Association, 50(269):196–208, March 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281106.

Sarhan:1956:BRB

[Sar56] A. E. Sarhan. Book review: [untitled]. Journal of the American Statistical Association, 51(276):671–672, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281506.

Savage:1951:BRB

[Sav51a] L. J. Savage. Book review: [untitled]. Journal of the American Statistical Association, 46(255):383–384, September 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280518.

Savage:1951:TSD

[Sav51b] L. J. Savage. The theory of statistical decision. Journal of the American Statistical Association, 46(253):55–67, March 1951. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280094.

Savage:1953:BNS

[Sav53] I. Richard Savage. Bibliography of nonparametric statistics and related topics: Introduction. Journal of the American Statistical Association, 48(264):844–849, December 1953. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281078.

Savage:1954:ETS

[Sav54a] L. J. Savage. Errata: The theory of statistical decision. Journal of the American Statistical Association, 49(268):906, December 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281552.

Saville:1954:CFF

[Sav54b] Lloyd Saville. Cyclical fluctuations in foundry activity. Journal of the American Statistical Association, 49(268):853–875, December 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2281545.

Savage:1955:BRB

[Sav55] Leonard J. Savage. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1352–1354, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281238.

Savage:1956:BRB

[Sav56] Leonard J. Savage. Book review: [untitled]. Journal of the American Statistical Association, 51(274):383–385, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281360.

Savage:1957:NS

[Sav57a] I. Richard Savage. Nonparametric statistics. Journal of the American Statistical Association, 52(279):331–344, September 1957. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280903.

Savage:1957:ITR

[Sav57b] Richard Savage. On the independence of tests of randomness and other hypotheses. Journal of the American Statistical Association, 52(277):53–57, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281400.

Saville:1957:BRB

[Sav57c] Lloyd Saville. Book review: [untitled]. Journal of the American Statistical Association, 52(280):579–580, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281710.

Savage:1958:CBN

[Sav58a] I. Richard Savage. Corrigenda: Bibliography of nonparametric statistics and related topics. Journal of the American Statistical Association, 53(284):1031, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281969.

Savage:1958:BRB

[Sav58b] Leonard J. Savage. Book review: [untitled]. Journal of the American Statistical Association, 53(283):763, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282065.

Savage:1959:PMC

[Sav59a] I. Richard Savage. A production model and continuous sampling plan. Journal of the American Statistical Association, 54(285):231– 247, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2282149.

Savage:1959:BRB

[Sav59b] Leonard J. Savage. Book review: [untitled]. Journal of the American Statistical Association, 54(288):827, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282529.

Saxenian:1956:BRB

[Sax56] Hrand Saxenian. Book review: [untitled]. Journal of the American Statistical Association, 51(275):535–536, September 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281441.

Stewart:1952:BRB

[SB52] Paul W. Stewart and A. B. Blankenship. Book review: [untitled]. Journal of the American Statistical Association, 47(259):556–557, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281327.

Schumpeter:1950:BRB

[Sch50a] Joseph A. Schumpeter. Book review: [untitled]. Journal of the American Statistical Association, 45(249):140–142, March 1950.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280438.

Schustek:1950:BRB

[Sch50b] George W. Schustek. Book review: [untitled]. Journal of the American Statistical Association, 45(252):571, December 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280727.

Scheffe:1952:AVP

[Sch52a] Henry Scheffe. An analysis of variance for paired comparisons. Journal of the American Statistical Association, 47(259):381–400, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281310.

Schmitt:1952:SCM

[Sch52b] Robert C. Schmitt. Short-cut methods of estimating county population. Journal of the American Statistical Association, 47(258):232– 238, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280747.

Schaaf:1955:BRB

[Sch55] Albert H. Schaaf. Book review: [untitled]. Journal of the American Statistical Association, 50(270):627, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280993.

Schiff:1956:BRB

[Sch56a] Eric Schiff. Book review: [untitled]. Journal of the American Statistical Association, 51(275):554–555, September 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281458.

Schweiger:1956:BRB

[Sch56b] Irving Schweiger. Book review: [untitled]. Journal of the American Statistical Association, 51(275):550, September 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281455.

Schramm:1957:BRB

[Sch57a] Wilbur Schramm. Book review: [untitled]. Journal of the American Statistical Association, 52(279):410–411, September 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280926.

Schultz:1957:BRB

[Sch57b] Theodore W. Schultz. Book review: [untitled]. Journal of the American Statistical Association, 52(278):281–282, June 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280866.

Scheffe:1958:FSL

[Sch58] Henry Scheffe. Fitting straight lines when one variable is controlled. Journal of the American Statistical Association, 53(281):106–117, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2282571.

Scott:1958:BRB

[Sco58] Ira O. Scott. Book review: [untitled]. Journal of the American Statistical Association, 53(284):1046–1047, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281985.

Shiskin:1957:SAE

[SE57] Julius Shiskin and Harry Eisenpress. Seasonal adjustments by electronic computer methods. Journal of the American Statistical Association, 52(280):415–449, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281692.

Seal:1958:RPS

[Sea58] K. C. Seal. On ranking parameters of scale in Type III populations. Journal of the American Statistical Association, 53(281):164–175, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2282577.

Seal:1959:SSP

[Sea59] K. C. Seal. A single sampling plan for correlated variables with a single-sided specification limit. Journal of the American Statistical Association, 54(285):248–259, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282150.

Seiden:1956:BRB

[Sei56] Esther Seiden. Book review: [untitled]. Journal of the American Statistical Association, 51(273):205, March 1956. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2280836.

Selizer:1955:ISP

[Sel55] Lawrence H. Selizer. Interest as a source of personal income and tax revenue. Journal of the American Statistical Association, 50 (272):1248–1330, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281217.

Sewell:1951:BRB

[Sew51] William H. Sewell. Book review: [untitled]. Journal of the American Statistical Association, 46(254):274–275, June 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279938.

Sarhan:1957:TBL

[SG57] A. E. Sarhan and B. G. Greenberg. Tables for best linear estimates by order statistics of the parameters of single exponential distributions from singly and doubly censored samples. Journal of the American Statistical Association, 52(277):58–87, March 1957. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281401.

Strandskov:1951:BRB

[SH51] Herluf H. Strandskov and Bertram L. Hanna. Book review: [untitled]. Journal of the American Statistical Association, 46(253):142– 144, March 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280116.

Siegel:1952:SCU

[SH52] Jacob S. Siegel and C. Horace Hamilton. Some considerations in the use of the residual method of estimating net migration. Journal of the American Statistical Association, 47(259):475–500, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2281317.

Shapiro:1950:EBR

[Sha50] Sam Shapiro. Estimating birth registration completeness. Journal of the American Statistical Association, 45(250):261–264, June 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280684.

Shaul:1952:SSC

[Sha52a] J. R. H. Shaul. Sampling surveys in Central Africa. Journal of the American Statistical Association, 47(258):239–254, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280748.

Shaw:1952:BRB

[Sha52b] E. S. Shaw. Book review: [untitled]. Journal of the American Statistical Association, 47(257):101–103, March 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279988.

Shapiro:1957:BRB

[Sha57a] Theresa R. Shapiro. Book review: [untitled]. Journal of the American Statistical Association, 52(280):598–599, December 1957. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281727.

Shaw:1957:BRB

[Sha57b] E. S. Shaw. Book review: [untitled]. Journal of the American Statistical Association, 52(278):286–287, June 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280870.

Shellard:1952:EPS

[She52] G. D. Shellard. Estimating the product of several random variables. Journal of the American Statistical Association, 47(258):216–221, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280745.

Sheatsley:1958:BRB

[She58] Paul B. Sheatsley. Book review: [untitled]. Journal of the American Statistical Association, 53(283):769–772, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282071.

Sholes:1955:BRB

[Sho55] De Ver Sholes. Book review: [untitled]. Journal of the American Statistical Association, 50(271):990–992, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281189.

Shryock:1951:RP

[Shr51] Henry S. Shryock, Jr. Redistribution of population: 1940 to 1950. Journal of the American Statistical Association, 46(256):417–437, December 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280393.

Shryock:1954:BRB

[Shr54] Henry S. Shryock, Jr. Book review: [untitled]. Journal of the American Statistical Association, 49(266):406–408, June 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280956.

Siegel:1953:LPS

[Sie53] Irving H. Siegel. Labor productivity in the Soviet union. Journal of the American Statistical Association, 48(261):65–78, March 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280879.

Siegel:1957:BRB

[Sie57] Irving H. Siegel. Book review: [untitled]. Journal of the American Statistical Association, 52(280):593–594, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281723.

Siegel:1959:BRB

[Sie59] Jacob S. Siegel. Book review: [untitled]. Journal of the American Statistical Association, 54(288):817–820, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282520.

Silk:1956:BRB

[Sil56] Leonard Silk. Book review: [untitled]. Journal of the American Statistical Association, 51(275):544–545, September 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281451.

Silverstone:1957:ELC

[Sil57] H. Silverstone. Estimating the logistic curve. Journal of the American Statistical Association, 52(280):567–577, December 1957. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281707.

Simpson:1951:BRB

[Sim51a] Paul B. Simpson. Book review: [untitled]. Journal of the American Statistical Association, 46(255):396–398, September 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280527.

Simpson:1951:TFT

[Sim51b] Paul B. Simpson. Transformation functions in the theory of production indexes. Journal of the American Statistical Association, 46(254):225–232, June 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279923.

Simmons:1953:EIC

[Sim53a] Walt R. Simmons. The elements of an industrial classification policy. Journal of the American Statistical Association, 48(263): 429–439, September 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280999.

Simpson:1953:BRBa

[Sim53b] Paul B. Simpson. Book review: [untitled]. Journal of the American Statistical Association, 48(262):376–380, June 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281303.

Simpson:1953:BRBb

[Sim53c] Paul B. Simpson. Book review: [untitled]. Journal of the American Statistical Association, 48(262):380–381, June 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281304.

Simon:1954:SCC

[Sim54] Herbert A. Simon. Spurious correlation: a causal interpretation. Journal of the American Statistical Association, 49(267):467–479, September 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281124.

Simpson:1958:ANO

[Sim58] Paul B. Simpson. Approaches to national output measurement. Journal of the American Statistical Association, 53(284):948–962, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281962.

Sitgeaves:1953:BRB

[Sit53] Rosedith Sitgeaves. Book review: [untitled]. Journal of the American Statistical Association, 48(263):650–651, September 1953. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281016.

Sletto:1951:BRB

[Sle51] Raymond F. Sletto. Book review: [untitled]. Journal of the American Statistical Association, 46(254):275–276, June 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279939.

Sletto:1957:BRB

[Sle57] Raymond F. Sletto. Book review: [untitled]. Journal of the American Statistical Association, 52(277):128–129, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281427.

Slonim:1957:SN

[Slo57] Morris James Slonim. Sampling in a nutshell. Journal of the American Statistical Association, 52(278):143–161, June 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280841.

Slonim:1958:TDM

[Slo58] Morris James Slonim. The trentile deviation method of weather forecast evaluation. Journal of the American Statistical Association, 53(282):398–407, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281863.

Sherman:1950:SPF

[SM50] Jack Sherman and Winifred J. Morrison. Simplified procedures for fitting a Gompertz curve and a modified exponential curve. Journal of the American Statistical Association, 45(249):87–96, March 1950.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280430.

Schmid:1955:BPT

[SM55] Calvin F. Schmid and Earle H. MacCannell. Basic problems, techniques, and theory of isopleth mapping. Journal of the American Statistical Association, 50(269):220–239, March 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281108.

Slonim:1958:MAD

[SM58] Morris James Slonim and Chester H. McCall, Jr. A method of adjustment for defective data. Journal of the American Statistical Association, 53(283):736–740, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282051.

Smith:1950:EPM

[Smi50a] H. Fairfield Smith. Estimating precision of measuring instruments. Journal of the American Statistical Association, 45(251):447–451, September 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280299.

Smith:1950:BRBa

[Smi50b] John H. Smith. Book review: [untitled]. Journal of the American Statistical Association, 45(249):142–143, March 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280439.

Smith:1950:BRBb

[Smi50c] John H. Smith. Book review: [untitled]. Journal of the American Statistical Association, 45(251):457–459, September 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280303.

Smith:1952:BRB

[Smi52] M. Brewster Smith. Book review: [untitled]. Journal of the American Statistical Association, 47(257):97–98, March 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279985.

Smith:1954:TAT

[Smi54] R. Tynes Smith, III. Technical aspects of transportation flow data. Journal of the American Statistical Association, 49(266):227–239, June 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280931.

Smith:1955:BRBa

[Smi55a] H. Fairfield Smith. Book review: [untitled]. Journal of the American Statistical Association, 50(271):975–979, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281181.

Smith:1955:BRBb

[Smi55b] Walter L. Smith. Book review: [untitled]. Journal of the American Statistical Association, 50(271):986–987, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281185.

Smith:1958:BRBc

[Smi58a] Harvey L. Smith. Book review: [untitled]. Journal of the American Statistical Association, 53(283):768–769, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282070.

Smith:1958:BRBb

[Smi58b] John Smith. Book review: [untitled]. Journal of the American Statistical Association, 53(281):249–251, March 1958. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282611.

Smith:1958:BRBa

[Smi58c] Victor E. Smith. Book review: [untitled]. Journal of the American Statistical Association, 53(281):224–226, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282593.

Snedecor:1951:BRB

[Sne51] George W. Snedecor. Book review: [untitled]. Journal of the American Statistical Association, 46(254):255–257, June 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279927.

Snedecor:1955:BRB

[Sne55] George W. Snedecor. Book review: [untitled]. Journal of the American Statistical Association, 50(270):608–609, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280980.

Snyder:1956:BRB

[Sny56] Robert E. Snyder. Book review: [untitled]. Journal of the American Statistical Association, 51(276):695–696, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281518.

Steck:1959:PPD

[SO59] G. P. Steck and D. B. Owen. Percentage points for the distribution of outgoing quality. Journal of the American Statistical Association, 54(287):689–694, September 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282547.

Sobol:1959:PMP

[Sob59] Marion Gross Sobol. Panel mortality and panel bias. Journal of the American Statistical Association, 54(285):52–68, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282139.

Sogge:1954:ICU

[Sog54] Tillman M. Sogge. Industrial classes in the United States 1870 to 1950. Journal of the American Statistical Association, 49(266):251– 253, June 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280933.

Solow:1952:BRBb

[Sol52a] Robert Solow. Book review: [untitled]. Journal of the American Statistical Association, 47(260):698–700, December 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280792.

Solow:1952:BRBa

[Sol52b] Robert M. Solow. Book review: [untitled]. Journal of the American Statistical Association, 47(258):330–332, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280763.

Solomon:1955:BRBa

[Sol55a] Herbert Solomon. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1408–1409, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281270.

Solomon:1955:BRBb

[Sol55b] Herbert Solomon. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1415–1417, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281274.

Solow:1955:BRB

[Sol55c] Robert M. Solow. Book review: [untitled]. Journal of the American Statistical Association, 50(271):995–996, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281191.

Solomon:1956:BRB

[Sol56a] David Solomon. Book review: [untitled]. Journal of the American Statistical Association, 51(274):416–417, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281387.

Solow:1956:BRB

[Sol56b] Robert M. Solow. Book review: [untitled]. Journal of the American Statistical Association, 51(274):398–400, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281372.

Somerville:1957:OSB

[Som57] Paul N. Somerville. Optimum sampling in binomial population. Journal of the American Statistical Association, 52(280):494–502, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281698.

Spellman:1957:BRB

[Spe57] John R. Spellman. Book review: [untitled]. Journal of the American Statistical Association, 52(278):273, June 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280860.

Spiegelman:1951:BRB

[Spi51] Mortimer Spiegelman. Book review: [untitled]. Journal of the American Statistical Association, 46(254):271–273, June 1951. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2279936.

Spiegelman:1955:BRB

[Spi55] Mortimer Spiegelman. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1373–1374, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2281250.

Spielmans:1957:BRB

[Spi57] John V. Spielmans. Book review: [untitled]. Journal of the American Statistical Association, 52(279):407–409, September 1957. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280924.

Spoerl:1951:ASS

[Spo51] Charles A. Spoerl. Actuarial science — a survey of theoretical development. Journal of the American Statistical Association, 46 (255):334–344, September 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280511.

Sprowls:1950:SDM

[Spr50] R. Clay Sprowls. Statistical decisions by the method for minimum risk: An application. Journal of the American Statistical Association, 45(250):238–248, June 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280681.

Sprowls:1953:BRB

[Spr53] R. Clay Sprowls. Book review: [untitled]. Journal of the American Statistical Association, 48(262):368–370, June 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281299.

Sprinkel:1955:ERP

[Spr55] Beryl Wayne Sprinkel. The 1955 economic report of the president. Journal of the American Statistical Association, 50(269):240–248, March 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2281109.

Sprott:1959:BRB

[Spr59a] D. A. Sprott. Book review: [untitled]. Journal of the American Statistical Association, 54(285):331–332, March 1959. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282171.

Sprowls:1959:BRB

[Spr59b] R. Clay Sprowls. Book review: [untitled]. Journal of the American Statistical Association, 54(285):315–316, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282156.

Spurr:1951:SCM

[Spu51] William A. Spurr. A short-cut measure of correlation. Journal of the American Statistical Association, 46(253):89–94, March 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280097.

Steinhaus:1950:BRB

[Ste50] H. W. Steinhaus. Book review: [untitled]. Journal of the American Statistical Association, 45(252):584–585, December 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280736.

Steiner:1951:BRB

[Ste51a] Peter O. Steiner. Book review: [untitled]. Journal of the American Statistical Association, 46(254):270–271, June 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279935.

Steiner:1951:SBO

[Ste51b] Peter O. Steiner. A source of bias in one of the samples of the 1950 census. Journal of the American Statistical Association, 46 (253):110–113, March 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280100.

Steinberg:1952:BRB

[Ste52a] Gunther T. Steinberg. Book review: [untitled]. Journal of the American Statistical Association, 47(258):335–339, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280765.

Steiner:1952:BRB

[Ste52b] Peter O. Steiner. Book review: [untitled]. Journal of the American Statistical Association, 47(258):340–342, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280766.

Steiner:1953:BRB

[Ste53] Peter O. Steiner. Book review: [untitled]. Journal of the American Statistical Association, 48(264):923–924, December 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281091.

Stephan:1954:BRB

[Ste54] Frederick F. Stephan. Book review: [untitled]. Journal of the American Statistical Association, 49(267):672–673, September 1954. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281149.

Steinhaus:1955:BRB

[Ste55a] H. W. Steinhaus. Book review: [untitled]. Journal of the American Statistical Association, 50(270):623–625, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280991.

Stevens:1955:EBC

[Ste55b] W. L. Stevens. Estimation of the Brazilian coffee harvest by sampling survey. Journal of the American Statistical Association, 50 (271):775–787, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281165.

Steiner:1956:BRB

[Ste56] Peter O. Steiner. Book review: [untitled]. Journal of the American Statistical Association, 51(275):548–550, September 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281454.

Stegun:1957:BRB

[Ste57a] I. A. Stegun. Book review: [untitled]. Journal of the American Statistical Association, 52(278):272, June 1957. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280858.

Stewart:1957:BRB

[Ste57b] Charles D. Stewart. Book review: [untitled]. Journal of the American Statistical Association, 52(277):111–112, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281414.

Stern:1958:BRB

[Ste58] Johanna Stern. Book review: [untitled]. Journal of the American Statistical Association, 53(284):1040–1043, December 1958. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281981.

Steel:1959:MCS

[Ste59a] Robert G. D. Steel. A multiple comparison sign test: Treatments versus control. Journal of the American Statistical Association, 54 (288):767–775, December 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282500.

Sterling:1959:PDT

[Ste59b] Theodore D. Sterling. Publication decisions and their possible effects on inferences drawn from tests of significance–or vice versa. Journal of the American Statistical Association, 54(285):30–34, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2282137.

Stigler:1951:BRB

[Sti51] George J. Stigler. Book review: [untitled]. Journal of the American Statistical Association, 46(255):403–405, September 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280530.

Stigler:1952:BRB

[Sti52] George J. Stigler. Book review: [untitled]. Journal of the American Statistical Association, 47(258):344–345, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280768.

Stoller:1954:UTP

[Sto54] David S. Stoller. Univariate two-population distribution-free discrimination. Journal of the American Statistical Association, 49 (268):770–777, December 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281538.

Stoller:1955:CUT

[Sto55a] David S. Stoller. Corrigenda: Univariate two-population distribution-free discrimination. Journal of the American Statistical Association, 50(272):1332–1333, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281226.

Stolnitz:1955:UCV

[Sto55b] George J. Stolnitz. Uses of crude vital rates in the analysis of reproductivity. Journal of the American Statistical Association, 50 (272):1215–1234, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281215.

Stone:1955:BRB

[Sto55c] M. H. Stone. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1420–1422, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281277.

Stolnitz:1958:BRB

[Sto58a] George J. Stolnitz. Book review: [untitled]. Journal of the American Statistical Association, 53(282):617–618, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281895.

Stouffer:1958:KPA

[Sto58b] Samuel A. Stouffer. Karl Pearson — an appreciation on the 100th anniversary of his birth. Journal of the American Statistical Association, 53(281):23–27, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282562.

Strodtbeck:1953:BRB

[Str53] Fred L. Strodtbeck. Book review: [untitled]. Journal of the American Statistical Association, 48(263):666–667, September 1953. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281025.

Stuart:1952:PTD

[Stu52] Alan Stuart. The power of two difference-sign tests. Journal of the American Statistical Association, 47(259):416–424, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2281312.

Stuart:1954:ARE

[Stu54] Alan Stuart. Asymptotic relative efficiencies of distribution-free tests of randomness against normal alternatives. Journal of the American Statistical Association, 49(265):147–157, March 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281041.

Stuart:1956:ETR

[Stu56] Alan Stuart. The efficiencies of tests of randomness against normal regression. Journal of the American Statistical Association, 51(274):285–287, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281348.

Suits:1953:BRB

[Sui53] Daniel B. Suits. Book review: [untitled]. Journal of the American Statistical Association, 48(263):651–654, September 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281017.

Suits:1957:UDV

[Sui57] Daniel B. Suits. Use of dummy variables in regression equations. Journal of the American Statistical Association, 52(280):548–551, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281705.

Suits:1958:BRB

[Sui58] Daniel B. Suits. Book review: [untitled]. Journal of the American Statistical Association, 53(282):610–611, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281890.

Suppes:1955:BRB

[Sup55] Patrick Suppes. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1348–1352, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281237.

Suppes:1958:BRB

[Sup58] Patrick Suppes. Book review: [untitled]. Journal of the American Statistical Association, 53(283):755–756, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282057.

Swan:1954:BRB

[Swa54] A. W. Swan. Book review: [untitled]. Journal of the American Statistical Association, 49(265):184–186, March 1954. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281046.

Szatrowski:1954:BRB

[Sza54] Z. Szatrowski. Book review: [untitled]. Journal of the American Statistical Association, 49(266):385–386, June 1954. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2280943.

Taeuber:1956:BRB

[Tae56] Irene B. Taeuber. Book review: [untitled]. Journal of the American Statistical Association, 51(276):705–706, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281526.

Tate:1955:ACM

[Tat55] R. F. Tate. Applications of correlation models for biserial data. Journal of the American Statistical Association, 50(272):1078–1095, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281207.

Teichroew:1955:NAR

[Tei55] D. Teichroew. Numerical analysis research unpublished statistical tables. Journal of the American Statistical Association, 50 (270):550–556, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280975.

Teichroew:1956:BRB

[Tei56] D. Teichroew. Book review: [untitled]. Journal of the American Statistical Association, 51(274):390–391, June 1956. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281365.

Teichroew:1957:BRB

[Tei57] D. Teichroew. Book review: [untitled]. Journal of the American Statistical Association, 52(278):272–273, June 1957. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2280859.

Teper:1950:BRB

[Tep50] Lazare Teper. Book review: [untitled]. Journal of the American Statistical Association, 45(250):298–300, June 1950. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2280694.

Terry:1956:BRBa

[Ter56a] Milton E. Terry. Book review: [untitled]. Journal of the American Statistical Association, 51(274):382, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281359.

Terry:1956:BRBb

[Ter56b] Milton E. Terry. Book review: [untitled]. Journal of the American Statistical Association, 51(276):669–671, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281505.

Thompson:1952:CSM

[Tho52] Lorin A. Thompson. Comparison of selected measures of ability of communities to bear tax burdens. Journal of the American Statistical Association, 47(258):263–266, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280750.

Thompson:1959:BRB

[Tho59] Warren S. Thompson. Book review: [untitled]. Journal of the American Statistical Association, 54(285):329, March 1959. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282168.

Tick:1958:BRBa

[Tic58a] Leo J. Tick. Book review: [untitled]. Journal of the American Statistical Association, 53(281):221–222, March 1958. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282590.

Tick:1958:BRBb

[Tic58b] Leo J. Tick. Book review: [untitled]. Journal of the American Statistical Association, 53(282):626–627, June 1958. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281902.

Tintner:1958:BRBa

[Tin58a] Gerhard Tintner. Book review: [untitled]. Journal of the American Statistical Association, 53(281):219–220, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282588.

Tintner:1958:BRBb

[Tin58b] Gerhard Tintner. Book review: [untitled]. Journal of the American Statistical Association, 53(283):756–757, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282058.

Tippett:1952:BRB

[Tip52] L. H. C. Tippett. Book review: [untitled]. Journal of the American Statistical Association, 47(259):551–553, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281324.

Tippett:1955:BRB

[Tip55] L. H. C. Tippett. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1364–1365, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281244.

Tippett:1959:BRB

[Tip59] L. H. C. Tippett. Book review: [untitled]. Journal of the American Statistical Association, 54(287):711, September 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282557.

Thurstone:1957:ROM

[TJ57] L. L. Thurstone and Lyle V. Jones. The rational origin for measuring subjective values. Journal of the American Statistical Association, 52(280):458–471, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281694.

Tate:1959:OCI

[TK59] R. F. Tate and G. W. Klett. Optimal confidence intervals for the variance of a normal distribution. Journal of the American Statistical Association, 54(287):674–682, September 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282545.

Topp:1955:FJS

[TL55] Chester W. Topp and Fred C. Leone. A family of J-shaped frequency functions. Journal of the American Statistical Association, 50(269):209–219, March 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281107.

Tingey:1959:MRS

[TM59] F. H. Tingey and J. A. Merrill. Minimum risk specification limits. Journal of the American Statistical Association, 54(285):260–274, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2282151.

Tolles:1955:BRB

[Tol55] N. Arnold Tolles. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1389–1391, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281260.

Torgerson:1952:BRB

[Tor52] Warren S. Torgerson. Book review: [untitled]. Journal of the American Statistical Association, 47(258):350–351, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280773.

Toulouse:1955:SIU

[Tou55] Julian H. Toulouse. A study of industrial use of probability statistics in the physical sciences. Journal of the American Statistical Association, 50(272):1014–1021, December 1955. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281201.

Trueblood:1956:BRB

[Tru56] Robert M. Trueblood. Book review: [untitled]. Journal of the American Statistical Association, 51(275):537–538, September 1956. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281444.

Tukey:1950:CCA

[Tuk50a] John W. Tukey. Chairman’s closure. Acceptance sampling. A symposium. Journal of the American Statistical Association, 45(250): 150–155, June 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tukey:1950:SSS

[Tuk50b] John W. Tukey. Some sampling simplified. Journal of the American Statistical Association, 45(252):501–519, December 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280719.

Tukey:1952:RBE

[Tuk52] John W. Tukey. Review of The Extrapolation, Interpolation and Smoothing of Stationary Time Series with Engineering Applications, by N. Wiener. Journal of the American Statistical Association, 47(258):319–321, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280758.

Tukey:1954:UPE

[Tuk54] John W. Tukey. Unsolved problems of experimental statistics. Journal of the American Statistical Association, 49(268):706–731, December 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281535.

Tullock:1959:PDT

[Tul59] Gordon Tullock. Publication decisions and tests of significance — a comment. Journal of the American Statistical Association, 54 (287):593, September 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282539.

Turnbull:1957:BRB

[Tur57] John G. Turnbull. Book review: [untitled]. Journal of the American Statistical Association, 52(277):117–118, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281421.

Tukey:1952:RBM

[TW52] John W. Tukey and J. Wolfowitz. Review of Statistische Methoden für Naturwissenschafter, Mediziner, und Ingenieure [Statistical Methods for Natural Scientists, Physicians, and Engineers], by A. Linder. Journal of the American Statistical Association, 47 (259):554–556, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281326.

Tyler:1954:BRB

[Tyl54] Leona E. Tyler. Book review: [untitled]. Journal of the American Statistical Association, 49(266):404–406, June 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280955.

Tyler:1958:BRB

[Tyl58] Ralph W. Tyler. Book review: [untitled]. Journal of the American Statistical Association, 53(284):1039–1040, December 1958. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281980.

Ullman:1952:BRB

[Ull52] Morris B. Ullman. Book review: [untitled]. Journal of the American Statistical Association, 47(260):697–698, December 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280791.

Ulmer:1950:BRB

[Ulm50] Melville J. Ulmer. Book review: [untitled]. Journal of the American Statistical Association, 45(250):300–301, June 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280695.

Valinsky:1958:BRB

[Val58] David Valinsky. Book review: [untitled]. Journal of the American Statistical Association, 53(281):226–227, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282594.

VanVechten:1955:BRB

[Van55] C. C. Van Vechten. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1382–1383, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281255.

Villard:1952:BRB

[Vil52] Henry H. Villard. Book review: [untitled]. Journal of the American Statistical Association, 47(259):574–575, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281337.

Villard:1953:BRB

[Vil53] Henry H. Villard. Book review: [untitled]. Journal of the American Statistical Association, 48(261):159–162, March 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280894.

Vining:1951:BRB

[Vin51] Rutledge Vining. Book review: [untitled]. Journal of the American Statistical Association, 46(253):134–136, March 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280109.

Vining:1953:DEA

[Vin53] Rutledge Vining. Delimitation of economic areas: Statistical conceptions in the study of the spatial structure of an economic system. Journal of the American Statistical Association, 48(261):44– 64, March 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280878.

Voight:1952:SPP

[VK52] Robert B. Voight and Martin Kriesberg. Some principles of processing census and survey data. Journal of the American Statistical Association, 47(258):222–231, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280746.

Votaw:1952:ATP

[Vot52] D. F. Votaw, Sr. An aptitude test for predicting success in a first course in statistics. Journal of the American Statistical Association, 47(259):516–523, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281319.

W:1951:BRBa

[W.51a] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 46(254):276, June 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279940.

W:1951:BRBb

[W.51b] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 46(255):394–395, September 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280524.

W:1951:BRBc

[W.51c] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 46(255):395, September 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280525.

W:1952:BRBa

[W.52a] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 47(257):92, March 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279981.

W:1952:BRBb

[W.52b] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 47(257):96–97, March 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279984.

W:1952:BRBc

[W.52c] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 47(257):110, March 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2279994.

W:1952:BRBd

[W.52d] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 47(259):551, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281323.

W:1952:BRBe

[W.52e] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 47(259):579–580, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281339.

W:1953:BRB

[W.53] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 48(261):155, March 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280891.

W:1954:BRBa

[W.54a] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 49(265):184, March 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281045.

W:1954:BRBb

[W.54b] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 49(265):201, March 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281056.

W:1954:BRBc

[W.54c] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 49(265):201–202, March 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281057.

W:1954:BRBd

[W.54d] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 49(265):202–203, March 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281059.

W:1954:BRBe

[W.54e] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 49(267):667, September 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281145.

W:1955:BRBa

[W.55a] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 50(269):262–263, March 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281116.

W:1955:BRBb

[W.55b] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 50(269):264, March 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281118.

W:1955:BRBc

[W.55c] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 50(270):617, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280986.

W:1955:BRBd

[W.55d] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 50(271):987–988, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281186.

W:1955:BRBe

[W.55e] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 50(271):1003, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281196.

W:1956:BRBa

[W.56a] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 51(273):175–176, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280818.

W:1956:BRBb

[W.56b] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 51(274):392–393, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281367.

W:1956:BRBc

[W.56c] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 51(275):536, September 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281442.

W:1956:BRBd

[W.56d] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 51(275):538, September 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281445.

W:1956:BRBe

[W.56e] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 51(275):538–539, September 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281446.

W:1957:BRBa

[W.57a] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 52(277):111, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281413.

W:1957:BRBb

[W.57b] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 52(277):113, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281416.

W:1957:BRBc

[W.57c] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 52(279):387, September 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280910.

W:1958:BRBa

[W.58a] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 53(281):214–215, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282583.

W:1958:BRBb

[W.58b] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 53(281):215, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282584.

W:1958:BRBc

[W.58c] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 53(281):217, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282586.

W:1958:BRBd

[W.58d] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 53(281):220–221, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282589.

W:1958:BRBe

[W.58e] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 53(282):600, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281882.

W:1958:BRBf

[W.58f] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 53(282):600, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281883.

W:1958:BRBg

[W.58g] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 53(283):768, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282069.

W:1959:BRB

[W.59] W. A. W. Book review: [untitled]. Journal of the American Statistical Association, 54(285):333, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2282173.

Wagner:1958:DSL

[Wag58] Harvey M. Wagner. On the distribution of solutions in linear programming problems. Journal of the American Statistical Association, 53(281):161–163, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282576.

Wagner:1959:BRB

[Wag59a] Harvey M. Wagner. Book review: [untitled]. Journal of the American Statistical Association, 54(285):324–325, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282165.

Wagner:1959:LPT

[Wag59b] Harvey M. Wagner. Linear programming techniques for regression analysis. Journal of the American Statistical Association, 54 (285):206–212, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282146.

Walsh:1950:CBC

[Wal50a] John E. Walsh. Correction to “On the Best Choice of Sample Sizes for a t-Test When the Ratio of Variances is Known”. Journal of the American Statistical Association, 45(249):111, March 1950. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280432.

Walsh:1950:LST

[Wal50b] John E. Walsh. Large sample tests and confidence intervals for mortality rates. Journal of the American Statistical Association, 45(250):225–237, June 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280680.

Walsh:1951:LSS

[Wal51] John E. Walsh. A large sample t-statistic which is insensitive to non-randomness. Journal of the American Statistical Association, 46(253):79–88, March 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280096.

Walsh:1952:OCT

[Wal52a] John E. Walsh. Operating characteristics for tests of the stability of a normal population. Journal of the American Statistical Association, 47(258):191–202, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280743.

Walsh:1952:SNT

[Wal52b] John E. Walsh. Some nonparametric tests for Student’s hypothesis in experimental designs. Journal of the American Statistical Association, 47(259):401–415, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281311.

Walsh:1954:BRB

[Wal54] John E. Walsh. Book review: [untitled]. Journal of the American Statistical Association, 49(266):408–409, June 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280957.

Wallace:1955:BRB

[Wal55] David L. Wallace. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1367–1368, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281246.

Walker:1956:BRB

[Wal56a] Helen M. Walker. Book review: [untitled]. Journal of the American Statistical Association, 51(273):163–165, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280812.

Wallace:1956:BRB

[Wal56b] David A. Wallace. Book review: [untitled]. Journal of the American Statistical Association, 51(274):417–418, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281388.

Walsh:1956:AEN

[Wal56c] John E. Walsh. Asymptotic efficiencies of a nonparametric life test for smaller percentiles of a gamma distribution. Journal of the American Statistical Association, 51(275):467–480, September 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2281436.

Walker:1957:BRB

[Wal57] Charls E. Walker. Book review: [untitled]. Journal of the American Statistical Association, 52(277):114–115, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281419.

Walker:1958:CKP

[Wal58a] Helen M. Walker. The contributions of Karl Pearson. Journal of the American Statistical Association, 53(281):11–22, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2282561.

Wallace:1958:BRB

[Wal58b] D. L. Wallace. Book review: [untitled]. Journal of the American Statistical Association, 53(283):757, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282059.

Wallace:1959:SBA

[Wal59a] David L. Wallace. Simplified beta-approximations to the Kruskal– Wallis H test. Journal of the American Statistical Association, 54(285):225–230, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282148.

Walsh:1959:CSS

[Wal59b] John E. Walsh. Comments on “the simplest signed-rank tests”. Journal of the American Statistical Association, 54(285):213–224, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2282147.

Warntz:1956:MSA

[War56] William Warntz. Measuring spatial association with special consideration of the case of market orientation of production. Journal of the American Statistical Association, 51(276):597–604, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2281472.

Watkins:1956:CF

[Wat56] Ralph J. Watkins. Confession of faith: 1955. Journal of the American Statistical Association, 51(273):1–11, March 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280800.

Webber:1959:BRB

[Web59] Irving L. Webber. Book review: [untitled]. Journal of the American Statistical Association, 54(285):330–331, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282170.

Weil:1950:NDI

[Wei50] Ulric H. Weil. A note on the derivation of income estimates by source of income of persons making less than $500 per annum, 1944–1948. Journal of the American Statistical Association, 45 (251):439–446, September 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280298.

Weimer:1951:BRB

[Wei51] A. M. Weimer. Book review: [untitled]. Journal of the American Statistical Association, 46(253):138–140, March 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280112.

Weiler:1953:URC

[Wei53] H. Weiler. The use of runs to control the mean in quality control. Journal of the American Statistical Association, 48(264):816–825, December 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281075.

Weiler:1954:NTC

[Wei54] H. Weiler. A new type of control chart limits for means, ranges, and sequential runs. Journal of the American Statistical Association, 49(266):298–314, June 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280936.

Weinstein:1955:IEL

[Wei55] Abbott S. Weinstein. Increasing the effective length of short timeseries for the purpose of estimating autoregressive parameters. Journal of the American Statistical Association, 50(271):909–922, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281176.

Weiler:1956:CUR

[Wei56] H. Weiler. Corrigenda: The use of runs to control the mean in quality control. Journal of the American Statistical Association, 51 (276):652, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281490.

Weintraub:1957:BRB

[Wei57] Robert E. Weintraub. Book review: [untitled]. Journal of the American Statistical Association, 52(280):611–612, December 1957. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281734.

Weinstein:1958:ADS

[Wei58] Abbott S. Weinstein. Alternative definitions of the serial correlation coefficient in short autoregressive sequences. Journal of the American Statistical Association, 53(284):881–892, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281958.

Weingartner:1959:BRB

[Wei59] H. M. Weingartner. Book review: [untitled]. Journal of the American Statistical Association, 54(285):322–323, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282163.

Welch:1954:BRB

[Wel54] Bernard L. Welch. Book review: [untitled]. Journal of the American Statistical Association, 49(266):378–379, June 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280940.

Welch:1956:LCS

[Wel56] B. L. Welch. On linear combinations of several variances. Journal of the American Statistical Association, 51(273):132–148, March 1956.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280809.

Welch:1957:BRB

[Wel57] B. L. Welch. Book review: [untitled]. Journal of the American Statistical Association, 52(280):588–589, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281717.

Welch:1958:SSS

[Wel58] B. L. Welch. ‘student’ and small sample theory. Journal of the American Statistical Association, 53(284):777–788, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281951.

Weston:1951:BRB

[Wes51] J. F. Weston. Book review: [untitled]. Journal of the American Statistical Association, 46(253):137–138, March 1951. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2280111.

West:1952:RVC

[Wes52] Vincent I. West. Replacing variables in correlation problems. Journal of the American Statistical Association, 47(258):185–190, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280742.

Westoff:1955:BRB

[Wes55] Charles F. Westoff. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1374–1376, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281251.

Westfield:1958:BRB

[Wes58] Fred M. Westfield. Book review: [untitled]. Journal of the American Statistical Association, 53(283):751–752, September 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282054.

Waugh:1957:GC

[WF57] Frederick V. Waugh and Karl A. Fox. Graphic computation of R_{1.23}. Journal of the American Statistical Association, 52(280):

479–481, December 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281696.

Waugh:1958:CGC

[WF58] Frederick V. Waugh and Karl A. Fox. Corrigenda: Graphic computation of R_{1.23}. Journal of the American Statistical Association, 53(284):1031, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281970.

Whelpton:1950:CMK

[Whe50] P. K. Whelpton. Comments on Mr. Karmel’s note. Journal of the American Statistical Association, 45(249):125–135, March 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280435.

White:1954:ESA

[Whi54] Helen R. White. Empirical study of the accuracy of selected methods of projecting state populations. Journal of the American Statistical Association, 49(267):480–498, September 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281125.

Wickens:1953:SPI

[Wic53] Aryness Joy Wickens. Statistics and the public interest. Journal of the American Statistical Association, 48(261):1–14, March 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280875.

Williams:1950:CNW

[Wil50] C. Arthur Williams, Jr. On the choice of the number and width of classes for the chi-square test of goodness of fit. Journal of the American Statistical Association, 45(249):77–86, March 1950. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280429.

Wilks:1951:USE

[Wil51] S. S. Wilks. Undergraduate statistical education. Journal of the American Statistical Association, 46(253):1–18, March 1951. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280089.

Wilcox:1952:BRB

[Wil52] Clair Wilcox. Book review: [untitled]. Journal of the American Statistical Association, 47(257):100–101, March 1952. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2279987.

Willcox:1954:MAS

[Wil54] Walter F. Willcox. Methods of apportioning seats in the house of representatives. Journal of the American Statistical Association, 49(268):685–695, December 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281533.

Williams:1955:RHP

[Wil55a] Robert M. Williams. The relationship of housing prices and building costs in los angeles, 1900- 1953. Journal of the American Statistical Association, 50(270):370–376, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280967.

Wilson:1955:BRB

[Wil55b] Edwin B. Wilson. Book review: [untitled]. Journal of the American Statistical Association, 50(272):1354–1357, December 1955. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281239.

Wilk:1956:BRB

[Wil56a] M. B. Wilk. Book review: [untitled]. Journal of the American Statistical Association, 51(275):536–537, September 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281443.

Willcox:1956:CMA

[Wil56b] Walter F. Willcox. Corrigenda: Methods of apportioning seats in the house of representatives. Journal of the American Statistical Association, 51(276):652, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281492.

Willis:1959:LBF

[Wil59] Richard H. Willis. Lower bound formulas for the mean intercorrelation coefficient. Journal of the American Statistical Association, 54(285):275–280, March 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2282152.

Windle:1959:ACL

[Win59] Charles Windle. The accuracy of census literary statistics in Iran. Journal of the American Statistical Association, 54(287):578–581, September 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2282537.

Wiseman:1958:BRB

[Wis58] J. Wiseman. Book review: [untitled]. Journal of the American Statistical Association, 53(281):237–238, March 1958. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2282603.

Withey:1956:BRB

[Wit56] S. B. Withey. Book review: [untitled]. Journal of the American Statistical Association, 51(273):194–195, March 1956. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2280829.

Wilk:1955:FMR

[WK55] M. B. Wilk and O. Kempthorne. Fixed, mixed, and random models. Journal of the American Statistical Association, 50(272):1144–1167, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281212.

Wilk:1956:CFM

[WK56] M. B. Wilk and O. Kempthorne. Corrigenda: Fixed, mixed, and random models. Journal of the American Statistical Association, 51(276):652, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281491.

Wilk:1957:NAL

[WK57] M. B. Wilk and Oscar Kempthorne. Non-additives in a Latin square design. Journal of the American Statistical Association, 52 (278):218–236, June 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280847.

Wolfe:1951:BRB

[Wol51] Cedric Wolfe. Book review: [untitled]. Journal of the American Statistical Association, 46(255):402–403, September 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280529.

Wolfe:1952:BRB

[Wol52] Cedric Wolfe. Book review: [untitled]. Journal of the American Statistical Association, 47(259):568–569, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281333.

Wolfle:1955:BRB

[Wol55] Dael Wolfle. Book review: [untitled]. Journal of the American Statistical Association, 50(270):619, June 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280988.

Wolman:1959:BRB

[Wol59] William Wolman. Book review: [untitled]. Journal of the American Statistical Association, 54(286):518–520, June 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281794.

Woodbury:1951:RASa

[Woo51a] Max A. Woodbury. Reprints of abstracts in statistical methodology. Journal of the American Statistical Association, 46(254):251–254, June 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279926.

Woodbury:1951:RASb

[Woo51b] Max A. Woodbury. Reprints of abstracts in statistical methodology. Journal of the American Statistical Association, 46(255):379–380, September 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280516.

Woodbury:1951:RASc

[Woo51c] Max A. Woodbury. Reprints of abstracts in statistical methodology. Journal of the American Statistical Association, 46(256):516, December 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280402.

Woodbury:1952:RASa

[Woo52a] Max A. Woodbury. Reprints of abstracts in statistical methodology. Journal of the American Statistical Association, 47(257):91, March 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2279980.

Woodbury:1952:RASb

[Woo52b] Max A. Woodbury. Reprints of abstracts in statistical methodology. Journal of the American Statistical Association, 47(258):313–314, June 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280755.

Woodbury:1952:RASc

[Woo52c] Max A. Woodbury. Reprints of abstracts in statistical methodology. Journal of the American Statistical Association, 47(259):550, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2281322.

Woodruff:1952:CIM

[Woo52d] Ralph S. Woodruff. Confidence intervals for medians and other position measures. Journal of the American Statistical Association, 47(260):635–646, December 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2280781.

Woytinsky:1950:BRB

[Woy50] W. S. Woytinsky. Book review: [untitled]. Journal of the American Statistical Association, 45(250):301–303, June 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280696.

Wright:1951:BRB

[Wri51] C. Ashley Wright. Book review: [untitled]. Journal of the American Statistical Association, 46(256):518–521, December 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280404.

Wright:1952:BRB

[Wri52] William W. Wright. Book review: [untitled]. Journal of the American Statistical Association, 47(257):94–96, March 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2279983.

Wright:1956:BRB

[Wri56] Kenneth M. Wright. Book review: [untitled]. Journal of the American Statistical Association, 51(275):555–557, September 1956. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2281459.

Wrong:1956:BRB

[Wro56] Dennis H. Wrong. Book review: [untitled]. Journal of the American Statistical Association, 51(274):411–412, June 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281382.

Waugh:1953:PB

[WW53] Dan F. Waugh and Frederick V. Waugh. On probabilities in bridge. Journal of the American Statistical Association, 48(261):79–87, March 1953. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280880.

Wyckoff:1958:BRB

[Wyc58] Viola Wyckoff. Book review: [untitled]. Journal of the American Statistical Association, 53(282):609–610, June 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281889.

Yates:1951:ISM

[Yat51] F. Yates. The influence of statistical methods for research workers on the development of the science of statistics. Journal of the American Statistical Association, 46(253):19–34, March 1951. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280090.

Yeracaris:1955:DMG

[Yer55] Constantine A. Yeracaris. Differential mortality, general and causespecific in buffalo, 1939-41. Journal of the American Statistical Association, 50(272):1235–1247, December 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281216.

Yeracaris:1956:CDM

[Yer56] Constantine A. Yeracaris. Corrigenda: Differential mortality, general and cause-specific in buffalo, 1939-41. Journal of the American Statistical Association, 51(276):652, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281493.

Yerushalmy:1957:BRB

[Yer57] J. Yerushalmy. Book review: [untitled]. Journal of the American Statistical Association, 52(277):124–126, March 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281425.

Yoder:1952:BRB

[Yod52] Dale Yoder. Book review: [untitled]. Journal of the American Statistical Association, 47(259):573–574, September 1952. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281336.

Youden:1951:FRM

[You51] W. J. Youden. The Fisherian revolution in methods of experimentation. Journal of the American Statistical Association, 46(253):47– 50, March 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-

274X (electronic). URL http://www.jstor.org/stable/2280092.

Zakrzewski:1956:PVI

[Zak56] Gustave Zakrzewski. Practical value of international educational statistics. Journal of the American Statistical Association, 51(276): 605–614, December 1956. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2281473.

Zarkovic:1954:SCL

[Zar54] S. S. Zarkovic. Sampling control of literacy data. Journal of the American Statistical Association, 49(267):510–519, September 1954. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2281127.

Zarkovic:1955:SMY

[Zar55] S. S. Zarkovic. Sampling methods in the Yugoslav 1953 census of population. Journal of the American Statistical Association, 50 (271):720–737, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2281162.

Zeisel:1959:BRB

[Zei59] Hans Zeisel. Book review: [untitled]. Journal of the American Statistical Association, 54(286):511, June 1959. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2281789.

Zelen:1957:AIB

[Zel57] Marvin Zelen. The analysis of incomplete block designs. Journal of the American Statistical Association, 52(278):204–217, June 1957. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2280846.

Zellner:1958:SAP

[Zel58] Arnold Zellner. A statistical analysis of provisional estimates of gross national product and its components, of selected national income components, and of personal saving. Journal of the American Statistical Association, 53(281):54–65, March 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2282565.

Ziola:1955:BRB

[Zio55] R. Ziola. Book review: [untitled]. Journal of the American Statistical Association, 50(271):1001–1002, September 1955. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2281194.

Zobel:1950:MPL

[Zob50] Sigmund P. Zobel. On the measurement of the productivity of labor. Journal of the American Statistical Association, 45(250):218–224, June 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2280679.

Zucker:1957:BRB

[Zuc57] Ruth Zucker. Book review: [untitled]. Journal of the American Statistical Association, 52(280):590, December 1957. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2281719.